### v1

# Orthogonal roots, quantum Hafnians, and generalized Rothe diagrams

Abstract.
Let $U$ be a set of positive roots of type $ADE$, and let $\Omega_{U}$ be the set of all maximum cardinality orthogonal subsets of $U$. For each element $R\in\Omega_{U}$, we define a generalized Rothe diagram whose cardinality we call the level, $\rho(R)$, of $R$. We define the generalized quantum Hafnian of $U$ to be the generating function of $\rho$, regarded as a $q$-polynomial in $U$. Several widely studied algebraic and combinatorial objects arise as special cases of these constructions, and in many cases, $\Omega_{U}$ has the structure of a graded partially ordered set with rank function $\rho$. A motivating example of the construction involves a certain set of $k^{2}$ roots in type $D_{2k}$, where the elements of $\Omega_{U}$ correspond to permutations in $S_{k}$, the generalized Rothe diagrams are the traditional Rothe diagrams associated to permutations, the level of a permutation is its length, the generalized quantum Hafnian is the $q$-permanent, and the partial order is the Bruhat order. We exhibit many other natural examples of this construction, including one involving perfect matchings, two involving labelled Fano planes, and one involving the invariant cubic form in type $E_{6}$.
Key words and phrases:
root system, Weyl group, Rothe diagram, quantum Hafnian, quasiparabolic set
1991 Mathematics Subject Classification:
Primary: 17B22; Secondary: 05E18, 20F55.
1. Introduction
Let $U$ be a subset of the positive roots of a Weyl group of type $ADE$, and let $\Omega_{U}$ be the set of orthogonal subsets of $U$ of maximal cardinality. For each $R\in\Omega_{U}$, we define a residue of $R$ with respect to $U$ to be an element $\gamma\in U$ such that $s_{\beta}(\gamma)$ is positive for all $\beta\in R$, where $s_{\beta}$ denotes the reflection associated to the vector $\beta$. The generalized Rothe diagram of $R$ is the set of residues of $R$ in $U$, which we denote by $\text{\rm Res}_{U}(R)$. We call $\rho(R):=|\text{\rm Res}_{U}(R)|$ the level of $R$, and we define the generalized quantum Hafnian of $U$ to be the element of ${\mathbb{Z}}[U][q]$ given by
$\sum_{R\in\Omega_{U}}q^{\rho(R)}\prod_{\beta_{i}\in R}x_{\beta_{i}},$
where the $x_{\beta}$ are commuting indeterminates indexed by the set $U$.
Although it is not a priori clear why one might want to study these definitions, which are new to the best of our knowledge, we will show that many choices for $U$ give rise to concise constructions of known algebraic and combinatorial objects, in some cases endowing them with additional structure. For example, we will use our methods to associate a generalized Rothe diagram in a canonical way to each of the $45$ terms of the invariant cubic polynomial in type $E_{6}$, which gives a nonrecursive combinatorial interpretation of the coefficients of the polynomial.
One of the main motivating examples of this phenomenon (Theorem 3.13) is the case where $U=\{\varepsilon_{i}+\varepsilon_{k+j}:1\leq i,j\leq k\}$ is a set of $k^{2}$ roots of type $D_{2k}$, corresponding to the positions $(i,j)$ on a $k\times k$ grid. Here, the set $\Omega_{U}$ corresponds to the set of configurations of $k$ nonattacking rooks, or equivalently, the elements of the symmetric group $S_{k}$. If we place the origin at the top left, then a vacant position $(a,b)$ is a residue for $R\in\Omega_{U}$ unless it is attacked from above or from its left; see Figure 1 and Remark 3.14 (i). The set of such positions is a well-studied object known as the Rothe diagram of the permutation $\tau$ associated to $R$, and it is a well-known result that the cardinality of the Rothe diagram is equal to the length of $\tau$. In this case, the level function is the length of $\tau$, the quantum Hafnian is the $q$-permanent of a $k\times k$ matrix $A$, and specializing $q$ to q-q gives the $q$-determinant of $A$.
$11223344\bigstar\bigstar\bigstar$
Figure 1. The Rothe diagram of $\underline{3142}\in S_{4}$, with residues marked by stars
Another motivating example (Theorem 3.10) is the case where $U=\{\varepsilon_{i}+\varepsilon_{j}:1\leq i<j<2k\}$ is a set of $\binom{2k}{2}$ roots of type $D_{2k}$ corresponding to the $2$-subsets of the set $2k]:=\{1,2,\ldots,2k\}$. Here, the set $\Omega_{U}$ corresponds to the set of perfect matchings of the set $[2k]$, and a pair $\{a,b\}$ (where $1\leq a<b\leq 2k$) is a residue of $R\in\Omega_{U}$ if and only if we have both $\tau(a)>b$ and $\tau(b)>a$, where $\tau$ is the fixed point free involution corresponding to the perfect matching. The residues in this case are the Rothe diagrams of fixed point free involutions as defined by Hamaker–Marberg–Pawlowski [9, (3.3)], the level function is a well-known statistic on perfect matchings, and the quantum Hafnian is precisely the $q$-Hafnian defined by Jing–Zhang [13]. Specializing $q$ to q-q gives the quantum Pfaffian defined by Strickland [16], and specializing $q$ to +1+1 and to 1-1 gives the Hafnian and Pfaffian of a matrix, respectively. Recall that the Pfaffian, $\text{\rm Pf}(A)$, of a skew-symmetric $2k\times 2k$ matrix $A$ is a polynomial in the entries of $A$ with the property that $(\text{\rm Pf}(A))^{2}=\det(A)$.
The two examples just mentioned have the additional property that the level function $\rho$ endows $\Omega_{U}$ with the structure of a quasiparabolic set in the sense of Rains–Vazirani [14]. This equips $\Omega_{U}$ with a $W_{I}$-action for some standard parabolic subgroup $W_{I}$ of the Weyl group $W$, and with a graded partial order whose rank function is $\rho$. In the first example, this produces the Bruhat order on $S_{k}$.
In Section 4, we construct many other examples of sets of roots $U$ whose level function gives $\Omega_{U}$ the structure of a quasiparabolic set. Our construction seems to be an unreasonably effective method of producing a quasiparabolic set from a set of roots. For root systems of types $D_{2k}$, $E_{7}$, or $E_{8}$, we can form a quasiparabolic set by taking $U={\Phi_{+}}$ to be the set of all positive roots (Theorem 4.3). We also discuss an abundance of examples of quasiparabolic sets $\Omega_{U}$ where $U=U(\Gamma,p)$ is a set of roots associated to a suitable simple root $\alpha_{p}$ in a simply laced root system. These include examples corresponding to every simple root in every root system of type $A$ or $D$ (Example 4.7, Remark 4.8 (i)), as well as a total of $12$ examples in type $E$ (Example 4.9 (i)). The residue sets for the examples of type $A$ and $D$ have interesting combinatorial interpretations in terms of rook configurations, generalizing the aforementioned motivating examples. One of the examples in type $E_{7}$, associated to the set $U=U(E_{7},2)$, corresponds in a precise way to the $30$ inequivalent $7$-labellings of the Fano plane, and one example in type $E_{8}$, $U=U(E_{8},2)$, corresponds to the $240$ packings of the projective space $PG(3,2)$.
In Section 5, we focus on two important examples arising from the construction of Section 4 in type $E$. The set $U$ in these cases are given by a certain set $U(E_{8},8)$ of $56$ roots in type $E_{8}$ and a certain set $U(E_{7},7)$ of $27$ roots in type $E_{7}$, both of which appear naturally as the set of weights of a suitable minuscule representation. The set $U=U(E_{7},7)$ gives rise to a quasiparabolic set $\Omega_{U}$ whose elements correspond to tritangent planes on a del Pezzo surface of degree 3 and whose Pfaffian gives the well-known invariant cubic polynomial in type $E_{6}$ (Theorem 5.19). The quasiparabolic order on $\Omega_{U}$ endows the tritangent planes with a natural partial order and yields a non-recursive interpretation of the coefficients of the invariant polynomial, both of which are new to our knowledge.
The paper is organized as follows. Section 2 recalls the necessary background regarding the combinatorics of root systems. In Section 3, we introduce the key definitions and summarize the theory of quasiparabolic sets developed by Rains–Vazirani, and we give details of the two motivating examples. Section 4 exhibits many other examples of our construction, including the cases where $U$ is the set of all positive roots of type $E_{7}$, $E_{8}$, or $D_{n}$ for $n$ even. Section 5 is an in-depth study of the two examples in type $E$ mentioned in the last paragraph. Section 6 discusses some possible directions for further research.
2. Orthogonal roots
2.1. Review of root systems
We recall the necessary background concerning root systems in this subsection. While many definitions in this paper work more generally for all (finite) root systems, we will focus on irreducible simply laced root systems, i.e., on the root systems of types $ADE$. We label the vertices in the Dynkin diagrams of root systems as shown in Figure 2 until Section 5.2, where we will switch to a different convention in type $E$ for reasons we will explain.
1122n1n-1nn
(a) $A_{n}$
1122n3n-3n2n-2n1n-1nn
(b) $D_{n}$
113344556622
(c) $E_{6}$
11334455667722
(d) $E_{7}$
1122334455667788
(e) $E_{8}$
Figure 2. Dynkin diagrams of irreducible simply laced Weyl groups
Let $\Phi=\Phi(\Gamma)$ be a simply laced root system with Dynkin diagram $\Gamma$. The vertices of $\Gamma$ index the simple roots $\Pi=\{\alpha_{i}:i\in\Gamma\}$, and the rank of $\Phi$ is $\text{\rm rank}(\Phi):=\lvert\Pi\rvert$. The root lattice ${\mathbb{Z}}\Pi$ is the free $\mathbb{Z}$}-module on $\Pi$, and it carries a symmetric $\mathbb{Z}$}-bilinear form $B$ given by
[…]
If $\alpha_{i}\in\Pi$, then the simple reflection $s_{i}$ is the self-inverse $\mathbb{Z}$}-linear operator ${\mathbb{Z}}\Pi\rightarrow{\mathbb{Z}}\Pi$ given by
(2.2) $s_{i}(\beta)=\beta-B(\alpha_{i},\beta)\alpha_{i}.$
The Weyl group $W=W(\Gamma)$ is the finite group generated by the set $S=S(\Gamma)=\{s_{i}:i\in\Gamma\}$, and we have $\Phi=\{w(\alpha_{i}):\alpha_{i}\in\Pi,w\in W\}$. Each element of $\Phi$ is called a root, and each root gives rise to a reflection, which is the self-inverse linear map $s_{\alpha}:{\mathbb{Z}}\Pi\rightarrow{\mathbb{Z}}\Pi$ given by the formula
(2.3) $s_{\alpha}(\beta)=\beta-B(\alpha,\beta)\alpha,$
generalizing Equation (2.2). We have $s_{\alpha}\in W$ for all roots $\alpha\in\Phi$. Moreover, the Weyl group acts transitively on the roots and the reflections $\{s_{\alpha}:\alpha\in\Phi\}$ form a single conjugacy class in $W$. The reflection representation of $W$ is the ${\mathbb{Q}}W$-module $V={\mathbb{Q}}\otimes_{\mathbb{Z}}{\mathbb{Z}}\Pi$, where each reflection acts by Equation (2.3). The simple roots form a basis of $V$, so that we have $\dim(V)=\text{\rm rank}(\Phi)$.
The pair $(W,S)$ is a Coxeter system, and the set of reflections $\{s_{\alpha}:\alpha\in\Phi\}$ coincides with the set $T=\{wsw^{-1}:s\in S,w\in W\}$. For each subset $I\subseteq S$, the standard parabolic subgroup $W_{I}\leq W$ is the subgroup $\langle I\rangle$ generated by $I$.
Every root is of the form $\alpha=\sum_{i\in\Gamma}c_{i}\alpha_{i}$ for some integers $c_{i}$ that are either all all nonnegative or all nonpositive. We say $\alpha$ is positive or negative in these cases, respectively. The sum of the coefficients $\sum_{i\in\Gamma}c_{i}$ is called the height of $\alpha$. Given any $\alpha\in\Phi$, the only multiples of $\alpha$ in $\Phi$ are $\alpha$ and $\alpha$.
There is a partial order on the root lattice ${\mathbb{Z}}\Pi$ defined by the condition that $\alpha\leq\beta$ if and only if $\beta-\alpha$ is a nonnegative linear combination of simple roots, so that a root $\alpha$ is positive if and only if $\alpha>0$ in this order. With respect to (the restriction of) this partial order, $\Phi$ has a unique maximal element, $\theta$, called the highest root. We denote the set of positive roots $\{\alpha\in\Phi:\alpha>0\}$ by $\Phi_{+}$}.
A subsystem of a root system $\Phi$ is a subset $\Psi\subseteq\Phi$ that is itself a root system. Equivalently, a subsystem of $\Phi$ is a nonempty subset that is closed under the operations $\{s_{\alpha}:\alpha\in\Psi\}$. The positive roots of $\Phi$ naturally induce a positive system, $\Psi_{+}$, and a simple system, $\Pi_{\Psi}$, in $\Psi$ (in the sense of [10, Section 1.3]; see also [8, Section 2.2]). We have $\Psi_{+}=\Psi\cap\Phi_{+}$, and $\Pi_{\Psi}$ can be characterized as the set the elements of $\Psi_{+}$ that cannot be expressed as nonnegative linear combinations of other elements of $\Psi_{+}$. The elements of $\Pi_{\Psi}$ may or may not be elements of $\Pi$, although it is true that $\Pi\cap\Psi\subseteq\Pi_{\Psi}$. We say that $\Psi$ is of type $\Gamma^{\prime}$ if $\Psi$ is isomorphic to a root system of type $\Gamma^{\prime}$.
The bilinear form $B$ on ${\mathbb{Z}}\Pi$ is nondegenerate. We have $B(\alpha,\beta)\in\{2,1,0,-1,-2\}$ for all roots $\alpha,\beta\in\Phi$ , with $B(\alpha,\beta)=2$ if and only if $\alpha=\beta$ ([10, §6.6]). We say two roots $\alpha$ and $\beta$ are orthogonal if $B(\alpha,\beta)=0.$ The roots orthogonal to a fixed root in any root system $\Phi$ naturally form a subsystem, and it is known that when $\Phi$ has type $E_{8}$, $E_{7}$, $E_{6}$, and $D_{n}$ for $n>4$, such a subsystem has type $E_{7}$, $D_{6}$, $A_{5}$, and $A_{1}+D_{n-2}$, respectively, for all roots in $\Phi$ ([7, Theorem 2.5]). The form $B$ is $W$-invariant in the sense that $B(\alpha,\beta)=B(w(\alpha),w(\beta))$ for all $w\in W$ and all $\alpha,\beta\in\Phi$. In particular, the action of $W$ on $\Phi$ preserves orthogonality.
The root systems of types $A$, $D$, and $E$ have the following well-known coordinate representations.
Let $\varepsilon_{1},\ldots,\varepsilon_{n}$ be the usual standard basis of the Euclidean space ${\mathbb{R}}^{n}$. The vectors $\{\varepsilon_{i}-\varepsilon_{j}:1\leq i\neq j\leq n\}$ form a root system of type $A_{n-1}$. The simple roots $\Pi=\{\alpha_{1},\alpha_{2},\ldots,\alpha_{n-1}\}$ are given by $\alpha_{i}=\varepsilon_{i}-\varepsilon_{i+1}$. A root $\varepsilon_{i}-\varepsilon_{j}$ is positive if $i<j$ and negative if $i>j$. The highest root is $\varepsilon_{1}-\varepsilon_{n}$. The Weyl group is isomorphic to the symmetric group $S_{n}$ and acts by permuting the basis $\varepsilon_{1},\ldots,\varepsilon_{n}$. The bilinear form $B$ is the usual dot product. Two roots are orthogonal if and only if they have disjoint support, where by the support of any element $v=\sum_{i=1}^{n}c_{i}\varepsilon_{i}\in{\mathbb{R}}^{n}$ we mean the set $\{i:c_{i}\neq 0\}$.
The vectors $\{\pm\varepsilon_{i}\pm\varepsilon_{j}:1\leq i<j\leq n\}$ form a root system of type $D_{n}$. The simple roots $\Pi=\{\alpha_{1},\alpha_{2},\ldots,\alpha_{n}\}$ are given by $\alpha_{i}=\varepsilon_{i}-\varepsilon_{i+1}$ for $i<n$ and $\alpha_{n}=\varepsilon_{n-1}+\varepsilon_{n}$. If $i<j$ then the roots $\varepsilon_{i}\pm\varepsilon_{j}$ are positive, and the roots $\varepsilon_{i}\pm\varepsilon_{j}$ are negative. The highest root is $\varepsilon_{1}+\varepsilon_{2}$. The Weyl group has order $2^{n-1}n!$ and acts by signed permutations of the orthonormal basis, with the restriction that each element effects an even number of sign changes. The bilinear form $B$ is the usual dot product. Two roots $\alpha$ and $\beta$ are orthogonal if and only if either (a) $\alpha$ and $\beta$ have disjoint support or (b) $\alpha$ and $\beta$ have the same support and $\alpha\neq\pm\beta$.
Note that the root system of type $A_{n-1}$ may be naturally identified as the subsystem $\Phi(D_{n})\cap{\mathbb{Z}}(\Pi\setminus\{\alpha_{n}\})$ of $\Phi(D_{n})$, and $W(A_{n-1})$ as the parabolic subgroup $W_{I}$ of the Weyl group $W=W(D_{n})$ where $I=S(D_{n})\setminus\{s_{n}\}=\{s_{1},\ldots,s_{n-1}\}$.
The root system of type $E_{8}$ may be regarded as a subset of ${\mathbb{R}}^{8}$ as follows. There are 112 roots of the form $\pm 2(\varepsilon_{i}\pm\varepsilon_{j})$ where $1\leq i\neq j\leq 8$, and there are 128 roots of the form $\sum_{i=1}^{8}\pm\varepsilon_{i}$, where the signs are chosen so that the total number of - is even. The simple roots are $\Pi=\{\alpha_{1},\alpha_{2},\ldots,\alpha_{8}\}$, where
$\alpha_{1}=\varepsilon_{1}-\varepsilon_{2}-\varepsilon_{3}-\varepsilon_{4}-\varepsilon_{5}-\varepsilon_{6}-\varepsilon_{7}+\varepsilon_{8},$
$\alpha_{2}=2(\varepsilon_{1}+\varepsilon_{2})$, and $\alpha_{i}=2(\varepsilon_{i-1}-\varepsilon_{i-2})$ for all $3\leq i\leq 8$. If $k$ is the largest integer such that $\varepsilon_{k}$ appears in $\alpha$ with nonzero coefficient $c$, then $\alpha$ is positive if and only if $c>0$. The highest root, $\theta_{8}$, is $2(\varepsilon_{7}+\varepsilon_{8})$, and it decomposes as the sum
(2.4) $\theta_{8}=2\alpha_{1}+3\alpha_{2}+4\alpha_{3}+6\alpha_{4}+5\alpha_{5}+4\alpha_{6}+3\alpha_{7}+2\alpha_{8}.$
The bilinear form $B$ is $1/4$ of the usual dot product. The highest root $\theta_{8}$ has the property that $B(\theta_{8},\alpha_{8})=1$ and $B(\theta_{8},\alpha_{i})=0$ for all $1\leq i\leq 7$.
If $\Phi=\Phi(E_{8})$ is the root system of type $E_{8}$ constructed above, then the root system $\Phi(E_{7})$ of type $E_{7}$ may be obtained from $\Phi$ as a subsystem, namely, as the set $\Phi\cap{\mathbb{Z}}(\Pi\setminus\{\alpha_{8}\})$; the highest root in $\Phi(E_{7})$ is $2(\varepsilon_{8}-\varepsilon_{7})$, and it decomposes as the sum
(2.5) $\theta_{7}=2\alpha_{1}+2\alpha_{2}+3\alpha_{3}+4\alpha_{4}+3\alpha_{5}+2\alpha_{6}+\alpha_{7}.$
Similarly, the root system $\Phi(E_{6})$ of type $E_{6}$ can be identified as the set $\Phi\cap{\mathbb{Z}}(\Pi\setminus\{\alpha_{8},\alpha_{7}\})$.
Definition 2.1.
Let $\Phi$ be any root system with Dynkin diagram and let $\Pi=\{\alpha_{i}:i\in\Gamma\}$ be the simple roots of $\Phi$. For each root $\alpha=\sum_{i\in\Gamma}c_{i}\alpha_{i}$ and each $i\in\Gamma$, we define the $i$-height, denoted $\text{\rm ht}_{i}(\alpha)$, to be the coefficient $c_{i}$. For all $i\in\Gamma$ and $j\in{\mathbb{Z}}_{\geq 0}$, we define
$\Phi_{{i},{j}}=\{\alpha\in\Phi_{+}:\text{\rm ht}_{i}(\alpha)=j\}.$
2.2. Orthogonal sets of a fixed size
We introduce the notion of $k$-roots and a natural Weyl group action on them in this subsection.
Definition 2.2.
Let $\Phi$ be a root system with Weyl group $W$, and let $k$ be a positive integer.
  1. (i)
    For each root $\alpha\in\Phi$, we define $\lvert\alpha\rvert$ to be the positive root in the set $\{\alpha,-\alpha\}$.
  2. (ii)
    We define a $k$-root of $\Phi$ (or of $W$) to be a set of $k$ mutually orthogonal roots in $\Phi$, and we say a $k$-root is positive if all its elements are positive. We denote the set of all $k$-roots by $\Phi^{k}$ and the set of all positive $k$-roots by $\Phi_{+}^{k}$. More generally, for any $U\subseteq\Phi$ we define $U^{k}$ and $U_{+}^{k}$ to be the sets of $k$-roots and positive $k$-roots whose elements all lie in $U$, respectively.
  3. (iii)
    We define
    $\lvert R\rvert=\{\lvert\alpha\rvert:\alpha\in R\}$
    for each $k$-root $R\in\Phi^{k}$, and we define the standard action of $W$ on the positive $k$-root to be the $W$-action given by
    (2.6) $w\cdot R=\lvert\{w\cdot\alpha:\alpha\in R\}\rvert=\{\lvert w\cdot\alpha\rvert:\alpha\in R\}$
    for all $w\in W$ and $R\in\Phi_{+}^{k}$.
Note that Equation (2.6) indeed defines a $W$-action on $\Phi_{+}^{k}$. Also note that since orthogonal sets in $R$ are linearly independent in the reflection representation $V$, $k$-roots cannot exist for $k>\dim(V)=\text{\rm rank}(\Phi)$. The notion of $k$-roots generalize the notions of $2$-roots and $n$-roots studied in [7, 8], where $n=\text{\rm rank}(\Phi)$; see [7, Remark 3.7] and [8, Section 1].
Lemma 2.3.
Let $\Phi$ be an irreducible simply laced root system of rank $n$, and let $W$ be the Weyl group of $\Phi$.
  • (i)
    There exists an $n$-root of $\Phi$ (i.e., the set $\Phi^{n}$ is nonempty) if and only $\Phi$ is of type $E_{7},E_{8}$, or $D_{n}$ for some even integer $n$. When this is the case, every set of mutually orthogonal roots of $\Phi$ can be extended to an $n$-root.
  • (ii)
    If $\Phi$ has type $E_{7},E_{8}$, or $D_{n}$ for $n$ even, then the element-wise action of $W$ on $\Phi^{n}$ given by
    $w\cdot\{\alpha_{1},\ldots,\alpha_{n}\}=\{w\cdot\alpha_{1},\ldots,w\cdot\alpha_{n}\}$
    is transitive, and so is the standard action of $W$ on the positive $n$-roots $\Phi_{+}^{n}$ given by Equation (2.6).
Remark 2.4.
  1. (i)
    By a perfect matching of the set $n]:=\{1,2,\ldots,n\}$ for an even integer $n=2k$, we mean a partition of the form $M=\{\{i_{1},j_{1}\},\ldots,\{i_{k},j_{k}\}\}$ where $\{i_{1},j_{1},\ldots,i_{k},j_{k}\}=\{1,2,\ldots,2k\}$.
  2. (ii)
    From now on, whenever we speak of “$n$-roots” of a root system $\Phi$ we will assume that $\Phi$ is a root system of rank $n$ and type $E_{7},E_{8}$, or $D_{n}$ for $n$ even. By Lemma 2.3 (i), in these cases the $n$-roots coincide with the maximal sets (with respect to set inclusion) of pairwise orthogonal roots, as well as with the bases of the reflection representation $V$ that consist of pairwise orthogonal roots. We will thus use the terms “$n$-roots”, “maximal orthogonal set”, and “orthogonal basis” interchangeably.
  3. (iii)
    The fact that any general $k$-root can be extended to an $n$-root makes $n$-roots very useful, as will be evident in Sections 4 and 5.
Proof of Lemma 2.3.
Part (i) is well known, and can be proved similarly to [8, Lemma 3.2] by an inductive argument. The first assertion of (ii) is [8, Lemma 3.2], and the second assertion then follows immediately. ∎
The following example on the $4$-roots in type $D_{4}$ turns out to be crucial for the general theory of $n$-roots.
Example 2.5.
The root system $\Phi=\Phi(D_{4})$ of type $D_{4}$ has exactly three positive 4-roots, namely, the sets
$Q_{1}=\{\varepsilon_{1}-\varepsilon_{2},\ \varepsilon_{1}+\varepsilon_{2},\ \varepsilon_{3}-\varepsilon_{4},\ \varepsilon_{3}+\varepsilon_{4}\},\quad Q_{2}=\{\varepsilon_{1}-\varepsilon_{3},\ \varepsilon_{1}+\varepsilon_{3},\ \varepsilon_{2}-\varepsilon_{4},\ \varepsilon_{2}+\varepsilon_{4}\},$
and
$Q_{3}=\{\varepsilon_{1}-\varepsilon_{4},\ \varepsilon_{1}+\varepsilon_{4},\ \varepsilon_{2}-\varepsilon_{3},\ \varepsilon_{2}+\varepsilon_{3}\}.$
They correspond to the perfect matchings $\{\{1,2\},\{3,4\}\}$, $\{\{1,3\},\{2,4\}\}$, and $\{\{1,4\},\{2,3\}\}$, respectively. The reflection $s_{\alpha}$ for the root $\alpha=\varepsilon_{2}-\varepsilon_{3}$ acts on roots as the transposition $(2,3)$ on indices, so it interchanges $Q_{1}$ and $Q_{2}$ while fixing $Q_{3}$ in the standard action of the Weyl group on $\Phi_{+}^{4}$.
Note that the 4-roots $Q_{1},Q_{2},Q_{3}$ form a partition of $\Phi_{+}$. Also note that decomposing each root as a sum of simple roots yields
$Q_{1}=\{\alpha_{1},\ \alpha_{3},\ \alpha_{4},\ \alpha_{1}+2\alpha_{2}+\alpha_{3}+\alpha_{4}\},\quad Q_{2}=\{\alpha_{1}+\alpha_{2},\ \alpha_{2}+\alpha_{3},\ \alpha_{2}+\alpha_{4},\ \alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4}\},$
and
$Q_{3}=\{\alpha_{2},\ \alpha_{1}+\alpha_{2}+\alpha_{3},\ \alpha_{1}+\alpha_{2}+\alpha_{4},\ \alpha_{2}+\alpha_{3}+\alpha_{4}\}.$
We will use these facts in Section 2.3.
2.3. Coplanar quadruples and $D_{4}$-subsystems
Let $\Phi$ be a root system of type $E_{7},E_{8}$, or $D_{n}$ for $n$ even, so that $n$-roots exist for $\Phi$, and let $W$ be the Weyl group of $\Phi$. In [8], we showed that certain quadruples of orthogonal roots called coplanar quadruples are key to understanding the $n$-roots of $\Phi$ and the action of $W$ on them. These quadruples remain useful for understanding the more general $k$-roots in this paper, so we further develop their properties in this subsection.
Definition 2.6.
Let $\Phi$ be a root system of type $E_{7},E_{8}$, or $D_{n}$ for $n$ even. We define a set of four distinct positive roots $Q=\{\beta_{1},\beta_{2},\beta_{3},\beta_{4}\}\subseteq\Phi$ to be a coplanar quadruple if $\beta_{1}+\beta_{2}+\beta_{3}+\beta_{4}=2\beta$ for some root $\beta\in\Phi$.
Example 2.7.
The 4-roots $Q_{1},Q_{2}$ and $Q_{3}$ of type $D_{4}$ from Example 2.5 all sum to twice a root, namely, twice the roots $\beta_{1}=\varepsilon_{1}+\varepsilon_{3},\beta_{2}=\varepsilon_{1}+\varepsilon_{2}$, and $\beta_{3}=\varepsilon_{1}+\varepsilon_{2}$, respectively, so they are all coplanar quadruples.
Although every orthogonal quadruple in type $D_{4}$ is coplanar, this is not true in general. For example, the four orthogonal simple roots $\{\alpha_{1},\alpha_{3},\alpha_{5},\alpha_{6}\}$ in type $D_{6}$ do not form a coplanar quadruple.
Coplanar quadruples are intimately connected to $D_{4}$-subsystems of $\Phi$. To make this precise, it is helpful to have the following definition.
Definition 2.8.
Let $\Phi$ be a root system of type $E_{7},E_{8}$, or $D_{n}$ for $n$ even. Let $R\subset\Phi$ be a set of mutually orthogonal positive roots of $\Phi$. For each integer $4\leq k\leq|R|$, we define $D_{k}(R)$ to be the set
[…]
We define $\mathcal{Q}(R)$ to be the set of all $D_{4}$-subsystems $\Psi$ of $\Phi$ such that $\Psi\cap R\in D_{4}(R)$.
The following proposition expands [8, Proposition 3.3] and details many useful properties of coplanar quadruples.
Proposition 2.9.
Let $\Phi$ be a root system of type $E_{7}$, $E_{8}$, or $D_{n}$ for $n\geq 4$ even, let $V$ be the reflection representation of the Weyl group of $\Phi$, and let $R\in\Phi_{+}^{n}$ be an orthogonal basis of $V$.
  • (i)
    If $\beta\in{\Phi_{+}}\setminus R$, then $\beta$ is orthogonal to all but four elements $Q=\{\beta_{1},\beta_{2},\beta_{3},\beta_{4}\}$ of $R$. Moreover, in this case we have $\beta=(\pm\beta_{1}\pm\beta_{2}\pm\beta_{3}\pm\beta_{4})/2$ for suitable choices of sign; the set
    (2.7) $\Psi_{Q}:=Q\cup(-Q)\cup\{(\pm\beta_{1}\pm\beta_{2}\pm\beta_{3}\pm\beta_{4})/2\}$
    is a subset of $\Phi$ that forms a subsystem of type $D_{4}$; and $\Psi_{Q}$ is the smallest subsystem of $\Phi$ containing both $Q$ and $\beta$.
  • (ii)
    The following are equivalent for a subset $Q\subseteq R$:
    • (a)
      $Q$ is a coplanar quadruple, i.e., we have $\sum_{\beta\in Q}\beta=2\alpha$ for some root $\alpha$;
    • (b)
      $Q$ is the support of some root $\alpha\in\Phi\backslash(R\cup-R)\subseteq V$ with respect to the basis $R$;
    • (c)
      $Q\in D_{4}(R)$;
    • (d)
      $Q=\Psi\cap R$ for some $\Psi\in\mathcal{Q}(R)$.
  • (iii)
    The maps
    $\phi:D_{4}(R)\rightarrow\mathcal{Q}(R),\ Q\mapsto\text{\rm Span}(Q)\cap R$
    and
    $\psi:\mathcal{Q}(R)\rightarrow D_{4}(R),\ \Psi\mapsto\Psi\cap R$
    are mutually inverse maps. Every root $\alpha\in\Psi\setminus(R\cup-R)$ lies in $\text{\rm Span}(Q)$ for a unique element $Q\in D_{4}(R)$, namely, the support of $\alpha$.
Remark 2.10.
For each maximal orthogonal set of roots $R$, part (ii) of Proposition 2.9 shows that the elements of $D_{4}(R)$ are precisely the coplanar quadruples in $R$, and part (iii) establishes natural bijections between the coplanar quadruples and the $D_{4}$-subsystems of $\Phi$ intersecting $R$ in an element of $D_{4}(R)$. Part (ii) also shows that each positive root $\alpha$ not in $R$ naturally gives rise to a coplanar quadruple in $R$, namely, the support of $\alpha$ with respect to $R$.
Proof of Proposition 2.9.
Part (i) follows from [8, Proposition 3.8] and the first paragraph of its proof.
To prove (ii), first note that (a) immediately implies (b). Suppose that $\alpha$ and $Q$ satisfy the conditions of (b). Since $R$ is orthogonal, for each root $\gamma=\sum_{\beta_{i}\in R}c_{i}\beta_{i}\in\Phi$ we have $c_{i}=B(\gamma,\beta_{i})/2$, so the support of $\gamma$ with respect to $R$ consists precisely of the elements in $R$ not orthogonal to $R$. It follows from (i) that $Q$ is a quadruple $Q=\{\beta_{1},\beta_{2},\beta_{3},\beta_{4}\}$, and that the $D_{4}$-subsystem $\Psi_{Q}$ defined in Equation (2.7) is contained in $\text{\rm Span}(Q)\cap\Phi$. Conversely, for any root $\gamma=c_{1}\beta_{1}+c_{2}\beta_{2}+c_{3}\beta_{3}+c_{4}\beta_{4}\in\text{\rm Span}(Q)$, we have
$2=B(\gamma,\gamma)=2c_{1}^{2}+2c_{2}^{2}+2c_{3}^{2}+2c_{4}^{2}$
and
$c_{i}=B(\gamma,\beta_{i})/2\in\{0,\pm 1,\pm(1/2)\}$
for each $i$ by Section 2.2, so either we have $c_{i}=1$ and $\gamma=\pm\beta_{i}$ for some $i$, or we have $c_{i}=\pm(1/2)$ for all $i$. It follows that $\text{\rm Span}(Q)\cap\Phi=\Psi_{Q}$, which proves (c). The same argument also shows that if a root $\gamma$ not in $R$ lies in $\text{\rm Span}(Q)$ for a quadruple $Q\subseteq R$, then $Q$ is necessarily the support of $\gamma$, which proves the last sentence in (iii).
Now assume the hypothesis of (c), so that $\Psi=\text{\rm Span}(Q)\cap\Phi$ is a $D_{4}$-subsystem of $\Phi$. This implies that $Q\subseteq\Psi\cap R$, and that $\text{\rm Span}(\Psi)$ is a 4-dimensional subspace of $V$. Since $R$ is an orthogonal basis for $V$, we must have $|\Psi\cap R|\leq 4=|Q|$, proving (d).
Finally, assume the hypothesis of (d), so that $Q=\Psi\cap R$. Any orthogonal quadruple $Q=\{\beta_{1},\beta_{2},\beta_{3},\beta_{4}\}$ in a $D_{4}$-subsystem has the property that $\gamma=(\beta_{1}+\beta_{2}+\beta_{3}+\beta_{4})/2$ is a root, which proves (a) and completes the proof of (ii).
The map $\psi$ sends $\mathcal{Q}(R)$ to $D_{4}(R)$ since (d) implies (c) in (ii), and the proof of the fact that (c) implies (d) shows both that $\phi$ sends $D_{4}(R)$ to $\mathcal{Q}(R)$ and that $\psi\circ\phi$ is the identity map. For each $\Psi\in\mathcal{Q}(R)$ and the quadruple $Q=\Psi\cap R$, the proof of the fact that (b) implies (c) shows that any positive root $\alpha\in\Psi\setminus Q$ must satisfy $B(\alpha,\beta)=\pm 1/2$ for all $\beta$ and have $Q$ as its support, and that $\phi(Q)=\text{\rm Span}(Q)\cap\Phi=\Psi_{Q}=\Psi$, where the last equality holds by (i). It follows that $\phi\circ\psi$ is the identity map on $\mathcal{Q}(R)$, and therefore $\phi$ and $\psi$ are mutual inverses. ∎
As in [8], it will be useful in this paper to further categorize coplanar quadruples into three mutually exclusive types. The definition of these types come from the combinatorics of perfect matchings.
Definition 2.11.
For any two blocks $\{a,b\}$ and $\{c,d\}$ in a perfect matching $M$ such that $a<b$ and $c<d$, we call the pair $P=\{\{a,b\},\{c,d\}\}$ an alignment if $a<b<c<d$, a crossing if $a<c<b<d$, and a nesting if $a<c<d<b$; see Figure 3. We also call each alignment, crossing, or nesting a feature.
It follows from Definition 2.11 that any pair of blocks in a perfect matching forms exactly one type of feature. For example, the matchings $\{\{1,2\},\{3,4\}\}$, $\{\{1,3\},\{2,4\}\}$, and $\{\{1,4\},\{2,3\}\}$ corresponding to the coplanar quadruples $Q_{1},Q_{2},Q_{3}$ in the root system of type $D_{4}$ from Example 2.7 form an alignment, crossing, and nesting, respectively. Note that by Example 2.5, the quadruples $Q_{1},Q_{2}$ and $Q_{3}$ can be distinguished by the numbers of simple roots in them, because they contain $3,0,$ and $1$ simple root(s), respectively. This motivates the following definition, which extends the notions of alignments, crossings, and nestings to the context of orthogonal roots.
aabbccdd
(a) alignment
aaccbbdd
(b) crossing
aaccddbb
(c) nesting
Figure 3. Features formed by two blocks $\{a,b\},\{c,d\}$ in a perfect matching
Definition 2.12.
Let $Q$ be a coplanar quadruple in $\Phi$ and let $\Psi=\text{\rm Span}(Q)\cap\Phi$ be the corresponding subsytem of type $D_{4}$. Let $\Pi_{\Psi}$ be the induced simple system of $\Psi$. We call $Q$ an alignment if $\lvert Q\cap\Pi_{\Psi}\rvert=3$, a crossing if $\lvert Q\cap\Pi_{\Psi}\rvert=0$, and a nesting if $\lvert Q\cap\Pi_{\Psi}\rvert=1$. We also call each alignment, crossing, or nesting a feature.
Remark 2.13.
Definition 2.12 resembles but differs from [8, Definition 3.9], where the classification of coplanar quadruples into the three types of possible features is characterized in a less concise way. However, the definitions are equivalent: in the notation of Example 2.7, each coplanar quadruple is necessarily one of $Q_{1},Q_{2}$ and $Q_{3}$, and both definitions classify $Q_{1}$ as an alignment, $Q_{2}$ as a crossing, and $Q_{3}$ as an alignment. In other words, both definitions guarantee that each coplanar quadruple $Q$ is an alignment, crossing, or nesting if and only if the perfect matching corresponding to it is a feature of the same type under the usual coordinates for the $D_{4}$-subsystem $\Psi=\text{\rm Span}(Q)\cap\Phi$ associated to $Q$.
[…]
Lemma 2.14.
Let $\Phi$ be a root system of type $E_{7}$, $E_{8}$, or $D_{n}$ for $n$ even. Let $R\in\Phi_{+}^{n}$ be a positive $n$-root, let $Q$ and $Q^{\prime}$ be features in $R$ such that $Q\neq Q^{\prime}$ and $Q\cap Q^{\prime}\neq\emptyset$, and define $H=Q\cup Q^{\prime}$.
  • (i)
    We have $|H|=6$, and $\Xi:=\text{\rm Span}(H)\cap\Phi$ is a subsystem of type $D_{6}.$ In other words, we have $H\in D_{6}(R)$.
  • (ii)
    The sets $Q$, $^{\prime}, QQ$$\cap Q^{\prime}$, and $Q\cup Q^{\prime}$ all consist of couples in $\Xi$.
  • (iii)
    If $\Phi$ has type $E_{8}$, then $R\setminus Q$ is also a coplanar quadruple.
Proof.
Part (i) and (ii) a restatement of [8, Proposition 3.20]. Part (iii) follows from [8, Lemma 3.18], which states that the features of $R$ form the blocks of a Steiner quadruple system $S(3,4,8)$, meaning that every $3$-element subset of $R$ is contained in a unique feature. More precisely, there is a unique Steiner quadruple system on $8$ points ([8, Remark 3.16]) and its blocks are closed under taking complements, so (iii) follows. ∎
3. Residues and quasiparabolic sets
In Section 3.1, we establish the key new definitions of residues, Rothe diagrams, the level statistic, as well as generalized Hafnians and Pfaffians and their $q$-analogues. We also recall the notion of a quasiparabolic set, due to Rains and Vazirani. Then, in Section 3.2, we give details of two of the motivating examples of residues: one based on perfect matchings of a set of size $2k$ (Theorem 3.10), and one based on permutations of $k$ objects (Theorem 3.13).
3.1. Key definitions
The following definition introduces the key notion of residues, which applies to all root systems. This notion is new to the best of our knowledge.
Definition 3.1.
Let $\Phi$ be an arbitrary root system, and let $U\subseteq\Phi_{+}$.
  1. (i)
    We define $\Omega_{U}$ to be the set of all $\kappa$-roots of $U$, where $\kappa$ is the maximum cardinality of a pairwise orthogonal subset of $U$.
  2. (ii)
    For any orthogonal set $R\in\Omega_{U}$, we define the set of residues of $R$ with respect to $U$ to be
    $\text{\rm Res}_{U}(R)=\{\gamma\in U:s_{\beta}(\gamma)>0\text{\ for\ all\ }\beta\in R\},$
    and we will also refer to $\text{\rm Res}_{U}(R)$ as the (generalized) Rothe diagram of $R$ with respect to $U$. We define $\rho_{U}(R):=\lvert\text{\rm Res}_{U}(R)\rvert$ to be the level of $R$ with respect to $U$.
  3. (iii)
    We identify an element $R=\{\beta_{1},\beta_{2},\ldots,\beta_{k}\}\in\Omega_{U}$ with the monomial $\prod_{\beta\in R}x_{\beta}$, where $\{x_{\beta}:\beta\in U\}$ is a set of commuting indeterminates. The (generalized) quantum Hafnian of $U$, $\text{\rm QHf}(U)=\text{\rm QHf}(U,q)$, is the element of ${\mathbb{Z}}[U][q]$ given by
    $\sum_{R\in\Omega_{U}}q^{\rho(R)}\prod_{\beta_{i}\in R}x_{\beta_{i}}.$
    The specializations $\text{\rm QPf}(U):=\text{\rm QHf}(U,-q)$, $\text{\rm Hf}(U):=\text{\rm QHf}(U,1)$, and $\text{\rm Pf}(U):=\text{\rm QHf}(U,-1)$ are called the (generalized) quantum Pfaffian, the (generalized) Hafnian, and the (generalized) Pfaffian of $U$, respectively.
  4. (iv)
    We define the Poincaré polynomial of $U$ to be the polynomial
    $PS_{U}(q):=\sum_{R\in\Omega_{U}}q^{\rho(R)}\in{\mathbb{Z}}[q]$
    obtained from $\text{\rm QHf}(U)$ by specializing all the indeterminates $x_{\beta}$ to $1$. If $PS_{U}(q)$ has the form $\prod_{d\in D}[d]_{q}$ for some multiset $D$ of elements from $\{$$\mathbb{N}\backslash\{0,1\}$, where $d]_{q}$ is the quantum integer
    $d]_{q}=\frac{q^{d}-1}{q-1}=1+q+q^{2}+\ldots+q^{d-1},$
    then we call the elements of $D$ the degrees of $U$.
Remark 3.2.
  1. (i)
    Since $s_{\gamma}(\gamma)=-\gamma<0$ for any positive root $\gamma$ in a root system, it follows that $\text{\rm Res}_{U}(R)$ cannot contain any element of $R$ in the setting of Definition 3.1.
  2. (ii)
    Factorization of a polynomial in ${\mathbb{Z}}[q]$ into quantum integers is unique up to order when possible, so the multiset $D$ in Definition 3.1 (iv) is well-defined.
Example 3.3.
Let $\Phi$ be the root system of type $D_{4}$, whose highest root is $\theta=\alpha_{1}+2\alpha_{2}+\alpha_{3}+\alpha_{4}$, and define $U={\Phi_{+}}$. There are three maximal orthogonal subsets of positive roots in $U$, namely,
$Q_{1}$ $=\{\alpha_{1},\ \alpha_{3},\ \alpha_{4},\ \alpha_{1}+2\alpha_{2}+\alpha_{3}+\alpha_{4}\},$
$Q_{2}$ $=\{\alpha_{1}+\alpha_{2},\ \alpha_{2}+\alpha_{3},\ \alpha_{2}+\alpha_{4},\ \alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4}\},\quad\text{and}$
$Q_{3}$ $=\{\alpha_{2},\ \alpha_{1}+\alpha_{2}+\alpha_{3},\ \alpha_{1}+\alpha_{2}+\alpha_{4},\ \alpha_{2}+\alpha_{3}+\alpha_{4}\}.$
Direct computation shows that
$\text{\rm Res}_{U}(Q_{1})=\emptyset,\quad\text{\rm Res}_{U}(Q_{2})=\{\theta\}=\{\varepsilon_{1}+\varepsilon_{2}\},\quad\text{and}\quad\text{\rm Res}_{U}(Q_{3})=\{\theta-\alpha_{2},\ \theta\}=\{\varepsilon_{1}+\varepsilon_{2},\ \varepsilon_{1}+\varepsilon_{3}\},$
from which it follows that $\rho_{U}(Q_{1})=0$, $\rho_{U}(Q_{2})=1$, and $\rho_{U}(Q_{3})=2$.
Using the coordinate system for roots of type $D_{4}$, and writing $u_{ij}$ as shorthand for $x_{\varepsilon_{i}+\varepsilon_{j}}x_{\varepsilon_{i}-\varepsilon_{j}}$, we have
$\text{\rm QHf}(U)$ $=u_{12}u_{34}+qu_{13}u_{24}+q^{2}u_{14}u_{23},$
$\text{\rm QPf}(U)$ $=u_{12}u_{34}-qu_{13}u_{24}+q^{2}u_{14}u_{23},$
$\text{\rm Hf}(U)$ $=u_{12}u_{34}+u_{13}u_{24}+u_{14}u_{23},\quad\text{\rm and}$
$\text{\rm Pf}(U)$ $=u_{12}u_{34}-u_{13}u_{24}+u_{14}u_{23}.$
It follows that $PS_{\Phi_{+}}(q)=1+q+q^{2}$, so that $3$ is the only degree of $U$.
Note that the sets $Q_{i}$ for $1\leq i\leq 3$ are exactly the three coplanar quadruples partitioning $\Phi_{+}$} from Example 2.7, with $Q_{1}$ being an alignment, $Q_{2}$ a crossing, and $Q_{3}$ a nesting. The highest root $\theta$ and the second highest root $\theta-\alpha_{2}$ are the only roots that may appear in $\text{\rm Res}_{U}(R)$ for any $R\in\Omega_{U}$. Note also that if $\alpha$ is a residue and $\alpha\in\{\theta,\theta-\alpha_{2}\}$, then $\alpha$ has full support with respect to any of the orthogonal bases $Q_{i}=\{\beta_{1},\beta_{2},\beta_{3},\beta_{4}\}$.
The next definition considers an important special case of Definition 3.1 (ii). Note that by Section 2.1, if $R$ is a $k$-root of a root system $\Phi$, then the set $\Psi:=\text{\rm Span}(R)\cap\Phi$ is a subsystem of $\Phi$, and the positive roots $\Phi_{+}$ of $\Phi$ induce a positive system $\Psi_{+}=\Psi\cap\Phi_{+}=\text{\rm Span}(R)\cap\Phi_{+}$ of $\Psi$.
Definition 3.4.
Let $\Phi$ be an arbitrary root system, let $R\subseteq\Phi_{+}$ be a $k$-root of $\Phi$, and let $\Psi_{R}=\text{\rm Span}(R)\cap\Phi$ be the subsystem of $\Phi$ induced by $R$. We define the set of residues of $R$ to be the set
$\text{\rm Res}(R)=\text{\rm Res}_{\Psi_{+}}(R),$
where $\Psi_{+}=\text{\rm Span}(R)\cap\Phi_{+}$ is the positive system of $\Psi$ induced by $\Phi_{+}$. We refer to $\text{\rm Res}(R)$ as the (generalized) Rothe diagram of $R$, and we define $\rho(R):=\lvert\text{\rm Res}(R)\rvert$ to be the level of $R$.
Remark 3.5.
  1. (i)
    It follows from Definitions 3.1 (ii) and 3.4 that for any $k$-root $R$ of a root system $\Phi$, we have $\text{\rm Res}(R)=\text{\rm Res}_{\Phi_{+}}(R)\cap\text{\rm Span}(R)$. In particular, if $R$ is an $n$-root where $n=\text{\rm rank}(\Phi)$, then $\text{\rm Res}(R)=\text{\rm Res}_{\Phi_{+}}(R)$ and $\rho(R)=\rho_{\Phi_{+}}(R)$.
  2. (ii)
    In the setting of Example 3.3, we have $\text{\rm Res}(Q_{i})=\text{\rm Res}_{\Phi_{+}}(Q_{i})=\text{\rm Res}_{U}(Q_{i})$ and $\rho(Q_{i})=\rho_{\Phi_{+}}(Q_{i})=\rho_{U}(Q_{i})$ for each $i$. Since the calculation of the residue sets in this example depends only on the choice of a simple system for a type $D_{4}$ root system, it follows that whenever $Q$ is a feature in a root system of type $E_{7},E_{8}$ or $D_{n}$ for $n$ even, we have $\rho(Q)=0$ if $Q$ is an alignment, $\rho(Q)=1$ if $Q$ is a crossing, and $\rho(Q)=2$ if $Q$ is a nesting.
Quasiparabolic sets were introduced by Rains and Vazirani for a general Coxeter system as follows.
Definition 3.6.
[14, Section 2, Section 5] Let $(W,S)$ be a Coxeter system with set of reflections $T$. A scaled $W$-set is a pair $(X,\lambda)$ where $X$ is a $W$-set and $\lambda:X\rightarrow{\mathbb{Z}}$ is a function satisfying $|\lambda(sx)-\lambda(x)|\leq 1$ for all $s\in S$. An element $x\in X$ is $W$-minimal if $\lambda(sx)\geq\lambda(x)$ for all $s\in S$ and is $W$-maximal if $\lambda(sx)\leq\lambda(x)$ for all $s\in S$.
A quasiparabolic set for $W$, or quasiparabolic $W$-set, is a scaled $W$-set $X$ satisfying the following two properties:
  • (QP1)
    for any $r\in T$ and $x\in X$, if $\lambda(rx)=\lambda(x)$, then $rx=x$;
  • (QP2)
    for any $r\in T$, $x\in X$, and $s\in S$, if $\lambda(rx)>\lambda(x)$ and $\lambda(srx)<\lambda(sx)$, then $rx=sx$.
For a quasiparabolic set $X$, we define the quasiparabolic order on $X$, $\leq_{Q}$, to be the weakest partial order such that $x\leq_{Q}rx$ whenever $x\in X$, $r\in T$, and $\lambda(x)\leq\lambda(rx)$.
Rains and Vazirani call $\lambda(x)$ the height of $x$ and $\leq_{Q}$ the Bruhat order, but we will refer to them as the level of $x$ and the quasiparabolic order because of the potential for confusion in the context of this paper.
For any quasiparabolic $W$-set $(X,\lambda)$, it follows from [14, Corollary 2.10] that each $W$-orbit in $X$ contains at most one $W$-minimal element, so if $X$ is a finite and transitive $W$-set then $X$ has a unique $W$-minimal element, $x_{0}$. Moreover, when this is the case, the level of any element $x\in X$ equals the minimum length of the elements $w\in W$ such that $w(x_{0})=x$ by [14, Corollary 2.13].
We may associate a quantum Hafnian and a Poincaré polynomial to a general quasiparabolic set by defining $\text{\rm QHf}(X)=\sum_{x\in X}q^{\rho(R)}x$ and $PS_{X}(q)=\sum_{x\in X}q^{\rho(x)}.$ However, we will not require this level of generality, because our focus for the rest of the paper will be on quasiparabolic sets that arise from sets of orthogonal roots via the following construction.
Definition 3.7.
Let $W$ be a finite Weyl group with root system $\Phi$ and generating set $S$, let $I\subseteq S$, and let $U\subseteq{\Phi_{+}}$. We say that $(W,I,U)$ is a quasiparabolic triple, or QP-triple, if the set $\Omega_{U}$ is a quasiparabolic set for $W_{I}$ under the restriction of the standard action (Definition 2.2) and level function $\lambda(R):=\rho_{U}(R)$. A QP-triple $(W,I,U)$ is transitive if $\Omega_{U}$ is a transitive quasiparabolic $W_{I}$-set.
3.2. Motivating examples
We now present the two motivating examples of residues: one based on perfect matchings, and the other based on permutations. The following definition will be helpful in the context of the perfect matching example.
Definition 3.8.
Let $M=\{\{i_{1},j_{1}\},\ldots,\{i_{k},j_{k}\}\}$ be a perfect matching of the set $[2k]$, where we have $k\geq 2$ and $i_{1}<\ldots<i_{k}$ and $i_{t}<j_{t}$ for all $1\leq t\leq k$. Following [12, §2], we define $\sigma_{M}\in S_{2k}$ to be the permutation whose two-line notation is given by
$\sigma_{M}=\begin{bmatrix}1&2&\ldots&2k-1&2k\\ i_{1}&j_{1}&\ldots&i_{k}&j_{k}\end{bmatrix}$.
We define $\tau_{M}\in S_{2k}$ to be the fixed point free involution whose disjoint cycle notation is given by
$\tau_{M}=(i_{1},j_{1})(i_{2},j_{2})\ldots(i_{k},j_{k}).$
Let
$I_{M}=\{(i,j):1\leq i<j\leq 2k\text{\ and\ }\sigma_{M}(i)>\sigma_{M}(j)\}$
be the set of inversions of $\sigma_{M}$, and define
$D_{M}=\{(i,j):1\leq i<j\leq 2k\text{\ and\ }\tau_{M}(i)>j\text{\ and\ }\tau_{M}(j)>i\}.$
Lemma 3.9.
Maintain the notation of Definition 3.8. There is a bijection $\psi_{M}:I_{M}\rightarrow D_{M}$ given by
$\psi_{M}((i,j))=(\tau_{M}\sigma_{M}(i),\sigma_{M}(j)).$
Proof.
For any integer $1\leq e\leq k$, the definitions of $\sigma_{M}$ and $I_{M}$ imply that $(2e-1,2e)\notin I_{M}$, and the definitions of $\tau_{M}$ and $D_{M}$ imply that $(\sigma_{M}(2e-1),\sigma_{M}(2e))\notin D_{M}$. It follows that every pair $(i,j)\in I_{M}$ corresponds to two distinct integers $1\leq e<f\leq k$ such that $i\in\{2e-1,2e\}$ and $j\in\{2f-1,2f\}$, and every pair $(i,j)\in I_{M}$ consists of elements from two blocks $\{\sigma_{M}(2e-1),\sigma_{M}(2e)\}$ and $\{\sigma_{M}(2f-1),\sigma_{M}(2f)\}$ for such two distinct integers $1\leq e<f\leq k$.
Suppose we have $1\leq e<f\leq k$, and let $a:=\sigma_{M}(2e-1)$, $b:=\sigma_{M}(2e)$, $c:=\sigma_{M}(2f-1)$ and $d:=\sigma_{M}(2f)$. It follows from the definition of $\sigma_{M}$ that we have $a<c<d$ and $a<b$. There are three possibilities for the relative order of $a,b,c,$ and $d$: $a<c<d<b$, $a<c<b<d$, and $a<b<c<d$. These correspond respectively to nestings, crossings, and alignments of the matching $M$.
Of the three cases just described, the first case gives two inversions of $\sigma_{M}$, namely (2e,2f1)(2e,2f-1) and $(2e,2f)$, and two elements of $D_{M}$, namely $(a,d)$ and $(a,c)$. The second case gives one inversion of $\sigma_{M}$, namely $(2e,2f)$, and one element of $D_{M}$, namely $(a,c)$. The third case does not give rise to any inversions of $\sigma_{M}$ or any elements of $D_{M}$. The map $\psi_{M}$ matches the inversions with the elements of $D_{M}$ in each case, which completes the proof. ∎
Theorem 3.10.
[…]
$U=\{\varepsilon_{i}+\varepsilon_{j}:1\leq i<j\leq 2k\},$
and write $u_{ij}=u_{ji}:=\varepsilon_{i}+\varepsilon_{j}$ for all $1\leq i<j\leq 2k$. Maintain the notation of Definition 3.8, and let $\mathcal{M}$ be the set of perfect matchings of $[2n]$.
  • (i)
    The triple $(W,I,U)$ is a QP-triple.
  • (ii)
    The set $\Omega_{U}$ has (2k1)!!(2k-1)!! elements, each of which has the form
    $R(M)=\{u_{i_{1}j_{1}},u_{i_{2}j_{2}},\ldots,u_{i_{k}j_{k}}\}$
    for some perfect matching $M\in\mathcal{M}$.
  • (iii)
    For each $M\in\mathcal{M}$, we have
    $\text{\rm Res}_{U}(R(M))=\{u_{ij}:1\leq i<j\leq 2k,\tau_{M}(i)>j,\text{and }\tau_{M}(j)>i\}.$
  • (iv)
    For each $M\in\mathcal{M}$, the set
    $\{(i,j):j<i\text{\ and\ }u_{ji}\in\text{\rm Res}_{U}(R(M))\}=\{(i,j):(j,i)\in D_{M}\}$
    is the Rothe diagram $\hat{D}_{\tt FPF}(\tau_{M})$ of the fixed point free involution $\tau_{M}$ in the sense of Hamaker–Marberg–Pawlowski [9, (3.3)].
  • (v)
    For each $M\in\mathcal{M}$, we have
    $\rho_{U}(R(M))=\ell(\sigma_{M})=(\ell(\tau_{M})-k)/2,$
    where $\ell$ denotes the length of a permutation.
  • (vi)
    The generalized quantum Hafnian $\text{\rm QHf}(U)$ is the $q$-Hafnian in the sense of Jing–Zhang [13, Proposition 3.6], and the generalized quantum Pfaffian $\text{\rm QPf}(U)$ is the $q$-Pfaffian in the sense of Strickland [16] and Jing–Zhang [13, Proposition 3.4].
  • (vii)
    The Hafnian $\text{\rm Hf}(U)$ is the Hafnian of a $2k\times 2k$ symmetric matrix $(a_{ij})$ satisfying $a_{ij}=u_{ij}$ for all $1\leq i<j\leq 2k$, and the Pfaffian $\text{\rm Pf}(U)$ is the Pfaffian of a $2k\times 2k$ skew-symmetric matrix $(a_{ij})$ satisfying $a_{ij}=u_{ij}$ for all $1\leq i<j\leq 2k$.
  • (viii)
    The Poincaré polynomial of $U$ is $\prod_{i=2}^{k}[2i-1]_{q}$, and its degrees are $3,5,7,\ldots,2k-1$.
Proof.
Part (ii) follows from the observation that two elements of $U$ are orthogonal if and only if they have disjoint support with respect to the $\varepsilon$-basis.
If $\beta,u_{ij}\in U$ satisfy $s_{\beta}(u_{ij})<0$, then we must either have $\beta=u_{aj}$ for some $a\leq i$, or $\beta=u_{ib}$ for some $b\leq j$. It follows that $u_{ij}$ is a residue of $R(M)$ if and only if we have both $\tau_{M}(i)>j$ and $\tau_{M}(j)>i$, proving (iii).
The set
$\{(j,i):j<i,\tau(i)>j\text{\ and\ }\tau(j)>i\}$
is the Rothe diagram $\hat{D}_{\tt FPF}(\tau)$ of the fixed point free involution $\tau$ as defined by Hamaker–Marberg–Pawlowski [9, (3.3)]. Part (iv) follows by applying (iii) to this definition.
It follows from [9, Proposition 3.6] that the level, $\rho_{U}(R(M))=|\hat{D}_{\tt FPF}(\tau_{M})|$, of $R$ is equal to $(\ell(\tau_{M})-k)/2$. Lemma 3.9 (iii) implies that $\rho_{U}(R(M))=\ell(\sigma_{M}$), and this completes the proof of (v).
Rains–Vazirani [14, §4] proved that the function $(\ell(\tau_{M})-k)/2$ is a quasiparabolic level function for the set of fixed point free involutions of $S_{2k}$, regarded as an $S_{2k}$-set in the natural way. Combining this result with part (v) proves (i).
It follows from (v) that we have
$\text{\rm QHf}(U)=\sum_{M}q^{\ell(\sigma_{M})}u_{i_{1}j_{1}}u_{i_{2}j_{2}}\ldots u_{i_{k}j_{k}},$
where the sum is over all perfect matchings of $[2k]$. This agrees with Jing–Zhang’s construction of the quantum Hafnian from [13, Proposition 3.6] All the assertions of (vi) and (vii) now follow from the relevant definitions.
Part (viii) is a standard result that has been rediscovered several times; see [8, Proposition 7.1]. ∎
Example 3.11.
In the case $n=6$ of Theorem 3.10, we have
$\text{\rm QHf}(U)=$ $\hphantom{+}u_{12}u_{34}u_{56}+qu_{13}u_{24}u_{56}+qu_{12}u_{35}u_{46}+q^{2}u_{13}u_{25}u_{46}+q^{2}u_{14}u_{23}u_{56}$
$+q^{2}u_{12}u_{36}u_{45}+q^{3}u_{14}u_{25}u_{36}+q^{3}u_{15}u_{23}u_{46}+q^{3}u_{13}u_{45}u_{26}+q^{4}u_{16}u_{23}u_{45}$
$+q^{4}u_{15}u_{24}u_{36}+q^{4}u_{14}u_{35}u_{26}+q^{5}u_{15}u_{26}u_{34}+q^{5}u_{16}u_{24}u_{35}+q^{6}u_{16}u_{25}u_{34}.$
Remark 3.12.
In certain cases, the product of the residues of a perfect matching, when regarded as a polynomial in the $\varepsilon_{i}$, has an interpretation in terms of Schubert polynomials; see [9, Theorem 1.3].
Theorem 3.13.
[…]
$U=\{\varepsilon_{i}+\varepsilon_{j+k}:i,j\in[k]\},$
write $v_{ij}:=\varepsilon_{i}+\varepsilon_{j+k}$ for all $1\leq i,j\leq k$.
  • (i)
    The triple $(W,I,U)$ is a QP-triple.
  • (ii)
    The set $\Omega_{U}$ has $k!$ elements, each of which has the form
    $R(\pi):=\{v_{1\pi(1)},v_{2\pi(2)},\ldots,v_{k\pi(k)}\}$
    for some permutation $\pi\in S_{k}$.
  • (iii)
    For each $\pi\in S_{k}$, we have
    $\text{\rm Res}_{U}(R(\pi))=\{v_{ij}:i,j\in[k],\pi(i)>j,\text{and }\pi^{-1}(j)>i\}.$
  • (iv)
    The set
    $\{(i,j):v_{ij}\in\text{\rm Res}_{U}(R(\pi))\}$
    is the Rothe diagram $D(\pi)$ of the permutation $\pi$.
  • (v)
    For each permutation $\pi\in S_{k}$, we have
    $\rho_{U}(R(M))=\ell(\pi),$
    where $\ell$ denotes the length of a permutation.
  • (vi)
    The generalized quantum Hafnian $\text{\rm QHf}(U)$ is the $q$-permanent of the matrix $A=(v_{ij})$ in the sense of [12, (2.7)], and the generalized quantum Pfaffian $\text{\rm QHf}(U)$ is the quantum row permanent $\text{\rm per}_{q}(A)$ of $A$ in the sense of [13, (3.18)].
  • (vii)
    The Hafnian $\text{\rm Hf}(U)$ and the Pfaffian $\text{\rm Pf}(U)$ are the permanent $\text{\rm per}(A)$, and the determinant $\det(A)$, respectively.
  • (viii)
    The Poincaré polynomial of $U$ is $\prod_{i=2}^{k}[i]_{q}$, and its degrees are 2,3,4,,k2,3,4,...,k.
Proof.
Parts (ii) and (iii) follow by restriction from Theorem 3.10 (ii) and (iii), with each $k$-root $R(\pi)$ defined in (ii) corresponding to the perfect matching $M\{\{i,\pi(i)+k\}:i\in[k]\}$. Part (iv) then follows from the definition of the Rothe diagram $D(\pi)$ [9, (2.7)]. Part (v), which is well known, follows from the fact that there is a bijection between the inversions of $\pi$ and the Rothe diagram of $\pi$ given by $(i,j)\mapsto(i,\pi(j))$.
If we identify $\Omega_{U}$ with $S_{k}$ and $W_{I}$ with $S_{k}\times S_{k}$, then the $W_{I}$-action on $\Omega_{U}$ is given by
$(w_{1},w_{2})\cdot w=w_{1}w(w_{2})^{-1}.$
By [14, Theorem 3.1], this makes $S_{k}$ into a quasiparabolic set with height function $\ell$, proving (i).
Parts (vi) and (vii) follow from the relevant definitions, and (viii) follows from the well-known identity $\sum_{\pi\in S_{k}}q^{\ell(\pi)}=[k]_{q}!$. ∎
Remark 3.14.
  1. (i)
    It is sometimes helpful to identify $U$ as the positions on a $k\times k$ board and $\Omega_{U}$ with configurations of $k$ nonattacking rooks on the board, with each root $v_{ij}=\varepsilon_{i}\pm\varepsilon_{j+k}$ corresponding to the $(i,j)$-th position. A position on the board then corresponds to a residue if and only if it is vacant and not attacked either from the left or from above. Plotting a $k$-root and its residues on the board yields the original Rothe diagram of the permutation correspoding to the rook configuration; see Figure 1.
  2. (ii)
    Instead of regarding $\Omega_{U}$ as an $S_{k}\times S_{k}$-set, we may regard it as an $(S_{k},S_{k})$-biset, as in the proof of part (v) of the theorem, and from this perspective, $\Omega_{U}$ has a left $S_{k}$-action permuting the rows, and a right $S_{k}$-action permuting the columns. The actions commute with each other, and $\Omega_{U}$ is a quasiparaboic $S_{k}$-set under each action.
4. Further examples of QP-triples
In this section, we give details of many other examples of QP-triples. Section 4.1 studies the root systems of rank $n$ where $n$-roots exist, and we show that for these root systems we can form a QP-triple by taking $U$ to be the entire set of positive roots (Theorem 4.3). In Section 4.2, we show that it is often possible to construct examples of QP-triples by selecting those positive roots in which particular simple roots appear with odd coefficients.
4.1. Orthogonal bases of roots
Recall from Lemma 2.3 that a Weyl group $W$ of rank $n$ and type $ADE$ has a set of $n$ orthogonal roots if and only if $W$ has type $E_{7}$, $E_{8}$, or $D_{n}$ for $n$ even, and we will show that $(W,S,{\Phi_{+}})$ is a QP-triple in these cases. These examples are particularly useful because many other examples of QP-triples can be constructed from them by restriction. The next result develops some key properties of residues of $n$-roots in these cases. Recall from Remark 3.5 (i) that for each $n$-root $R$ of $\Phi$, we have $\text{\rm Res}(R)=\text{\rm Res}_{\Phi_{+}}(R)$ and $\rho(R)=\rho_{\Phi_{+}}(R)$.
Proposition 4.1.
Let $\Phi$ be a root system of type $E_{7}$, $E_{8}$, or $D_{n}$ for $n$ even, and let $R$ be a orthogonal set of $n$ positive roots in $\Phi$.
  • (i)
    The set of residues of $R$ can be expressed as the disjoint union
    $\text{\rm Res}(R)=\dot{\bigcup}_{Q\in D_{4}(R)}\text{\rm Res}(Q).$
  • (ii)
    If $\Phi$ is of type $E_{8}$, then with respect to each fixed feature $Q\in D_{4}(R)$, we have
    $\text{\rm Res}(R)=\text{\rm Res}(Q)\quad\dot{\cup}\quad\dot{\bigcup}_{Q\subset H\in D_{6}(R)}(\text{\rm Res}(H)\backslash\text{\rm Res}(Q))\quad\dot{\cup}\quad\text{\rm Res}(R\backslash Q).$
Proof.
Remark 3.2 (i) and Proposition 2.9 (ii) imply that every $\gamma\in\text{\rm Res}(R)$ lies in the support of a unique $Q\in D_{4}(R)$. It follows that $\gamma\in\text{\rm Res}(Q)$ in this case, which proves $\text{\rm Res}(R)\subseteq\dot{\bigcup}_{Q\in D_{4}(R)}\text{\rm Res}(Q)$. Conversely, any element $\gamma\in\text{\rm Res}(Q)$ is fixed by all reflections $s_{\beta}$ such that $\Q\beta\in R\backslash$ Q, which implies that $\gamma\in\text{\rm Res}(R)$ and establishes the reverse containment. The disjointness assertion follows from Proposition 2.9 (iii), which completes the proof of (i).
Suppose $\Phi$ has type $E_{8}$, and let $Q\in D_{4}(R)$. The set $R\setminus Q$ is a feature by Lemma 2.14 (iii), and it is the only element of $D_{4}(R)$ that is disjoint from $Q$ by cardinality considerations. For any $Q^{\prime}\in D_{4}(R)\setminus\{Q,R\setminus Q\}$, Lemma 2.14 (i) and (ii) imply that the set $H=Q\cup Q^{\prime}$ lies in $D_{6}(R)$ and contains three elements of $D_{4}(R)$, namely, $Q$, $Q^{\prime}$, and their symmetric difference, $Q^{\prime\prime}=Q\ \Delta\ Q^{\prime}$. Applying part (i) to the type $D_{6}$ subsystem $\Xi=\text{\rm Span}(H)\cap\Phi$ yields
$\text{\rm Res}(H)=\text{\rm Res}(Q)\ \dot{\cup}\ \text{\rm Res}(Q^{\prime})\ \dot{\cup}\ \text{\rm Res}(Q^{\prime\prime})$
and thus
$\text{\rm Res}(H)\backslash\text{\rm Res}(Q)=\text{\rm Res}(Q^{\prime})\ \dot{\cup}\ \text{\rm Res}(Q^{\prime\prime}).$
The conclusions of (ii) now follow from (i). ∎
Corollary 4.2.
Let $\Phi$ be a root system of type $E_{7},E_{8}$, or $D_{n}$ for $n$ even. If $R$ is a maximal orthogonal set of positive roots in $\Phi$, then $\rho(R)=C(R)+2N(R)$, where $C(R)$ and $N(R)$ are the numbers of crossings and nestings in $R$, respectively.
Proof.
This is immediate from Proposition 4.1 (i) and Remark 3.5 (ii). ∎
Theorem 4.3.
Let $W$ be a Weyl group of type $E_{7}$, $E_{8}$, or $D_{n}$ for $n=2k\geq 4$, with root system $\Phi$ and generating set $S$. The triple $(W,S,{\Phi_{+}})$ is a transitive QP-triple.
Proof.
By Lemma 2.3 and the paragraph preceding it, the set $\Omega_{\Phi_{+}}$ consists of the positive $n$-roots of $\Phi$, and the standard action of $W_{S}=W$ on $\Omega_{\Phi_{+}}$ is transitive. It is known [8, Theorem 4.5] that $\Omega_{\Phi_{+}}$ is a quasiparabolic set for $W$ under the standard action and level function $C+2N$, where $C$ and $N$ count the numbers of crossings and nestings in each $n$-root. Since $C(R)+2N(R)=\rho(R)=\rho_{\Phi_{+}}(R)$ for all $R\in\Omega_{\Phi_{+}}$ by Corollary 4.2, it follows that $(W,S,\Phi_{+})$ is a transitive QP-triple, which completes the proof. ∎
Corollary 4.4.
Maintain the assumptions of Theorem 4.3. Let $R$ be a positive $n$-root of $\Phi$, let $\R\alpha\in{\Phi_{+}}\backslash$ R, and let $Q$ be the support of $\alpha$ with respect to $R$. If $\rho(s_{\alpha}Q)>\rho(Q)$, then the number
$d:=(\rho(s_{\alpha}R)-\rho(R))-(\rho(s_{\alpha}Q)-\rho(Q))$
is a nonnegative even number.
Proof.
This follows from Remark 3.12 and Proposition 4.7 of [8]. By that remark and Remark 3.5 (ii), if $\rho(s_{\alpha}Q)>\rho(Q)$ then we have (a) $Q$ is an alignment and $s_{\alpha}Q$ is a crossing, or (b) $Q$ is a crossing and $s_{\alpha}Q$ is a nesting, or (c) $Q$ is an alignment and $s_{\alpha}Q$ is a nesting. In these cases, the value of $\rho(s_{\alpha}(Q))-\rho(Q)$ is 1, 1, and 2, respectively, and $d$ is a nonnegative even number by parts (ii), (iii), and (iv) of [8, Proposition 4.7], respectively. ∎
Remark 4.5.
The paper [8] details many interesting properties of the quasiparabolic sets produced by Theorem 4.3. We highlight some of them that are directly related to this paper below.
  • (i)
    In type $D_{2k}$, the quasiparabolic set $\Omega_{\Phi_{+}}$ for the Weyl group $W=W(D_{2k})$ is essentially the same as the quasiparabolic set $\Omega_{U}$ for the parabolic subgroup $W_{I}\cong S_{2k}$ from Theorem 3.10 (where $I=S(D_{2k}\setminus\{s_{2k}\}$), in the sense that there is a natural bijection from $\Omega_{U}$ to $\Omega_{\Phi_{+}}$ that is $W_{I}$-equivariant and preserves level, namely, the map
    $\phi:\Omega_{U}\rightarrow\Omega_{\Phi_{+}},\;\{\varepsilon_{i_{1}}+\varepsilon_{j_{1}},\ldots,\varepsilon_{i_{k}}+\varepsilon_{j_{k}}\}\mapsto\{\varepsilon_{i_{1}}+\varepsilon_{j_{1}},\varepsilon_{i_{1}}-\varepsilon_{j_{1}},\ldots,\varepsilon_{i_{k}}+\varepsilon_{j_{k}},\varepsilon_{i_{k}}-\varepsilon_{j_{k}}\}.$
  • (ii)
    The Poincaré polynomial of $PS_{\Phi_{+}}(q)$ factors into quantum integers, with the multiset of degrees being $\{3,5,\ldots,2k-1\}$ in type $D_{2k}$, $\{3,5,9\}$ in type $E_{7}$, and $\{3,5,9,15\}$ in type $E_{8}$ ([8, Proposition 7.1 (i)]). In particular, the cardinality of $\Omega_{\Phi_{+}}$ is $3\cdot 5\cdot 9=135$ in type $E_{7}$ and $3\cdot 5\cdot 9\cdot 15=2025$ in type $E_{8}$.
  • (iii)
    The unique $W$-minimal element in type $E_{8}$ is the set
    $R=\{\alpha_{2},\alpha_{3},\alpha_{5},\alpha_{7},\theta_{4},\theta_{6},\theta_{7},\theta_{8}\},$
    where $\theta_{4}$, $\theta_{6}$, $\theta_{7}$, and $\theta_{8}$ are the highest roots in the standard parabolic subsystems of types $D_{4}$, $D_{6}$, $E_{7}$, and $E_{8}$ of $\Phi$, respectively. This follows from [8, §6.3], and can also be checked directly using the definition of residues. Since the lowest-degree term of $PS_{\Phi_{+}}(q)$ is the constant 1 by (ii), the $8$-root $R$ has no residues with respect to $\Phi_{+}$.
4.2. An effective construction
The constructions of the following definition turn out to provide extremely effective methods for constructing QP-triples from the root systems of type $ADE$, and many of the resulting QP-triples provide constructions of known objects, in some cases providing extra insight into their structure. We will record the examples without giving any detailed proofs in this section, but we note that the examples in types $A$ and $D$ can all be interpreted in terms of the combinatorics of rook configurations, and we will study the examples $T(E_{8},8)$ and $T(E_{7},7)$ in full detail in Section 5.
Definition 4.6.
Let $W$ be a Weyl group of type $ADE$ with Dynkin diagram $\Gamma$, generating set $S$, and root system $\Phi$. Let $\alpha_{p}$ be a simple root of $\Phi$. We define $U(\Gamma,p)\subseteq{\Phi_{+}}$ to be the set of all roots in which $\alpha_{p}$ appears with odd positive coefficient. We define $\Omega(\Gamma,p):=\Omega_{U(\Gamma,p)}$ and define $T(\Gamma,p)$ to be the triple
[…]
Example 4.7.
  • (i)
    The triple $(W,I,U)$ appearing in Theorem 3.10 is $T(D_{2k},2k)$. In type $D_{n}$, a simple root can only appear with coefficient $\pm 2$, $\pm 1$, or 0, so $U(D_{2k},2k)$ is the set of roots in which $\alpha_{2k}$ appears with coefficient $1$.
  • (ii)
    The triples $T(D_{2k},2k-1),T(D_{2k-1},2k-1),$ and $T(D_{2k-1},2k-2)\}$ are also QP-triples.
    The above three triples all give rise to essentially the same quasiparolic set as $T(D_{2k},2k)$, in the sense that their associated quasiparabolic sets $\Omega_{U(\Gamma,p)}$ all admit canonical level-preserving bijections to the quasiparabolic set $\Omega(D_{2k},2k)$ from (i). This bijection is induced by the symmetry of the Dynkin diagram for $\Omega(D_{2k},2k-1)$. For the case of $\Omega=\Omega(D_{2k-1},2k-1)$, each element of $\Omega$ corresponds to a (non-perfect) matching of $\{1,2,\dots,2k-1\}$ into (k1)(k-1) pairs and one unmatched point, and the canonical bijection from $\Omega$ to $\Omega(D_{2k},2k)$ sends such an element to the $2k$-root in $\Omega(D_{2k},2k)$ corresponding to the unique perfect matching containing those (k1)(k-1) pairs as blocks.
  • (iii)
    The triple $T(A_{n},p)$ is a QP-triple for all integers $n\geq 1$ and $1\leq p\leq n$.
    The residue sets arising from the triple $T(A_{n},p)$ resemble the Rothe diagrams mentioned in Remark 3.14 (i). More specifically, the set $U(A_{n},p)$ consists of the roots of the form $\varepsilon_{i}-\varepsilon_{j}$ where $1\leq i\leq p<j\leq n+1$, so we may identify $U$ with the positions on a $p\times(n+1-p)$ board with rows labeled by 1 to $p$ from top to bottom and columns labelled by $p+1$ to n+1pn+1-p from left to right. Each element of $R\in\Omega_{U(A_{n},p)}$ consists of $\min(p,n+1-p)$ orthogonal roots, and if we view $R$ as an involution in the natural way, as the product $w_{R}\in S_{n+1}$ of the 2-cycles $(i,j)$ for which $\varepsilon_{i}-\varepsilon_{j}\in R$, then a root $\varepsilon_{a}-\varepsilon_{b}\in U(A_{n},p)$ is a residue of $R$ with respect to $U(A_{n},p)$ if and only if we have both $w_{R}(a)<b$ and $w_{R}(b)>a$. Equivalently, a position $(a,b)$ on the board is a residue $R$ if and only if it is not attacked by a rook to its right or above it in the rook configuration corresponding to $R$, as shown in Figure 4, where the residues are marked by stars as in Figure 1.
[…]
    Figure 4. The generalized Rothe diagram of $\{\varepsilon_{1}-\varepsilon_{6},\varepsilon_{2}-\varepsilon_{7},\varepsilon_{3}-\varepsilon_{5}\}\in\Omega_{U(A_{7},3)}$
    Using the rook configuration model, it is possible to deduce that the Poincaré polynomial of $\Omega(A_{n},p)$ factors into quantum integers, as the product $PS_{U(A_{n},p)}=\prod_{D}[d]$ where $D$ is the set (with no repetition) containing each of the $\min(p,n+1-p)$ consecutive integers leading up to (and including) $\max(p,n+1-p)$. The row–column symmetry of the model also implies that there is a natural level-preserving bijection between the quasiparabolic sets $\Omega(A_{n},p)$ and $\Omega(A_{n},n+1-p)$; in particular, the Poincaré polynomials of $U(A_{n},p)$ and $U(A_{n},n+1-p)$ agree.
  • (iv)
    For all integers $n\geq 4$ and $1\leq p\leq n$, the triple $T(D_{n},p)$ is a QP-triple whenever we have $1\leq p\leq\lfloor n/2\rfloor$ (where $\lfloor\cdot\rfloor$ denotes the floor function), but not when $\lfloor n/2\rfloor<p<n-2$. However, if we relax the setting of Definition 4.6 and consider the set $U^{\prime}_{p}=U(D_{n},p)\cap U(D_{n},n)$, then the triple $T^{\prime}(\Gamma,P)=(W(D_{n}),S(D_{n})\setminus\{s_{p}\},U^{\prime}_{p})$ forms a QP-triple for all $1\leq p\leq n-2$. Note that the QP-triple $(W,I,U)$ from Theorem 3.13 is a special case of the triple $T^{\prime}(\Gamma,p)$ from (v), namely, $T^{\prime}(D_{2k},k)$.
    We note that for all the triples $T(D_{n},p)$ mentioned in the last paragraph, the elements $R\in\Omega_{U(D_{n},p)}$ and their residues can be interpreted in terms of rook configurations on a $p\times(n-p)$ board as in (iii), with the caveat that a position on the board may now correspond to either one of two roots $\varepsilon_{i}\pm\varepsilon_{j}$. The fact that $T(D_{n},p)$ fails to be a QP-triple when $\lfloor n/2\rfloor<p<n-2$ has to do with this caveat and the fact that in these cases $R$ does not contribute a rook to every row.
Remark 4.8.
  • (i)
    To summarize Example 4.7, every simple root $\alpha_{p}$ in a root system of type $A_{n}$ or $D_{n}$ gives rise to one or two QP-triples: $T(A_{n},p)$ in type $A_{n}$, $T^{\prime}(D_{n},p)$ in type $D_{n}$, and $T(D_{n},p)$ if $p\leq\lfloor n/2\rfloor$.
  • (ii)
    The rook configuration model described in Example 4.7 (iii) is inspired by a similar model used by Garsia and Remmel in [3] to study a $q$-analogue of the Stirling numbers of the second kind. Garsia and Remmel place rooks on a general Ferrers board, and our model corresponds to the special case where the board is a $a\times b$ rectangle containing $\min(a,b)$ non-attacking rooks, with the residues in our construction corresponding to the positions marked with a circle on the board; see [3, Figure (I.7)].
  • (iii)
    The rook configuration models mentioned in Example 4.7 allow combinatorial proofs of the Poincaré polynomials of the quasiparabolic triples. In particular, it follows from the models that for all the quasiparabolic sets mentioned in the example, the Poincaré polynomials factor into quantum integers (that have combinatorial interpretations). For example, the quasiparabolic sets associated to $T(A_{n},p)$ and $T(A_{n},n+1-p)$ have the same Poincaré polynomial for any $1\leq p\leq\lfloor n+1-p\rfloor$, namely, $\prod_{d=n-2p+2}^{n+1-p}[d]_{q}$.
Example 4.9.
  1. (i)
    Table 1 gives the complete list of cases for which $T(E_{n},p)$ is a QP-triple, where $n\in\{6,7,8\},1\leq p\leq n$, and $\kappa$ stands for the cardinality of the elements in $\Omega_{U}$ in each case. We have verified that this is the complete list by computation, and we will give a detailed treatment of $T(E_{8},8)$ and $T(E_{7},7)$, two examples that have natural connections to minuscule representations and del Pezzo surfaces, but we do not have a combinatorial model or a conceptual proof of the quasiparabolic structure for all the examples. We also note that by our computation, in the cases where $T(E_{n},p)$ is not a QP-triple, it always happens that $T(E_{n},p)$ is not even a scaled $W_{I}$-set.
    $T(\Gamma,p)$ $|U|$ $|\Omega_{U}|$ $\kappa$ degrees
    $T(E_{6},1)$ 16 40 2 5, 8
    $T(E_{6},2)$ 20 30 4 2, 3, 5
    $T(E_{6},6)$ 16 40 2 5, 8
    $T(E_{7},1)$ 32 120 4 3, 5, 8
    $T(E_{7},2)$ 35 30 7 2, 3, 5
    $T(E_{7},5)$ 35 30 7 2, 3, 5
    $T(E_{7},7)$ 27 45 3 5, 9
    $T(E_{8},1)$ 64 240 8 2, 3, 5, 8
    $T(E_{8},2)$ 64 240 8 2, 3, 5, 8
    $T(E_{8},5)$ 64 240 8 2, 3, 5, 8
    $T(E_{8},6)$ 64 240 8 2, 3, 5, 8
    $T(E_{8},8)$ 56 630 4 5, 9, 14
    Table 1. Table of QP-triples of the form $T(E_{n},p)$
  2. (ii)
    The structures of the QP-triple $T(E_{7},2)$ and the associated quasiparolic set $\Omega(E_{7},2)$ are known by [8, Section 6.2]: the parabolic subgroup $W_{I}$ has type $A_{6}$, the set $\Omega(E_{7},2)$ coincides with the set of positive $7$-roots of $\Phi(E_{7})$ containing no alignments, and $W_{I}\cong S_{7}$ acts transitively on $\Omega(E_{7},2)$ in the standard action. The set $\Omega(E_{7},2)$ is also in a natural bijection with the $30$ inequivalent labellings of the Fano plane. Under the correspondence between the $7$-roots and labelled Fano planes, the generalized Rothe diagram of a labelled Fano plane has a simple combinatorial description: a residue is a triple of non-collinear points such that each point has a higher label than the point collinear with the other two. For example, the unique $W_{I}$-minimal labelling
    $\{136,145,127,235,246,347,567\}$
    is the only labelling with no residues, and the unique $W_{I}$-maximal labelling
    $\{123,145,246,257,347,356,167\}$
    has seven residues: $267$, $357$, $367$, $456$, $457$, $467$, and $567$. The generalized Rothe diagram of each labelling can be interpreted as the set of non-positions in the context of games arising from the Steiner system $S(2,3,7)$; see [11].
  3. (iii)
    For the triple $T(E_{8},2)$, the results of [8, Section 6.3] show that the corresponding subgroup $W_{I}$ has type $A_{7}$, the set $\Omega(E_{8},2)$ coincides with the set of positive $8$-roots of $\Phi(E_{8})$, and $W_{I}\cong S_{8}$ acts transitively on $\Omega(E_{8},2)$ in the standard action. The set $\Omega(E_{8},2)$ is in a natural bijection with the set of $240$ inequivalent labellings of the Fano plane by the set $\{1,2,\ldots,8\}$, as well as with the 240 packings of the projective space $PG(3,2)$. The generalized Rothe diagram of each $8$-root in $\Omega(E_{8},2)$ has a characterization similar to the one in (ii), although the details are more complicated. This particularly intricate example is the subject of the paper [6].
5. The QP-triples $T(E_{8},8)$ and $T(E_{7},7)$
This section studies the QP-triples
$T(E_{8},8)=(W(E_{8}),S(E_{8})\setminus\{s_{8}\},U(E_{8},8))\quad\text{and}\quad T(E_{7},7)=(W(E_{7}),S(E_{7})\setminus\{s_{7}\},U(E_{7},7))$
from Example 4.9 in detail. The sets $U(E_{8},8)$ and $U(E_{7},7)$ have interesting natural connections to minuscule representations and del Pezzo surfaces. We give a non-computational proof that $T(E_{8},8)$ is a QP-triple in Section 5.1, and we use this result to deduce that $T(E_{7},7)$ is a QP-triple in Section 5.2. In Section 5.3, we show that the Pfaffian of $T(E_{7},7)$ coincides with the invariant polynomial associated to a 27-dimensional minuscule representation of type $E_{6}$, and we also describe the generalized Rothe diagrams of the elements of $\Omega(E_{7},7)$, which we expect may have applications to Schubert varieties or del Pezzo surfaces.
5.1. $T(E_{8},8)$
Let $\Phi$ be the root system of type $E_{8}$ throughout this subsection. Let $\theta$ be the highest root of $\Phi$, and let $U_{i}=\Phi_{8,i}\subseteq\Phi$ be the set of positive roots of $8$-height $i$ for each $i\in\{0,1,2\}$. It follows from Section 2.1 that $\theta$ is the only element of $U_{2}$, so we have $U(E_{8},8)=U_{1}$, and to prove that $T(E_{8},8)$ is a QP-triple it suffices to show that $\Omega_{U_{1}}$ is a quasiparabolic set for $W_{I}\cong W(E_{7})$, where $I=S(E_{8})\setminus\{s_{8}\}$.
As we will see in Corollary 5.4, the elements of $\Omega_{U}$ turn out to be precisely the features in $\Phi$ whose elements all come from $U_{1}$, and they can also be characterized as the intersections of the form $Q=R\cap U_{1}$ for the $8$-roots $R$ of $\Phi$ that do not contain $\theta$. Our strategy for proving the quasiparabolic structure of $\Omega_{U_{1}}$ is to leverage the quasiparabolic structure of the set $\Omega_{\Phi_{+}}$ of the $8$-roots, so inclusions of the form $Q\subseteq R$ where $Q\in\Omega_{U_{1}}$ and $R\in\Omega_{\Phi_{+}}$ will play an important role in our arguments, as will chains of the form $Q\subseteq H\subseteq R$ where $R$ is an $8$-root, $H\in D_{6}(R)$, and $Q$ is a feature in $R$. More specifically, it will be useful to study the relationship between the residues of $Q,H,$ and $R$ with various $8$-heights in such a chain. We will do so in lemmas 5.25.3 and 5.55.8.
Definition 5.1.
Let $\Phi$ be the root system of type $E_{8}$, and let $R$ be a positive $8$-root of $\Phi$. Suppose that $H\in D_{4}(R)$, or $H\in D_{6}(R)$, or $H=R$. For each $i\in\{0,1,2\}$, we define
$\mathrm{Res}^{(i)}(H):=\text{\rm Res}(H)\cap U_{i}\text{\; and \;}\rho^{({i})}(H)=\lvert\mathrm{Res}^{(i)}(H)\rvert$.
Recall from Section 2.1 that for any $H\in D_{6}(R)$ where $R$ is an $8$-root, the $D_{6}$ subsystem $\Xi:=\text{\rm Span}(H)\cap\Phi$ has positive roots $\{\varepsilon_{i}\pm\varepsilon_{j}:1\leq i<j\leq 6\}$, simple roots $\{\varepsilon_{1}-\varepsilon_{2},\varepsilon_{2}-\varepsilon_{3},\ldots,\varepsilon_{5}-\varepsilon_{6},\varepsilon_{5}+\varepsilon_{6}\}$, and highest root $\varepsilon_{1}+\varepsilon_{2}$ in the coordinate realization of $\Xi$. A simple root of $\Phi$ is necessarily also simple in $\Xi$. By Lemma 2.14, every $Q\subseteq H$ consists of two couples and has the form $\{\varepsilon_{a}\pm\varepsilon_{b},\varepsilon_{c}\pm\varepsilon_{d}\}$ for distinct elements $a,b,c,d\in[6]$, and the set $H\setminus Q$ consists of the couple $\{\varepsilon_{e}\pm\varepsilon_{f}\}$ where $\{a,b,c,d,e,f\}=[6]$. Also recall from Section 2.1 that the highest root $\theta$ has the property that $B(\theta,\alpha_{8})=1$ while $B(\theta,\alpha_{i})=0$ for all the simple roots $\alpha_{i}$ where $1\leq i\leq 7$, so a root $\gamma\in\Phi_{+}$ lies in $U_{i}$ if and only if $B(\theta,\gamma)=i$. Finally, note that for any roots $\beta\in U_{i}$ and $\gamma\in U_{j}$ where $0\leq i<j$, the root $s_{\beta}(\gamma)$ must be a positive root because it has positive $8$-height. We will frequently use the facts mentioned in this paragraph without further comment in the proofs of this subsection.
Lemma 5.2.
Let $R$ be a positive $8$-root of $\Phi$ such that $\theta\notin R$. Suppose that $H\in D_{4}(R)$ or $H\in D_{6}(R)$, and let $\Xi=\text{\rm Span}(H)\cap\Phi$ be the $D_{4}$ or $D_{6}$ subsystem of $\Phi$ corresponding to $H$.
  1. (i)
    We have $\mathrm{Res}^{(2)}(R)=\{\theta\}$ and $\theta\in\text{\rm Res}_{\Phi_{+}}(H)$.
  2. (ii)
    If $\lvert H\cap U_{0}\rvert<4$, then $\mathrm{Res}^{(0)}(H)=\emptyset$.
  3. (iii)
    If $\lvert H\cap U_{1}\rvert<4$, then $\theta\notin\Xi$ and $\mathrm{Res}^{(2)}(H)=\emptyset$.
Proof.
Part (i) holds because $\theta$ has $8$-height 2 while all other positive roots have strictly smaller heights.
We prove (ii) by proving its contrapositive. Suppose that $\mathrm{Res}^{(0)}(H)$ contains a root $\gamma$. Applying Proposition 4.1 (i) to $\Xi=\text{\rm Span}(H)\cap\Phi$ shows that $\gamma$ is the highest or second highest root in a $D_{4}$ subsystem $\Psi\subseteq\Xi$ corresponding to some feature $Q\subseteq H$. Each simple root in a $D_{4}$ subsystem appears with nonzero coefficient in each of the two highest roots, so it follows from the assumption $\gamma\in U_{0}$ that $\Psi\subseteq U_{0}$. In particular, the four elements of $Q$ lie in $U_{0}$, so $\lvert H\cap U_{0}\rvert\geq 4$. The conclusion of (ii) now follows.
Suppose $\lvert H\cap U_{1}\rvert<4$, and suppose for a contradiction that $\theta=\varepsilon_{1}+\varepsilon_{2}\in\Xi$. The assumption $\theta\not\in R$ implies that the perfect matching corresponding to $H$ contains two distinct blocks $\{1,j\}$ and $\{2,k\}$. We then have
$2\theta=(\varepsilon_{1}+\varepsilon_{j})+(\varepsilon_{1}-\varepsilon_{j})+(\varepsilon_{2}+\varepsilon_{k})+(\varepsilon_{2}-\varepsilon_{k})$
in $\Xi$, which implies that each of the four roots on the right hand side has $8$-height 11. This contradicts the assumption that $|H\cap U_{1}|<4$, so we have $\theta\notin\Xi$. It follows that $\theta\notin\text{\rm Res}(H)$, so we have $\mathrm{Res}^{(2)}(H)=\emptyset$, proving (iii). ∎
Lemma 5.3.
Let $R$ be a positive $8$-root of $\Phi$ such that $\theta\not\in R$.
  • (i)
    We have $R=Q_{1}\ \dot{\cup}\ Q_{0}$, where $Q_{i}$ is a feature consisting of roots in $U_{i}$ for each $i$.
  • (ii)
    In the setting of (i), we have $\theta\in\text{\rm Res}(Q_{1})$. We also have
    $\mathrm{Res}^{(2)}(R)=\{\theta\},\quad\mathrm{Res}^{(1)}(R)=\text{\rm Res}_{U_{1}}(Q_{1}),\ \text{and}\quad\mathrm{Res}^{(0)}(R)=\text{\rm Res}(Q_{0}).$
Proof.
Since $\theta\not\in R$, every element of $R$ has $8$-height 0 or $1$. By Proposition 2.9, the support of $\theta$ with respect to $R$ is a feature $Q_{1}:=\{\beta_{1},\beta_{2},\beta_{3},\beta_{4}\}$ such that $\pm\beta_{1}\pm\beta_{2}\pm\beta_{3}\pm\beta_{4}=2\theta$, and $\theta$ is orthogonal to the other four elements of the set $Q_{0}:=R\setminus Q_{1}$. Since $\theta$ has $8$-height 2, it must be the case that $\beta_{1}+\beta_{2}+\beta_{3}+\beta_{4}=2\theta$, so $Q_{1}$ is a feature and we have $Q_{1}\subseteq U_{1}$. Lemma 2.14 (iii) implies that $Q_{0}$ is a feature because $Q_{0}=R\setminus Q_{1}$, and we have $Q_{0}\subseteq U_{0}$ because $B(\theta,\gamma)=0$ for all $\gamma\in Q_{0}$. This proves (i).
The last paragraph shows that $\theta$ lies in the $D_{4}$-subsystem $\Phi\cap\text{\rm Span}(Q_{1})$. Since $\theta\in\text{\rm Res}_{\Phi_{+}}(Q_{1})$ by Lemma 5.2 (i), it follows that $\theta\in\text{\rm Res}(Q_{1})$. This proves the first assertion of (ii).
The equation $\mathrm{Res}^{(2)}(R)=\{\theta\}$ holds by Lemma 5.2 (ii). The fact that $Q_{1}\subseteq R$ implies that $\mathrm{Res}^{(1)}(R)=\text{\rm Res}_{\Phi_{+}}(R)\cap U_{1}\subseteq\text{\rm Res}_{U_{1}}(Q_{1})$. Conversely, we have $\text{\rm Res}_{U_{1}}(Q_{1})\subseteq\text{\rm Res}_{\Phi_{+}}(R)\cap U_{1}$ because $s_{\beta}(\gamma)>0$ for all $\beta\in Q_{0}\subseteq U_{0}$ and $\gamma\in U_{1}$, so we have $\mathrm{Res}^{(1)}(R)=\text{\rm Res}_{U_{1}}(Q_{1})$. Finally, we have $\text{\rm Res}(R)=\bigcup_{Q\in D_{4}(R)}\text{\rm Res}(Q)$ by Proposition 4.1 (i) and therefore $\mathrm{Res}^{(0)}(R)=\bigcup_{Q\in D_{4}(R)}\mathrm{Res}^{(0)}(Q)$. Every feature $Q$ other than $Q_{0}$ appearing in the union satisfies $\lvert Q\cap U_{0}\rvert<4$ and thus $\mathrm{Res}^{(0)}(Q)=\emptyset$ by Lemma 5.2 (ii), so it follows that $\mathrm{Res}^{(0)}(R)=\mathrm{Res}^{(0)}(Q_{0})$. This completes the proof of (ii). ∎
Corollary 5.4.
We have $\Omega_{U_{1}}=\{R\cap U_{1}:R\in\Omega_{\Phi_{+}},\theta\notin R\}=\{Q\subseteq U_{1}:Q\text{ is a feature}\}$.
Proof.
We have $B(\gamma,\theta)=1\neq 0$ for all $\gamma\in U_{1}$, so it follows from Lemma 2.3 that every element $Q\in\Omega_{U_{1}}$ can be extended to an $8$-root $R$ not containing $\theta$. The asserted equalities now follow from Lemma 5.3 (i). ∎
Lemma 5.5.
Let $R$ be a positive $8$-root of $\Phi$ such that $\theta\notin R$. Let $H\in D_{6}(R)$, let $\Xi=\text{\rm Span}(H)\cap\Phi$ be the associated $D_{6}$ subsystem of $\Phi$, and suppose that we have $|H\cap U_{1}|=2$ and $|H\cap U_{0}|=4$.
  • (i)
    We have $\theta\notin\Xi$.
  • (ii)
    The highest root $\psi:=\varepsilon_{1}+\varepsilon_{2}$ of $\Xi$ has $8$-height 11.
  • (iii)
    The simple root $\beta_{1}:=\varepsilon_{1}-\varepsilon_{2}$ of $\Xi$ has $8$-height 11; the other simple roots of $\Xi$ have $8$-height 0.
  • (iv)
    A root $\gamma\in\Xi_{+}$ has $8$-height 11 if $\gamma=\varepsilon_{1}\pm\varepsilon_{i}$ for some $i>1$, and has $8$-height 0 otherwise.
Proof.
Part (i) follows from Lemma 5.2 (iii) and the assumption that $\lvert H\cap U_{1}\rvert=2<4$.
Since $\theta\not\in\Xi$, the highest root $\psi$ of $\Xi$ has $8$-height at most 1. On the other hand, since $H\cap U_{1}\neq\emptyset$ by assumption, $H$ contains a root of $8$-height 1 and hence so does $\Xi$. It follows that $\psi\in U_{1}$, which proves (ii).
We now prove the claim that $\beta_{1}\in U_{1}$. The set $Q_{0}:=H\cap U_{0}$ has size 4 by assumption and is thus a feature in $U_{0}$ by Lemma 5.3 (i). Lemma 2.14 (ii) implies that $Q_{0}$ contains two couples in $\Xi$ while $H\cap U_{1}=H\setminus U_{0}$ consists of one couple. We have just proved that $\psi=\varepsilon_{1}+\varepsilon_{2}$ lies in $U_{1}$, so if $\psi\in H$ then we have $H\cap U_{1}=\{\varepsilon_{1}\pm\varepsilon_{2}\}$ and thus $\beta_{1}\in U_{1}$. If $\psi\notin H$, then the perfect matching of the set $[6]$ corresponding to $H$ contains two distinct pairs $\{1,j\}$ and $\{2,k\}$. These four roots $\varepsilon_{1}\pm\varepsilon_{j},\varepsilon_{2}\pm\varepsilon_{k}\in H$ sum to $2\psi$ in this case, so either $\{1,j\}$ corresponds to a pair of $8$-height 11 and $\{2,k\}$ corresponds to a pair of $8$-height 0, or vice versa. One the other hand, the root $\beta_{1}\in\Xi_{+}\subseteq\Phi_{+}$ has nonnegative $8$-height and satisfies
$2\beta_{1}=((\varepsilon_{1}+\varepsilon_{j})+(\varepsilon_{1}-\varepsilon_{j}))-((\varepsilon_{2}+\varepsilon_{k})+(\varepsilon_{2}-\varepsilon_{k})).$
It follows that $\varepsilon_{1}\pm\varepsilon_{j}\in U_{1},\varepsilon_{2}\pm\varepsilon_{k}\in U_{0}$, and $\beta_{1}\in U_{1}$, as claimed.
Since every simple root of $\Xi$ appears in the decomposition of $\psi$ into simple roots and $\psi$ and $\beta_{1}$ have the same $8$-height, it follows that all simple roots of $\Xi$ other than $\beta$ lie in $U_{0}$, proving (iii).
Part (iv) follows from (iii) by expressing $\gamma$ as a linear combination of simple roots. ∎
Lemma 5.6.
Let $R$ be a positive $8$-root in type $E_{8}$ such that $\theta\notin R$, and let $R=Q_{0}\ \dot{\cup}\ Q_{1}$ be the decomposition of Lemma 5.3 (i). Suppose that $Q$ is a feature in $R$ such that $\lvert Q\cap Q_{0}\rvert=\lvert Q\cap Q_{1}\rvert=2$, and let $H=Q\cup Q_{0}$.
  • (i)
    All residues of $Q$ have $8$-height 1, so that $\text{\rm Res}(Q)=\mathrm{Res}^{(1)}(Q)$.
  • (ii)
    We have $H\in D_{6}(R)$, and we also have
    $\mathrm{Res}^{(2)}(H)=\{\theta\},\quad\mathrm{Res}^{(1)}(H)=\text{\rm Res}_{U_{1}}(Q),\ \text{and}\quad\mathrm{Res}^{(0)}(H)=\text{\rm Res}(Q_{0}).$
Proof.
The assumptions that $\lvert Q\cap Q_{0}\rvert=2<4$ and $\lvert Q\cap Q_{1}\rvert=2<4$ imply that $\mathrm{Res}^{(0)}(Q)=\emptyset$ and $\mathrm{Res}^{(2)}(Q)=\emptyset$ by Lemma 5.2 (ii) and (iii), respectively, and (i) follows.
The fact that $H\in D_{6}(R)$ follows from Lemma 2.14 (i). Let $\Xi=\text{\rm Span}(H)\cap\Phi$ be the corresponding $D_{6}$ subsystem in $\Phi$. We have $\theta\notin\Xi$ by Lemma 5.5 (i), so $\theta\notin\text{\rm Res}(H)$ and $\mathrm{Res}^{(2)}(H)=\emptyset$, proving the first equality in (ii). The two remaining equalities follow from Proposition 4.1 (i) and Lemma 5.2 (ii) in the same way as the last two equalities in Lemma 5.3 (ii). ∎
Lemma 5.7.
Maintain the setting of Lemma 5.6, and suppose that $Q$ is the support of a root $\alpha\in\Phi\setminus R$. Let $r=s_{\alpha}$ be the reflection corresponding to $\alpha$.
  • (i)
    If $\rho(Q)<\rho(rQ)$ , then we have
    (5.1) $\rho^{({1})}(rH)-\rho^{({1})}(H)\geq\rho(rQ)-\rho(Q).$
  • (ii)
    If $\alpha$ is a simple root of $\Phi$, then we have $\rho(Q_{0})=\rho(rQ_{0})$.
Proof.
Let $\Xi=\text{\rm Span}(H)\cap\Phi$, as in the proof of Lemma 5.6. Lemma 5.5 (iv) implies that in the coordinate realization of $\Xi$ we have $\Xi\cap U_{1}=\{\varepsilon_{1}\pm\varepsilon_{i}:2\leq i\leq 6\}$, so it follows that
(5.2) $H=\{\varepsilon_{1}\pm\varepsilon_{a},\varepsilon_{b}\pm\varepsilon_{c},\varepsilon_{d}\pm\varepsilon_{e}\}$
where $b<c,d<e$, and $\{1,a,b,c,d,e\}=[6]$, and where we may assume that $Q=\{\varepsilon_{1}\pm\varepsilon_{a},\varepsilon_{b}\pm\varepsilon_{c}\}$. The assumptions on $\alpha$ and $Q$ implies that $Q$ and $rQ$ are two distinct features in the same $D_{4}$ system $\text{\rm Span}(Q)\cap\Phi$, so we have $rQ=\{\varepsilon_{1}\pm\varepsilon_{a^{\prime}},\varepsilon_{b^{\prime}}\pm\varepsilon_{c^{\prime}}\}$ where $\{a^{\prime},b^{\prime},c^{\prime}\}=\{a,b,c\}$ but $a\neq a^{\prime}$.
Let $\beta\in\Xi\cap U_{1}$. If we have $\beta=\varepsilon_{1}-\varepsilon_{i}$ for some $2\leq i\leq 6$ or $\beta=\varepsilon_{1}+\varepsilon_{i}$ for some $i>a$, then we have $s_{\gamma}(\beta)<0$ for $\gamma=\varepsilon_{1}+\varepsilon_{a}\in H$, so $\beta\notin\text{\rm Res}(H)$. On the other hand, if $\beta=\varepsilon_{1}+\varepsilon_{i}$ for some $1<i<a$ then we have $s_{\gamma}(\beta)>0$ for all $\gamma\in H$ and hence $\beta\in\text{\rm Res}(H)$. It follows that $\mathrm{Res}^{(1)}(H)=\{\varepsilon_{1}+\varepsilon_{i}:1<i<a\}$ and $\rho^{({1})}(H)=a-2$. Similarly, we have $\rho^{({1})}(rH)=a^{\prime}-2$, so that $\rho^{({1})}(rH)-\rho^{({1})}(H)=a^{\prime}-a$. By Remarks 2.13 and 3.5 (ii), the assumption that $\rho(rQ)>\rho(Q)$ implies that we have one of the following situations:
  1. (1)
    $Q$ is an alignment and $rQ$ is a crossing, so that $a<b=a^{\prime}<c$ and $\rho(rQ)-\rho(Q)=1$;
  2. (2)
    $Q$ is crossing and $rQ$ is a nesting, so that $b<a<c=a^{\prime}$ and $\rho(rQ)-\rho(Q)=1$;
  3. (3)
    $Q$ is an alignment and $rQ$ is a nesting, so that $a<b<c=a^{\prime}$ and $\rho(rQ)-\rho(Q)=2$.
We have $\rho^{({1})}(rH)-\rho^{({1})}(H)=a^{\prime}-a\geq\rho(rQ)-\rho(Q)$ in all these cases, which proves (i). Now suppose that $\alpha$ is a simple root of $\Phi$. This implies that $\alpha$ is also a simple root of $\Xi$, so the reflection $r$ induces a basic transposition $s_{i}:=(i,i+1)$ on the perfect matching $M:=\{\{1,a\},\{b,c\}\}$ corresponding to $Q$, with $\{i,i+1\}\subseteq\{1,a,b,c\}$ because $Q$ is the support of $\alpha$. If $\{i,i+1\}\in M$, then $rH=H$ and (ii) follows. Otherwise, $s_{i}$ must move either $1$ or $a$ to an adjacent integer, and in either case the numbers $(b,c,d,e)$ and $(s_{i}(b),s_{i}(c),s_{i}(d),s_{i}(e))$ appearing in the matchings corresponding to $Q_{0}$ and $rQ_{0}$ must be in the same relative order. It follows that $\rho(Q_{0})=\rho(rQ_{0})$, proving (ii). ∎
Lemma 5.8.
Maintain the hypotheses of Lemma 5.7, and suppose further that $\alpha$ is a root in $\Phi(E_{7})\subseteq\Phi$. If $\rho(Q)<\rho(rQ)$, then we have $\rho_{U_{1}}(Q_{1})<\rho_{U_{1}}(rQ_{1})$.
Proof.
We first note that the assumption that $\alpha\in\Phi(E_{7})$ implies that $r$ cannot change the $8$-height of a root, so we have $\theta\notin rR$ and $rR=rQ_{0}\ \dot{\cup}\ rQ_{1}$ is the decomposition of $rR$ into a feature in $U_{0}$ and a feature in $U_{1}$ as in Lemma 5.3 (i). The conclusions of Lemma 5.6 therefore still hold if we replace $Q,Q_{0},H$ and $R$ with $rQ,rQ_{0},rH$ and $rR$, respectively. Recall that $H=Q\cup Q_{0}$.
To prove $\rho_{U_{1}}(Q_{1})<\rho_{U_{1}}(rQ_{1})$, it suffices to prove that $\rho^{({1})}(R)<\rho^{({1})}(rR)$ by Lemma 5.3 (ii). We have
$\text{\rm Res}(R)=\text{\rm Res}(Q)\ \dot{\cup}\ \text{\rm Res}(R\setminus Q)\ \dot{\cup}\ \dot{\bigcup}_{Q\subseteq H^{\prime}\in D_{6}(R)}(\text{\rm Res}(H^{\prime})\setminus\text{\rm Res}(Q))$
by Proposition 4.1 (ii). Taking intersections with $U_{1}$ then implies that
$\mathrm{Res}^{(1)}(R)=\mathrm{Res}^{(1)}(Q)\ \dot{\cup}\ \mathrm{Res}^{(1)}(R\setminus Q)\ \dot{\cup}\ \dot{\bigcup}_{Q\subseteq H^{\prime}\in D_{6}(R)}(\mathrm{Res}^{(1)}(H^{\prime})\setminus\text{\rm Res}(Q),$
where we have $\mathrm{Res}^{(1)}(Q)=\text{\rm Res}(Q)$ by Lemma 5.6 (i) and $\text{\rm Res}(Q)=\mathrm{Res}^{(1)}(Q)\subseteq\mathrm{Res}^{(1)}(H^{\prime})$ for all $Q\subseteq H^{\prime}\in D_{6}(R)$ by Proposition 4.1 (i). Taking cardinalities of the sets in the last equation then yields the equation
(5.3) $\rho^{({1})}(R)=\rho(Q)+\rho^{({1})}(R\setminus Q)+(\rho^{({1})}(H)-\rho(Q))+\sum_{Q\subseteq H^{\prime}\in D_{6}(R)\setminus\{H\}}(\rho^{({1})}(H^{\prime})-\rho(Q)).$
By the first paragraph of this proof, we also similarly have
(5.4) $\rho^{({1})}(rR)=\rho(rQ)+\rho^{({1})}(rR\setminus rQ)+(\rho^{({1})}(rH)-\rho(rQ))+\sum_{rQ\subseteq H^{\prime}\in D_{6}(R)\setminus\{rH\}}(\rho^{({1})}(H^{\prime})-\rho(rQ)).$
We have $rR\setminus rQ=R\setminus Q$ since $r$ fixes every element in $R\setminus Q$, and we have
$\{H^{\prime}:rQ\subseteq H^{\prime}\in D_{6}(R)\setminus\{rH\}\}=\{rH^{\prime}:Q\subseteq H^{\prime}\in D_{6}(R)\setminus\{H\}\},$
so subtracting Equation (5.3) from Equation (5.4) yields that
(5.5) $\rho^{({1})}(rR)-\rho^{({1})}(R)$ $=$ $(\rho(rQ)-\rho(Q))+((\rho^{({1})}(rH)-\rho^{({1})}(H))-(\rho(rQ)-\rho(Q)))$
$+\sum_{Q\subseteq H^{\prime}\in D_{6}(R)\setminus\{H\}}\left((\rho^{({1})}(rH^{\prime})-\rho^{({1})}(H^{\prime}))-(\rho(rQ)-\rho(Q))\right).$
On the right side of Equation (5.5), we have $\rho(rQ)-\rho(Q)>0$ by assumption, and we have
$(\rho^{({1})}(rH)-\rho^{({1})}(H))-(\rho(rQ)-\rho(Q))>0$
by Lemma 5.7 (i). For any $Q\subseteq H^{\prime}\subseteq D_{6}(R)\setminus\{H\}$, we have $\lvert H^{\prime}\cap U_{0}\rvert=\lvert rH^{\prime}\cap U_{0}\rvert<4$ and therefore that $\mathrm{Res}^{(0)}(H^{\prime})=\mathrm{Res}^{(0)}(rH^{\prime})=\emptyset$ by Lemma 5.2 (ii). Furthermore, we have either $H^{\prime}=Q_{1}\cup Q$ or $\lvert H^{\prime}\cap Q_{1}\rvert<4$. In the former case, Lemma 5.3 (ii) and Proposition 4.1 (i) imply that $\theta\in\text{\rm Res}(Q_{1})\subseteq\text{\rm Res}(H^{\prime})$ and similarly $\theta\in\text{\rm Res}(rQ_{1})\subseteq\text{\rm Res}(rH^{\prime})$, so we have $\mathrm{Res}^{(2)}(H^{\prime})=\mathrm{Res}^{(2)}(rH^{\prime})=\{\theta\}$. In the latter case, we have $\mathrm{Res}^{(2)}(H^{\prime})=\mathrm{Res}^{(2)}(rH^{\prime})=\emptyset$ by Lemma 5.2 (iii). It follows that $\rho^{({0})}(H^{\prime})=\rho^{({0})}(rH^{\prime})$ and $\rho^{({2})}(H^{\prime})=\rho^{({2})}(H^{\prime})$ in all cases, so we have
$(\rho^{({1})}(rH^{\prime})-\rho^{({1})}(H^{\prime}))-(\rho(rQ)-\rho(Q))=(\rho(rH^{\prime})-\rho(H^{\prime}))-(\rho(rQ)-\rho(Q))>0$
for any $Q\subseteq H^{\prime}\in D_{6}(R)\setminus\{H\}$, where the inequality holds by an application of Corollary 4.4 to the $6$-root $H^{\prime}$ of the $D_{6}$ system $\text{\rm Span}(H^{\prime})\cap\Phi$. Equation (5.5) now implies that $\rho^{({1})}(rR)>\rho^{({1})}(R)$, as desired. ∎
Theorem 5.9.
[…]
Proof.
Let $Q_{1}\in\Omega_{U}$, and let $r=s_{\beta}$ be a reflection in $W_{I}$ corresponding to a positive root $\beta\in\Phi(E_{7})=\Phi_{8,0}$. Corollary 5.4 and Lemma 5.3 imply that we can extend $Q_{1}$ to an $8$-root $R=Q_{1}\ \dot{\cup}\ Q_{0}$ where $\theta\notin R$ and $Q_{i}$ is a feature in $U_{i}$ for each $i$. Since $r$ cannot change the $8$-heights of roots, we also have $rR=rQ_{0}\ \dot{\cup}\ rQ_{1}$, where $\theta\notin rR$ and $rQ_{i}$ is a feature in $U_{i}$ for each $i$.
Note that if $rQ_{1}\neq Q_{1}$, then since $rQ_{1}=rR\cap U_{1}$ and $Q_{1}=R\cap U_{1}$ we must have $rR\neq R$, so that $\beta\notin R$. It then follows from Proposition 2.9, Remark 2.13, and Remark 3.5 (ii) that the support of $\beta$ with respect to the basis $R$ is a feature in $R$, $Q$, such that $\rho(rQ)\neq\rho(Q)$. The assumption $rQ_{1}\neq Q_{1}$ also implies that $Q\neq Q_{0}$, because if $Q=Q_{0}$ then $r$ would fix every element of $Q_{1}$, so Lemma 2.14 (i) implies that either $Q=Q_{1}$ or $Q$ is a feature in $R$ such that $\lvert Q\cap Q_{0}\rvert=\lvert Q\cap Q_{1}\rvert=2$. Moreover, if $Q=Q_{1}$, then we have $rQ_{0}=Q_{0}$, so the set equalities asserted in Lemma 5.3 (ii) imply that
(5.6) $\lvert\rho_{U}(sQ_{1})-\rho_{U}(Q_{1})\rvert=\lvert(\rho(sR)-1-\rho(sQ_{0}))-(\rho(R)-1-\rho(Q_{0}))\rvert=\lvert\rho(sR)-\rho(R)\rvert$.
Suppose that $s=s_{\alpha}$ is a simple reflection in $W_{I}$ such that $sQ_{1}\neq Q_{1}$, and let $Q^{\prime}$ be the support of $\alpha$ with respect to $R$. By the last paragraph, we have either $Q^{\prime}=Q_{1}$ or $Q^{\prime}$ is a feature such that $\lvert Q^{\prime}\cap Q_{0}\rvert=\lvert Q^{\prime}\cap Q_{1}\rvert=2$. In the former case we have $sQ_{0}=Q_{0}$, and in the latter case we have $\rho(Q_{0})=\rho(sQ_{0})$ by Lemma 5.7 (ii). Lemma 5.3 (ii) implies Equation (5.6) in both these cases, so Theorem 4.3 implies that
$\lvert\rho_{U}(sQ_{1})-\rho_{U}(Q_{1})\rvert=\lvert\rho(sR)-\rho(R)\rvert\leq 1.$
It follows that $\Omega_{U}$ is a scaled set.
Suppose $rQ_{1}\neq Q_{1}$, and let $Q$ be the support of $\beta$ with respect to $R$. By the second paragraph, we have $\rho(rQ)\neq\rho(Q)$, and we have either $Q=Q_{1}$ or $Q$ is a feature such that $\lvert Q\cap Q_{0}\rvert=\lvert Q\cap Q_{1}\rvert=2$. Replacing $Q$ with $rQ$ if necessary, we may assume that $\rho(Q)<\rho(rQ)$. This implies that $\rho(R)<\rho(rR)$ by Corollary 4.4. If $Q=Q_{1}$, then we have $\rho_{U}(Q_{1})-\rho_{U}(rQ_{1})=\rho(R)-\rho(rR)<0$ by Equation (5.6). Otherwise, we have $\rho_{U}(Q_{1})<\rho_{U}(rQ_{1})$ by Lemma 5.8. It follows that $\Omega_{U}$ satisfies condition (QP1).
For any reflection $r=s_{\beta}\in W_{I}$ corresponding to a positive root $\beta\notin R$, Corollary 4.4 implies that we have $\rho(Q)<\rho(rQ)$ if and only if we have $\rho(R)<\rho(rR)$. The case discussion in the last paragraph thus implies that for any reflection $r\in W_{I}$ and any $8$-root $R=Q_{0}\ \dot{\cup}\ Q_{1}$ extending an element $Q_{1}\in\Omega_{U}$, if we have $\rho_{U}(rQ_{1})>\rho_{U}(Q_{1})$ then we have $\rho(rR)>\rho(R)$.
Now suppose that $s=s_{\alpha}$ is a simple reflection of $W_{I}$ and $r=r_{\beta}$ a reflection of $W_{I}$ such that $\rho_{U}(rQ_{1})>\rho_{U}(Q_{1})$ and $\rho_{U}(srQ_{1})<\rho_{U}(sQ_{1})$. We have $srQ_{1}\in\Omega_{U}$ and $sQ_{1}=(srs)(srQ_{1})$, where srssrs is also a reflection in $W_{I}$, so by the last paragraph we have $\rho(rR)>\rho(R)$ and $\rho(sR)>\rho(srR)$. This implies that $sR=rR$ by Theorem 4.3, and taking intersections with $U_{1}$ yields $sQ_{1}=rQ_{1}$, so $\Omega_{U}$ satsfies condition (QP2). This completes the proof. ∎
Remark 5.10.
  1. (i)
    Using Remark 4.5 (ii) and Lemma 5.3 (i), one can show that exactly 135 of the positive $8$-roots in $\Omega_{\Phi_{+}}$ contain $\theta$ and that the remaining 1890 positive $8$-roots are in a 3-to-1 correspondence with the elements of the set $\Omega_{U}$ from Theorem 5.9. The 33-to-11 ratio comes from the fact the each element $Q_{1}\in\Omega_{U}$ can be extended in three ways to a positive $8$-root $R=Q_{1}\ \dot{\cup}\ Q_{0}$, where $Q_{0}$ is one of the three features in the $D_{4}$-subsystem given by $\Psi=\{\beta\in\Phi:B(\beta,\gamma)=0\text{ for all }\gamma\in Q_{1}\}$. Combining this fact with Lemma 5.3 (ii), Remark 3.5 (ii), and Remark 4.5 (ii) shows that the Poincaré polynomial of $\Omega_{U}$ is $PS_{U}(q)=[5]_{q}[9]_{q}[14]_{q}$, which agrees with Table 1. In particular, the set $\Omega_{U}$ contains $5\cdot 9\cdot 14=630$ triples.
  2. (ii)
    Recall from Remark 4.5 (iii) that the $8$-root
    $R=\{\alpha_{2},\alpha_{3},\alpha_{5},\alpha_{7},\theta_{4},\theta_{6},\theta_{7},\theta_{8}\}$
    is the unique $W(E_{8})$-minimal element of $\Omega_{\Phi_{+}}$ and has no residues with respect to $\Phi_{+}$. It follows from Theorem 4.3 that $s_{8}(R)$ has a single residue, namely, $\theta_{8}$. Lemma 5.3 (ii) then implies that the unique $W(E_{7})$-minimal element of $\Omega_{U}$ is the quadruple $\{s_{8}(\alpha_{7}),s_{8}(\theta_{6}),s_{8}(\theta_{7}),s_{8}(\theta)\}$.
5.2. $T(E_{7},7)$
In this subsection, we maintain the notation of Theorem 5.9, identify the root system $\Phi(E_{7})$ with the subsystem $\Phi^{\prime}=\Phi_{8,0}$ of $\Phi$ in the usual way, and identify the set $U^{\prime}=U(E_{7},7)$ with the subset of $\Phi^{\prime}_{+}$ consisting of roots of positive odd $7$-height. Let $\theta^{\prime}=\theta_{7}$ be the highest root of $\Phi^{\prime}$. The $7$-height of $\theta^{\prime}$ is 1 by Equation (2.5), so it follows that $U^{\prime}=\Phi^{\prime}_{7,1}$. We will use Theorem 5.9 to prove that $T(E_{7},7)$ is a QP-triple in this subsection. In other words, we will prove that $\Omega_{U^{\prime}}$ is a quasiparabolic set for $W_{I^{\prime}}\cong W(E_{6})$ with level function $\rho_{U^{\prime}}$, where $I^{\prime}=S\setminus\{s_{8},s_{7}\}$.
We note that the set $U=U(E_{8},8)$ has a unique maximal element, which is also the second highest root of $\Phi$, namely, $\eta:=s_{8}(\theta)=\theta-\alpha_{8}$.
Lemma 5.11.
Let $Q\in\Omega_{U}$ and suppose that $\eta\in Q$.
  • (i)
    We have $\alpha_{8}\not\in\text{\rm Res}_{U}(Q)$.
  • (ii)
    Every element of $\text{\rm Res}_{U}(Q)$ is orthogonal to $\eta$.
Proof.
We have $B(\alpha_{8},\eta)=-1$, which implies that $\alpha_{8}\not\in Q$. The elements of $Q$ sum to $2\theta$, so we have $B(\alpha_{8},2\theta)=+2$. This can only happen if every $\beta\in Q\backslash\{\eta\}$ satisfies $B(\alpha_{8},\beta)=+1$, which implies that $s_{\beta}(\alpha_{8})<0$ for all $\beta\in Q\setminus\{\eta\}$, proving (i).
Let $\gamma\in\text{\rm Res}_{U}(Q)$. Remark 3.2 (i) implies that $\gamma\neq\eta$, so we have $B(\eta,\gamma)\neq 2$. If it were the case that $B(\eta,\gamma)=-1$, it would follow that $\eta+\gamma=\theta$ by considering $8$-heights. This would imply that $\gamma=\alpha_{8}$, contradicting (i). If we had $B(\eta,\gamma)=+1$, it would follow that $s_{\eta}(\gamma)<0$, and $\gamma$ would not be a residue. We conclude that $B(\eta,\gamma)=0$, and (ii) follows. ∎
Lemma 5.12.
  • (i)
    The simple reflection $s_{8}$ gives a bijection from $U^{\prime}$ to the set
    $\eta_{U}^{\perp}:=\{\beta\in U:B(\beta,\eta)=0\}.$
  • (ii)
    Each element $T\in\Omega_{U^{\prime}}$ has size 3, so that $\Omega_{U^{\prime}}$ consists of triples of pairwise orthogonal roots from $U^{\prime}$.
  • (iii)
    There is an injective $W(E_{6})$-equivariant map $\iota:\Omega_{U^{\prime}}\rightarrow\Omega_{U}$ given by
    $\iota(\{\beta_{1},\beta_{2},\beta_{3}\})=\{s_{8}(\beta_{1}),s_{8}(\beta_{2}),s_{8}(\beta_{3}),\eta\}.$
  • (iv)
    For any distinct roots $\beta,\gamma\in U^{\prime}$, we have $s_{\beta}(\gamma)<0$ if and only if $s_{s_{8}(\beta)}(s_{8}(\gamma))<0$.
  • (v)
    For every $T\in\Omega_{U^{\prime}}$, we have a bijection $\text{\rm Res}_{U^{\prime}}(T)\rightarrow\text{\rm Res}_{U}(\iota(T))$ sending each $\gamma\in\text{\rm Res}_{U^{\prime}}(T)$ to $s_{8}(\gamma)$. In particular, we have $\rho_{U^{\prime}}(T)=\rho_{U}(\iota(T))$.
Proof.
Let
$X_{i,j}=\{\alpha\in\Phi_{+}:\text{\rm ht}_{7}(\alpha)=i\text{ and }\text{\rm ht}_{8}(\alpha)=j\}$
for all $i,j\geq 0$. Direct computation shows that $s_{8}(X_{i,0})\subseteq X_{i,i}$ and $s_{8}(X_{i,1})\subseteq X_{i,i-1}$ for all $i\geq 0$, and we have $B(\beta,\eta)=B(s_{8}(\beta),\theta)=\text{\rm ht}_{8}(s_{8}(\beta))$ for any $\beta\in\Phi$. It follows that for any $\gamma\in U^{\prime}=X_{1,0}$, we have $s_{8}(\gamma)\in X_{1,1}\subseteq U$ and $B(s_{8}(\gamma),\eta)=\text{\rm ht}_{8}(\gamma)=0$, so $s_{8}(\gamma)\in\eta_{U}^{\perp}$. It also follows that for any $\beta\in\eta_{U}^{\perp}$ the element $\gamma=s_{8}(\beta)$ has $8$-height 0, which implies that $\gamma\in X_{1,0}=U^{\prime}$. Part (i) follows.
There exist orthogonal triples in $U^{\prime}$}, such as the triple $\{s_{8}(\alpha_{7}),s_{8}(\theta_{6}),s_{8}(\theta_{7})\}$ (see Remark 5.10 (ii)). For every $\beta\in U^{\prime}$, the root $s_{8}(\beta)$ has $8$-height 1 and 7-height 1, so (i) implies that if $T\in\Omega_{U^{\prime}}$ then we have $s_{8}(T)\cup\{\eta\}\in\Omega_{U}$. The set $s_{8}(T)\cup\{\eta\}$ has size 4 in this case by Corollary 5.4, so $T$ has size 3, proving (ii).
The last paragraph shows that the map $\iota$ given in (iii) sends elements of $\Omega_{U^{\prime}}$ to $\Omega_{U}$. The map $\iota$ is injective because $T=s_{8}(\iota(T)\setminus\{\eta\})$ for all $T\in\Omega_{U}$, and it is $W(E_{6})$-equivariant because every simple root in $S\setminus\{\alpha_{8},\alpha_{7}\}$ is orthogonal to both $\alpha_{8}$ and $\eta$. Part (iii) follows.
For any distinct roots $\beta,\gamma\in U^{\prime}$, we have
$s_{s_{8}(\beta)}(s_{8}(\gamma))=s_{8}(\gamma)-B(s_{8}(\beta),s_{8}(\gamma))s_{8}(\beta)=(\gamma+\alpha_{8})-B(\beta,\gamma)(\beta+\alpha_{8}).$
It follows that $s_{\beta}(\gamma)<0$ if and only if $s_{s_{8}(\beta)}(s_{8}(\gamma))<0$, because both conditions hold if and only if we have both $B(\beta,\gamma)=1$ and $\gamma<\beta$. This proves (iv).
Let $T=\{\beta_{1},\beta_{2},\beta_{3}\}\in\Omega_{U^{\prime}}$ and $\gamma\in\text{\rm Res}_{U^{\prime}}(T)$. Since $B(\eta,s_{8}(\gamma))=B(\theta,\gamma)=0$, we have $s_{s_{8}(\gamma)}(\eta)=\eta>0$. For each $i\in\{1,2,3\}$, we have $\gamma\neq\beta_{i}$ by Remark 3.2 (i), and we have $s_{\beta_{i}}(\gamma)>0$ since $\gamma\in\text{\rm Res}_{U^{\prime}}(T)$, so (iv) implies that $s_{s_{8}(\beta_{i})}(s_{8}(\gamma))>0$. It follows that $s_{8}(\gamma)\in\text{\rm Res}_{U}(\iota(T))$.
Conversely, for any $\delta\in\text{\rm Res}_{U}(\iota(T))$, we have $\delta\in\eta_{U}^{\perp}$ by Lemma 5.11 (ii), so we have $\delta=s_{8}(\gamma)$ for some $\gamma\in U^{\prime}$ by (i). The assumption that $\delta\in\text{\rm Res}_{U}(\iota(T))$ implies that $\delta\notin\iota(T)$ and, therefore, that $\gamma\notin T$. Since $\delta=s_{8}(\gamma)\in\text{\rm Res}_{U}(\iota(T))$, it follows from (iv) that $\gamma\in\text{\rm Res}_{U^{\prime}}(T)$, which proves the first assertion of (v). The second assertion follows from the first by taking cardinalities. ∎
Theorem 5.13.
[…]
Proof.
The set $\Omega_{U^{\prime}}$ is a $W_{I^{\prime}}$-set because elements of $I$ do not change $7$-heights of roots. Lemma 5.12 (iii) and (v) imply that for any $T\in\Omega_{U^{\prime}}$ and any reflection $r\in W_{I^{\prime}}$, we have $\rho_{U^{\prime}}(T)=\rho_{U}(\iota(T))$ and $\rho_{U^{\prime}}(rT)=\rho_{U}(\iota(rT))=\rho_{U}(r(\iota(T))$ for the element $\iota(T)\in\Omega_{U}$. The fact that $\Omega_{U^{\prime}}$ is a quasiparabolic $W(E_{6})$-set with respect to $\rho_{U^{\prime}}$ now follows from Theorem 5.9 and the injectivity of $\iota$ by a routine verification. ∎
Remark 5.14.
  1. (i)
    Recall from Remark 4.5 (ii) that there are $135$ positive $7$-roots in type $E_{7}$, and that the Poincaré polynomial of the associated to these 7-roots is $PS_{\Phi^{\prime}_{+}}(q)=[3]_{q}[5]_{q}[9]_{q}$. The $7$-roots are in a 3–to–1 correspondence with the elements of $\Omega_{U}^{\prime}$ similar to the correspondence mentioned in Remark 5.10 (i). One can thus use the correspondence to deduce that $PS_{U^{\prime}}(q)=[5]_{q}[9]_{q}$ and $\lvert\Omega_{U^{\prime}}\rvert=45$.
  2. (ii)
    Combining Remark 5.10 (ii) and Lemma 5.12 shows that the unique $W(E_{6})$-minimal element of $\Omega_{U^{\prime}}$ is the triple $\{\alpha_{7},\theta_{6},\theta_{7}\}$.
5.3. The invariant cubic polynomial in type $E_{6}$
The sets $U=U(E_{8},8)$ and $U^{\prime}=U(E_{7},7)$ from Theorems 5.9 and 5.13 have representation-theoretic interpretations. We recall that the exceptional Lie algebras of types $E_{6}$, $E_{7}$, and $E_{8}$ have three minuscule representations between them: two $27$-dimensional representations in type $E_{6}$, $L(E_{6},\omega_{1})$ and $L(E_{6},\omega_{6})$, and a $56$-dimensional representation in type $E_{7}$, $L(E_{7},\omega_{7})$, where each $\omega_{i}$ indicates the fundamental weight that appears as the highest weight of the corresponding representation (see, for example, [5, Theorem 5.1.5]). If one restricts the adjoint representation of the simple Lie algebra ${\mathfrak{g}}(E_{8})$ to ${\mathfrak{g}}(E_{7})$, then the set $U$ corresponds to the weights of $L(E_{7},\omega_{7})$ [5, Proposition 10.5.5]. Similarly, restricting the adjoint representation of ${\mathfrak{g}}(E_{7})$ to ${\mathfrak{g}}(E_{6})$ identifies $U^{\prime}$ with the set of weights of $L(E_{6},\omega_{1})$.
The sets $U$ and $U^{\prime}$ also have natural connections to algebraic geometry: the roots in $U$ are in bijection with the 56 exceptional curves on a del Pezzo surface $\mathcal{S}$ of degree 2, and the roots in $U^{\prime}$ are in bijection with the 27 lines on a del Pezzo surface $\mathcal{S}^{\prime}$ of degree 3 [5, Sections 9.3 and 10.1]. Under these bijections, two roots are orthogonal if and only if the intersection number between their corresponding curves is equal to $1$, so the triples of $\Omega_{U^{\prime}}$ are in a bijective correspondence with the tritangent planes of $\mathcal{S}^{\prime}$. The quadruples of $\Omega_{U}$ corresponds bijectively to the what are sometimes called “tetrads” on $\mathcal{S}$; see, for example, [4, Proposition 5.4]. Theorems 5.9 and 5.13 imply that the tritangent planes and tetrads carry natural partial orders induced by the quasiparabolic orders of $\Omega_{U^{\prime}}$ and $\Omega_{U}$, respectively. The existence of these induced orders are new to the best of our knowledge.
In the rest of this subsection, we discuss two applications of the quasiparabolic set $\Omega_{U^{\prime}}$ arising from the QP-triple $T(E_{7},7)$. First, we give explicit descriptions of the generalized Rothe diagrams associated to the elements $\Omega_{U^{\prime}}$ in Theorem 5.17. These descriptions are also new to our knowledge, and we expect they may have applications to Schubert varieties or del Pezzo surfaces. Secondly, we show in Theorem 5.19 that the Pfaffian of $\Omega_{U^{\prime}}$ agrees with the well-known invariant cubic polynomial associated to the minuscule representation $L(E_{6},\omega_{1})$. To state the theorems, we will use the following labelling conventions for the roots of $U$ adapted from the context of del Pezzo surfaces. We note that even though Theorems 5.17 and 5.19 both focus on properties of $\Omega_{U^{\prime}}$, it it more convenient to explain these conventions via the set $U$.
Let $\{\varepsilon_{1},\ldots,\varepsilon_{8}\}$ be an orthonormal basis for ${\mathbb{R}}^{8}$ as usual, and let $\{\alpha_{1},\ldots,\alpha_{8}\}$ be the simple roots of $\Phi=\Phi(E_{8})$, so that each element $\beta\in U$ has the form $\beta=\alpha_{8}+\sum_{i=1}^{7}d_{i}\alpha_{i}$ for some $d_{1},\ldots,d_{7}\geq 0$. In the rest of the paper, we will adopt the labelling convention for the Dynkin diagram of type $E_{7}$ shown in Figure 5, which is compatible with the realization of the root system $\Phi(E_{7})$ inside ${\mathbb{R}}^{8}$ where the simple roots are $\alpha^{\prime}_{i}:=4(\varepsilon_{i}-\varepsilon_{i+1})$ for $1\leq i\leq 6$ and
$\alpha^{\prime}_{7}:=(-2,-2,-2,2,2,2,2,-2).$
In this realization, the $126$ roots of type $E_{7}$ consist of the $56$ coordinate permutations of $4(\varepsilon_{1}-\varepsilon_{2})$ and the $70$ coordinate permutations of $\alpha^{\prime}_{7}$; a root is positive if and only if its rightmost nonzero coefficient is negative. More details about this realization can be found in [5, Example 8.1.7].
11223344556677
Figure 5. New labelling of the Dynkin diagram of type $E_{7}$
For any root $\beta=\alpha_{8}+\sum_{i=1}^{7}d_{i}\alpha_{i}\in U$, we define
$f(\beta)=(-1,-1,-1,-1,-1,-1,3,3)+\sum_{i=1}^{7}d_{i}\alpha^{\prime}_{\tau(i)},$
where $\tau(i)$ refers to the same vertex labeled by $i$ in the Dynkin diagram in the convention of Figure 2 (d). It is known that the map $f$ is a bijection between $U$ and the set of vectors $\{\pm v_{i,j}:1\leq i<j\leq 8\}$, where $v_{i,j}=4(\varepsilon_{i}+\varepsilon_{j})-\sum_{k=1}^{8}\varepsilon_{k}$ for all $i,j$ ([5, Example 8.1.7]). We will write $X_{ij}$ for $v_{i,j}$ and $Y_{ij}$ for $v_{i,j}$. We will also identify each vector $X_{ij}$ or $Y_{ij}$ with the corresponding root in $U$ from now on.
Let $a_{i}=X_{i7}$, $b_{i}=X_{i8}$, and $c_{ij}=c_{ji}=Y_{ij}$ for all $1\leq i\neq j\leq 6$. Routine verification shows that we have
$U^{\prime}=\{a_{i}:1\leq i\leq 6\}\ \dot{\cup}\ \{b_{i}:1\leq i\leq 6\}\ \dot{\cup}\ \{c_{ij}:1\leq i<j\leq 6\},$
and that three elements of $U^{\prime}$} form a triple in $\Omega_{U^{\prime}}$ if and only if they are of the form $a_{i}c_{ij}b_{j}$ for some $1\leq i\neq j\leq 6$ or of the form $c_{i_{1}j_{1}}c_{i_{2}j_{2}}c_{i_{3},j_{3}}$ corresponding to a perfect matching $\{\{i_{1},j_{1}\},\{i_{2},j_{2}\},\{i_{3},j_{3}\}\}$ of $\{1,2,\ldots,6\}$.
Example 5.15.
Recall from Remark 5.14 (ii) that the unique $W(E_{6})$-minimal element in $\Omega_{U^{\prime}}$ is the triple $T_{0}:=\{\alpha_{7},\theta_{6},\theta_{7}\}$ (written in the labelling convention of Figure 2 (d)). Direct translation of notation shows that we have $T_{0}=\{b_{6},c_{16},a_{1}\}$.
Lemma 5.16.
If $\beta$ and $\gamma$ are distinct elements of $U^{\prime}$}, then we have $s_{\gamma}(\beta)<0$ if and only if one of the following conditions holds:
  • (i)
    $\beta=a_{i}$ and $\gamma=a_{j}$ for some $j<i$;
  • (ii)
    $\beta=b_{i}$ and $\gamma=a_{j}$ for some $i\neq j$;
  • (iii)
    $\beta=b_{i}$ and $\gamma=c_{jk}$, where $i$, $j$, and $k$ are distinct;
  • (iv)
    $\beta=c_{ij}$ and $\gamma=c_{kj}$, where $k>i<j$ and $k\neq j$;
  • (v)
    $\beta=c_{ij}$ and $\gamma=c_{il}$ for any $i<j<l$;
  • (vi)
    $\beta=c_{ij}$ and $\gamma=a_{h}$, where $h$, $i$, and $j$ are distinct.
Proof.
This follows from a routine case by case check using the earlier description of the positive roots of type $E_{7}$ in the new coordinate realization. ∎
Theorem 5.17.
The generalized Rothe diagrams of the triples in $\Omega_{U^{\prime}}$ are given as follows.
  • (i)
    If $1<i<j<6$, then the triple $a_{i}c_{ij}b_{j}$ has level 5+ij5+i-j, and we have
    $\text{\rm Res}_{U^{\prime}}(a_{i}c_{ij}b_{j})=\{a_{h}:1\leq h<i\}\ \cup\ \{c_{ik}:j<k\leq 6\}.$
  • (ii)
    If $1<j<i<6$, then the triple $a_{i}c_{ij}b_{j}$ has level 4+ij4+i-j, and we have
    $\text{\rm Res}_{U^{\prime}}(a_{i}c_{ij}b_{j})=\{a_{h}:1\leq h<i\}\ \cup\ \{c_{ki}:j<k\leq 6\text{\ and\ }k\neq i\}.$
  • (iii)
    If $M:=\{\{i_{1},j_{1}\},\{i_{2},j_{2}\},\{i_{3},j_{3}\}\}$ is a perfect matching of $[6]$, then the triple $c_{i_{1}j_{1}}c_{i_{2}j_{2}}c_{i_{3}j_{3}}$ has level $(\ell(\tau)+9)/2$, where $\tau$ is the fixed point free involution associated to $M$ and $\ell(\tau)$ is the length of $\tau$. We have
    $\text{\rm Res}_{U^{\prime}}(c_{i_{1}j_{1}}c_{i_{2}j_{2}}c_{i_{3}j_{3}})=\{a_{i}:1\leq i\leq 6\}\ \cup\ \{c_{pq}:\tau(p)<q\text{\ and\ }\tau(q)<p\}.$
Proof.
For each triple $T\in\Omega_{U^{\prime}}$, we have
$\text{\rm Res}_{U^{\prime}}(T)=\{\beta\in U^{\prime}:s_{\gamma}(\beta)>0\text{ for all }\gamma\in T\}=\bigcap_{\gamma\in T}(U^{\prime}\setminus\{\beta\in U^{\prime}:s_{\gamma}(\beta)<0\}),$
where each set appearing in the intersection can be calculated using Lemma 5.16. The descriptions of the residues follow from the lemma by straightforward computations. Taking cardinalities of the generalized Rothe diagrams yields the levels of the triples in (i) and (ii), and the claim about the level in (iii) follows from Theorem 3.10 (v). ∎
Example 5.18.
Figure 6 depicts the generalized Rothe diagrams of the elements $T_{1}=a_{4}c_{24}b_{2}$ and $T_{2}=c_{13}c_{25}c_{46}$, in the following way. First, we identify the 27 roots of $U^{\prime}$ with the 27 square cells in the array shown in Figure 6 (a) as follows: identify the cells in the top row with $a_{1},a_{2},\ldots,$ and $a_{6}$ from left to right, the cells in the rightmost column with $b_{1},b_{2},\ldots,$ and $b_{6}$ from top to bottom, and the unique cell in the $(a_{j},b_{i})$-position (i.e., the cell in the same row as $b_{i}$ and same column as $a_{j}$) with $c_{ij}=c_{ji}$. If we represent each $T_{i}$ by marking the elements of $T$ with solid black dots, then the residues of $T_{i}$ are the cells marked with a star. For an element of the form $T=a_{i}c_{ij}b_{j}$ where $j<i$, such as $T_{1}$, Theorem 5.17 (ii) implies that $\text{\rm Res}_{U^{\prime}}(T)$ consists of the cells to the left of $a_{i}$, or below $c_{ji}$, or to the left of $b_{i}$ (which is below $b_{j}$ and not in $T$). For an element of the form $T=c_{i_{1}j_{1}}c_{i_{2}j_{2}}c_{i_{3}j_{3}}\in\Omega_{U^{\prime}}$, such as $T_{2}$, one can show using Theorem 5.17 (iii) that $\text{\rm Res}_{U^{\prime}}(T)$ consists of all the six roots in the top row and the roots $c_{ij}$ in the 27-cell grid that are neither attacked from the right nor from below by any of the roots in $T$ or their mirror images across the diagonal consisting of the $(a_{k},b_{k})$-positions for $k\in\{1,2,\ldots,6\}$. We have shaded the cell $b_{i}=b_{4}$ in Figure 6 (a) and the aforementioned mirror images in Figure 6 (b).
$665544332211\bigstar\bigstar\bigstar\bigstar\bigstar\bigstar$
(a) $\text{\rm Res}_{U^{\prime}}(a_{4}c_{24}b_{2})$
$665544332211\bigstar\bigstar\bigstar\bigstar\bigstar\bigstar\bigstar\bigstar$
(b) $\text{\rm Res}_{U^{\prime}}(c_{13}c_{25}c_{46})$
Figure 6. The generalized Rothe diagrams of two elements in $\Omega_{U^{\prime}}$
Theorem 5.17 allows us to compute the quantum Hafnian $\text{\rm QHf}(U^{\prime})$ of $\Omega_{U^{\prime}}$ explicitly. We have
$\text{\rm QHf}(U^{\prime})=$ $\hphantom{+}q^{0}a_{1}c_{16}b_{6}+q^{1}a_{2}c_{26}b_{6}+q^{1}a_{1}c_{15}b_{5}+q^{2}a_{2}c_{25}b_{5}+q^{2}a_{3}c_{36}b_{6}+q^{2}a_{1}c_{14}b_{4}+q^{3}a_{2}c_{24}b_{4}$
$+q^{3}a_{3}c_{35}b_{5}+q^{3}a_{4}c_{46}b_{6}+q^{3}a_{1}c_{13}b_{3}+q^{4}a_{2}c_{23}b_{3}+q^{4}a_{3}c_{34}b_{4}+q^{4}a_{4}c_{45}b_{5}+q^{4}a_{5}c_{56}b_{6}$
$+q^{4}a_{1}c_{12}b_{2}+q^{5}b_{1}c_{12}a_{2}+q^{5}b_{2}c_{23}a_{3}+q^{5}b_{3}c_{34}a_{4}+q^{5}b_{4}c_{45}a_{5}+q^{5}b_{5}c_{56}a_{6}+q^{6}b_{1}c_{13}a_{3}$
$+q^{6}b_{2}c_{24}a_{4}+q^{6}c_{12}c_{34}c_{56}+q^{6}b_{3}c_{35}a_{5}+q^{6}b_{4}c_{46}a_{6}+q^{7}b_{1}c_{14}a_{4}+q^{7}c_{13}c_{24}c_{56}+q^{7}b_{2}c_{25}a_{5}$
$+q^{7}c_{12}c_{35}c_{46}+q^{7}b_{3}c_{36}a_{6}+q^{8}b_{1}c_{15}a_{5}+q^{8}c_{14}c_{23}c_{56}+q^{8}c_{13}c_{25}c_{46}+q^{8}b_{2}c_{26}a_{6}+q^{8}c_{12}c_{36}c_{45}$
$+q^{9}b_{1}c_{16}a_{6}+q^{9}c_{15}c_{23}c_{46}+q^{9}c_{14}c_{25}c_{36}+q^{9}c_{13}c_{26}c_{45}+q^{10}c_{16}c_{23}c_{45}+q^{10}c_{15}c_{24}c_{36}$
$+q^{10}c_{14}c_{26}c_{35}+q^{11}c_{16}c_{24}c_{35}+q^{11}c_{15}c_{26}c_{34}+q^{12}c_{16}c_{25}c_{34}.$
We now show that the Pfaffian $\text{\rm Pf}({U^{\prime}})=\text{\rm QHf}(U^{\prime},-1)$ associated to the QP-triple $T(E_{7},7)$ agrees with the invariant cubic polynomial associated with the Lie algebra (or Chevalley group) of type $E_{6}$ as constructed by Vavilov–Luzgarev–Pevzner [17]. This provides a nonrecursive combinatorial interpretation of the coefficients of the polynomial in terms of the parity of the cardinalities of the associated generalized Rothe diagrams.
Theorem 5.19.
The Pfaffian $\text{\rm Pf}(U^{\prime})$ associated to the QP-triple $T(E_{7},7)$ is the invariant cubic polynomial $F$ constructed in Vavilov–Luzgarev–Pevzner [17, §12].
Proof.
In [17, §12], Vavilov–Luzgarev–Pevzner define the minimal $W(E_{6})$-element $\{b_{6},c_{16},a_{1}\}\in\Omega_{U^{\prime}}$ from Example 5.15 to be the “distinguished triple”, which they denote as $x_{0}=(\lambda_{0},\mu_{0},\nu_{0})$. They define the invariant cubic polynomial $F$ to be the linear combination of the $45$ elements in the $W(E_{6})$-orbit of $x_{0}$ in which the coefficient of any monomial $x_{T}$ is $(-1)^{\ell(w)}$, where $w$ is an element of minimum length such that $w(x_{0})=x_{T}$. As recalled in Section 3, the length of such an element is exactly the level of $x_{T}$ by general properties of quasiparabolic sets, so we have $F=\text{\rm Pf}(\Omega_{U^{\prime}})$. ∎
Remark 5.20.
  • (i)
    Vavilov–Luzgarev–Pevzner note in [17, §12] that if we apply a simple reflection $s_{i}$ on one of the monomials $x_{T}$ in the invariant cubic polynomial, then either $s_{i}$ fixes $x_{T}$, or the coefficients of $x_{T}$ and $s_{i}(x_{T})$ are the negatives of each other. This property also characterizes $\text{\rm Pf}(U^{\prime})$ and provides an alternative proof of Theorem 5.19.
  • (ii)
    The cubic polynomial $F$ is written down explicitly in [17, Table 11]. This makes it possible, after translating notation, to verify Theorem 5.19 by hand.
  • (iii)
    Let $\Gamma$ be the graph whose vertices are the elements of $\Omega_{U^{\prime}}$, and with adjacency given by disjointness. It can be shown that $\Gamma$ is the complement of the graph $U_{4}(3)$ described in [1, §10.17]. In particular, $\Gamma$ is strongly regular with parameters $(45,32,22,24)$.
6. Concluding remarks
Remark 6.1.
The original Rothe diagrams and the Rothe diagrams of fixed point free involutions are useful in the study of Schubert varieties. It would be interesting to know if this also applies to the generalized Rothe diagrams of Theorem 5.17, or to the generalized Rothe diagrams associated to the sets of all positive roots of types $E_{7}$ and $E_{8}$.
Remark 6.2.
In the setting of Section 5.3, the 56 vectors $X_{ij}$ and $Y_{ij}$ arising from the set $U$ can be grouped into 28 pairs $i,j]:=\{X_{ij},Y_{ij}\}$. These pairs correspond naturally to the 28 bitangents to a plane quartic curve, and the 630 quadruples of $\Omega_{U}$ correspond to the 315 “tetrads of bitangents” of the quartic curve [4, Remark 5.5]. It is possible to modify the Pfaffian of $\Omega_{U}$ in a precise way to produce an invariant quartic polynomial known as Dickson’s form (see [4, Definition 5.3] and [2]), and this fact may be viewed as an analogue of Theorem 5.19 for $\Omega_{U}$. We have chosen not to describe the modification here because it requires elaborate rescalings of coefficients for which we have no conceptual explanation. However, we note that the study of geometric objects such as the 630 tetrads corresponding to $\Omega_{U}$ and the 315 tetrads of bitangents has a long history dating back to the works of Cayley and Salmon (see, for example, [4, Remark 5.5] and [15, Section 262]). It would be interesting to know if the set $\Omega_{U}$ has further implications for these classical objects.
Remark 6.3.
For any Coxeter system $(W,S)$ and any subset $I\subseteq S$, the left cosets of the parabolic subgroup $W_{I}\leq W$ form a quasiparabolic $W$-set whose level function $\ell_{I}$ measures the lengths of shortest coset representatives, and using $\ell_{I}$ one can deform the natural action of $W$ on the set $W/W_{I}$ to construct a module for the Iwahori–Hecke algebra $\mathcal{H}$ of $W$. The defining properties of a quasiparabolic set $X$ generalize the properties of $W/W_{I}$ and allows a similar deformation of the $W$-action on $X$ [14, Theorem 7.1]. It would be interesting to use $k$-roots to study the $\mathcal{H}$-modules obtained via the deformation of the quasiparabolic sets $X=\Omega_{U}$ constructed in this paper.
Remark 6.4.
The examples of quasiparabolic sets given in this paper all share some additional structural features that are not required by the definitions or needed for the proofs, which suggests the possible existence of an undiscovered underlying structure. For example, we have no general conceptual explanation for why the Poincaré polynomial $\sum_{R\in\Omega_{U}}q^{\rho_{U}(R)}$ always factorizes as the product of quantum integers. Furthermore, the largest such integer always turns out to be equal to $|U|/\kappa$, where $\kappa$ is the size of a maximal orthogonal subset of $U$. It is often (but not always) the case that $U$ has a partition whose blocks are elements of $\Omega_{U}$, which gives a combinatorial proof that $\kappa$ divides $|U|$. In particular, such a partition always exists for the quasiparabolic sets of the form $T(E_{n},p)$, except in the cases $T(E_{6},2)$, $T(E_{7},2)$, and $T(E_{8},5)$.
Remark 6.5.
One can create a directed acyclic graph (DAG) from a QP-triple $(W,I,U)$ by taking $U$ to be the set of vertices, and installing an edge from $u_{1}$ to $u_{2}$ if $u_{2}-u_{1}$ is a positive root of $W_{I}$. If $T\in\Omega_{U}$ satisfies $\rho(T)=0$, then $T$ is a a kernel of the DAG, meaning that (a) no two elements of $T$ are joined by an edge, and (b) for all $v\in U\backslash T$, there is an edge from $v$ to some element $u\in U$.
We can use the DAG to create a two-player game, as in Example 4.9 (ii), so that the kernel is the kernel in the sense of game theory [18, §51]. The vertices of the graph are the game positions, the directed edges are the legal moves, and playing a move that leads into the kernel constitutes a winning strategy.

### v2

@@
 Abstract.
-Let $U$ be a set of positive roots of type $ADE$, and let $\Omega_{U}$ be the set of all maximum cardinality orthogonal subsets of $U$.
+Let $U$ be a set of positive roots of type $ADE$, and let $\Omega_{U}$ be the set of all maximum-cardinality orthogonal subsets of $U$.
 For each element $R\in\Omega_{U}$, we define a generalized Rothe diagram whose cardinality we call the level, $\rho(R)$, of $R$.
-We define the generalized quantum Hafnian of $U$ to be the generating function of $\rho$, regarded as a $q$-polynomial in $U$.
-Several widely studied algebraic and combinatorial objects arise as special cases of these constructions, and in many cases, $\Omega_{U}$ has the structure of a graded partially ordered set with rank function $\rho$.
-A motivating example of the construction involves a certain set of $k^{2}$ roots in type $D_{2k}$, where the elements of $\Omega_{U}$ correspond to permutations in $S_{k}$, the generalized Rothe diagrams are the traditional Rothe diagrams associated to permutations, the level of a permutation is its length, the generalized quantum Hafnian is the $q$-permanent, and the partial order is the Bruhat order.
-We exhibit many other natural examples of this construction, including one involving perfect matchings, two involving labelled Fano planes, and one involving the invariant cubic form in type $E_{6}$.
+We define the generalized quantum Hafnian of $U$ to be the generating function for $\rho$, regarded as a $q$-polynomial in $U$.
+In this paper, we study a large number of examples of sets of the form $\Omega_{U}$, and we explore their connections with a variety of widely studied algebraic and combinatorial objects.
+One of our motivating examples involves a certain set of $k^{2}$ roots in type $D_{2k}$, where the elements of $\Omega_{U}$ can be identified with permutations in $S_{k}$, the generalized Rothe diagrams are the traditional Rothe diagrams associated to permutations, and the generalized quantum Hafnian is the $q$-permanent.
+More generally, all our examples in types $A$ and $D$ are closely related to perfect matchings and rook configurations, and our examples in type $E$ have applications to del Pezzo surfaces and labellings of Fano planes.
+Each of our examples also gives rise to a matroid, and a large family of our examples is in a natural correspondence with the equal-rank simply-laced symmetric pairs.
 Key words and phrases:
-root system, Weyl group, Rothe diagram, quantum Hafnian, quasiparabolic set
+root system, Weyl group, Rothe diagram, quantum Hafnian, symmetric space
 1991 Mathematics Subject Classification:
@@
 1. Introduction
-Let $U$ be a subset of the positive roots of a Weyl group of type $ADE$, and let $\Omega_{U}$ be the set of orthogonal subsets of $U$ of maximal cardinality.
-For each $R\in\Omega_{U}$, we define a residue of $R$ with respect to $U$ to be an element $\gamma\in U$ such that $s_{\beta}(\gamma)$ is positive for all $\beta\in R$, where $s_{\beta}$ denotes the reflection associated to the vector $\beta$.
-The generalized Rothe diagram of $R$ is the set of residues of $R$ in $U$, which we denote by $\text{\rm Res}_{U}(R)$.
-We call $\rho(R):=|\text{\rm Res}_{U}(R)|$ the level of $R$, and we define the generalized quantum Hafnian of $U$ to be the element of ${\mathbb{Z}}[U][q]$ given by
+Let $U$ be a subset of the positive roots of a Weyl group of type $ADE$, and let $\Omega_{U}$ be the set of orthogonal subsets of $U$ of maximum cardinality.
+For each $R\in\Omega_{U}$, we say that a root $\alpha\in U$ dominates $R$, or that $\alpha$ is $R$-dominating, if $s_{\beta}(\gamma)$ is positive for all $\beta\in R$, where $s_{\beta}$ denotes the reflection associated to $\beta$.
+We define the generalized Rothe diagram, $\mathscr{D}_{U}(R)$, to be the set of all $R$-dominating roots in $U$, and we call $\rho(R):=|\mathscr{D}_{U}(R)|$ the level of $R$.
+We define the generalized quantum Hafnian of $U$ to be the element of ${\mathbb{Z}}[U][q]$ given by
 $\sum_{R\in\Omega_{U}}q^{\rho(R)}\prod_{\beta_{i}\in R}x_{\beta_{i}},$
@@
 Although it is not a priori clear why one might want to study these definitions, which are new to the best of our knowledge, we will show that many choices for $U$ give rise to concise constructions of known algebraic and combinatorial objects, in some cases endowing them with additional structure.
-For example, we will use our methods to associate a generalized Rothe diagram in a canonical way to each of the $45$ terms of the invariant cubic polynomial in type $E_{6}$, which gives a nonrecursive combinatorial interpretation of the coefficients of the polynomial.
-One of the main motivating examples of this phenomenon (Theorem 3.13) is the case where $U=\{\varepsilon_{i}+\varepsilon_{k+j}:1\leq i,j\leq k\}$ is a set of $k^{2}$ roots of type $D_{2k}$, corresponding to the positions $(i,j)$ on a $k\times k$ grid.
-Here, the set $\Omega_{U}$ corresponds to the set of configurations of $k$ nonattacking rooks, or equivalently, the elements of the symmetric group $S_{k}$.
-If we place the origin at the top left, then a vacant position $(a,b)$ is a residue for $R\in\Omega_{U}$ unless it is attacked from above or from its left; see Figure 1 and Remark 3.14 (i).
+For instance, our examples in types $A$ and $D$ are naturally related to perfect matchings and rook configurations (Section 4.2), one family of our examples has a close connection with Cartan’s celebrated classification of Riemannian symmetric spaces and symmetric pairs (Section 4.3), and one of our examples in type $E_{6}$ gives a non-recursive combinatorial interpretation for the 45 terms of a well-known invariant polynomial (Theorem 5.17).
+A key motivating example of our constructions is the case where $U=\{\varepsilon_{i}+\varepsilon_{k+j}:1\leq i,j\leq k\}$ is a set of $k^{2}$ roots of type $D_{2k}$ (Theorem 3.14).
+Here, the set $U$ corresponds to the positions $(i,j)$ on a $k\times k$ grid, and the set $\Omega_{U}$ corresponds to the set of configurations of $k$ nonattacking rooks, or equivalently, the elements of the symmetric group $S_{k}$.
+If we place the origin at the top left and place a dot at each position corresponding to a root in an element $R\in\Omega_{U}$, then a vacant position $(a,b)$ corresponds to an $R$-dominating root unless it is attacked from above or from its left by a dot; see Figure 1 (where the $R$-dominating roots are marked by stars) and Remark 3.15 (i).
 The set of such positions is a well-studied object known as the Rothe diagram of the permutation $\tau$ associated to $R$, and it is a well-known result that the cardinality of the Rothe diagram is equal to the length of $\tau$.
-In this case, the level function is the length of $\tau$, the quantum Hafnian is the $q$-permanent of a $k\times k$ matrix $A$, and specializing $q$ to q-q gives the $q$-determinant of $A$.
+In this case, the level function is the length of $\tau$, the generalized quantum Hafnian is the $q$-permanent of a $k\times k$ matrix $A$, and specializing $q$ to q-q gives the $q$-determinant of $A$.
 $11223344\bigstar\bigstar\bigstar$
-Figure 1. The Rothe diagram of $\underline{3142}\in S_{4}$, with residues marked by stars
-Another motivating example (Theorem 3.10) is the case where $U=\{\varepsilon_{i}+\varepsilon_{j}:1\leq i<j<2k\}$ is a set of $\binom{2k}{2}$ roots of type $D_{2k}$ corresponding to the $2$-subsets of the set $2k]:=\{1,2,\ldots,2k\}$.
-Here, the set $\Omega_{U}$ corresponds to the set of perfect matchings of the set $[2k]$, and a pair $\{a,b\}$ (where $1\leq a<b\leq 2k$) is a residue of $R\in\Omega_{U}$ if and only if we have both $\tau(a)>b$ and $\tau(b)>a$, where $\tau$ is the fixed point free involution corresponding to the perfect matching.
-The residues in this case are the Rothe diagrams of fixed point free involutions as defined by Hamaker–Marberg–Pawlowski [9, (3.3)], the level function is a well-known statistic on perfect matchings, and the quantum Hafnian is precisely the $q$-Hafnian defined by Jing–Zhang [13].
-Specializing $q$ to q-q gives the quantum Pfaffian defined by Strickland [16], and specializing $q$ to +1+1 and to 1-1 gives the Hafnian and Pfaffian of a matrix, respectively.
+Figure 1. The Rothe diagram of $\underline{3142}\in S_{4}$
+Another motivating example is the case where $U=\{\varepsilon_{i}+\varepsilon_{j}:1\leq i<j<2k\}$ is a set of $\binom{2k}{2}$ roots of type $D_{2k}$ corresponding to the $2$-subsets of the set $2k]:=\{1,2,\ldots,2k\}$ (Theorem 3.11).
+Here, the set $\Omega_{U}$ corresponds to the set of perfect matchings of the set $[2k]$ and also to the fixed-point-free involutions of the symmetric group $S_{2k}$.
+For each $R\in\Omega_{U}$, a pair $\{a,b\}$ (where $1\leq a<b\leq 2k$) corresponds to an $R$-dominating root if and only if we have both $\tau(a)>b$ and $\tau(b)>a$, where $\tau$ is the fixed-point-free involution corresponding to $R$.
+The generalized Rothe diagrams in this case recover the Rothe diagrams of fixed-point-free involutions as defined by Hamaker–Marberg–Pawlowski [13, (3.3)], the level function is a well-known statistic on perfect matchings, and the generalized quantum Hafnian is precisely the $q$-Hafnian defined by Jing–Zhang [18].
+Specializing $q$ to q-q gives the quantum Pfaffian defined by Strickland [21], and specializing $q$ to +1+1 and to 1-1 gives the Hafnian and Pfaffian of a suitable matrix, respectively.
 Recall that the Pfaffian, $\text{\rm Pf}(A)$, of a skew-symmetric $2k\times 2k$ matrix $A$ is a polynomial in the entries of $A$ with the property that $(\text{\rm Pf}(A))^{2}=\det(A)$.
@@
 In Section 4, we construct many other examples of sets of roots $U$ whose level function gives $\Omega_{U}$ the structure of a quasiparabolic set.
-Our construction seems to be an unreasonably effective method of producing a quasiparabolic set from a set of roots.
-For root systems of types $D_{2k}$, $E_{7}$, or $E_{8}$, we can form a quasiparabolic set by taking $U={\Phi_{+}}$ to be the set of all positive roots (Theorem 4.3).
-We also discuss an abundance of examples of quasiparabolic sets $\Omega_{U}$ where $U=U(\Gamma,p)$ is a set of roots associated to a suitable simple root $\alpha_{p}$ in a simply laced root system.
-These include examples corresponding to every simple root in every root system of type $A$ or $D$ (Example 4.7, Remark 4.8 (i)), as well as a total of $12$ examples in type $E$ (Example 4.9 (i)).
-The residue sets for the examples of type $A$ and $D$ have interesting combinatorial interpretations in terms of rook configurations, generalizing the aforementioned motivating examples.
-One of the examples in type $E_{7}$, associated to the set $U=U(E_{7},2)$, corresponds in a precise way to the $30$ inequivalent $7$-labellings of the Fano plane, and one example in type $E_{8}$, $U=U(E_{8},2)$, corresponds to the $240$ packings of the projective space $PG(3,2)$.
-In Section 5, we focus on two important examples arising from the construction of Section 4 in type $E$.
-The set $U$ in these cases are given by a certain set $U(E_{8},8)$ of $56$ roots in type $E_{8}$ and a certain set $U(E_{7},7)$ of $27$ roots in type $E_{7}$, both of which appear naturally as the set of weights of a suitable minuscule representation.
-The set $U=U(E_{7},7)$ gives rise to a quasiparabolic set $\Omega_{U}$ whose elements correspond to tritangent planes on a del Pezzo surface of degree 3 and whose Pfaffian gives the well-known invariant cubic polynomial in type $E_{6}$ (Theorem 5.19).
+For root systems of types $D_{2k}$, $E_{7}$, or $E_{8}$, we prove in Section 4.1 that we can construct a quasiparabolic set by taking $U$ to be the set of all positive roots (Theorem 4.3).
+In Section 4.2, we show that by setting $U$ to be the set $U(\Gamma,p)$ of all positive roots (where $\Gamma$ stands for the Dynkin type of the root system) in which a fixed simple root $\alpha_{p}$ appears with odd coefficient, we obtain an abundance of further examples where $\Omega_{U}$ forms a quasiparabolic set (examples 4.7, 4.9, 4.10, 4.11).
+In types $A$ and $D$, the generalized Rothe diagrams associated to the elements of $\Omega_{U(\Gamma,p)}$ have interesting combinatorial interpretations in terms of rook configurations, generalizing the aforementioned motivating examples.
+In type $E$, setting $U=U(E_{7},2)$ and $U=U(E_{8},2)$ yields quasiparabolic sets whose elements are in canonical bijections with labellings of the Fano plane using 7 symbols and 8 symbols, respectively, thus endowing the labellings with natural partial orders induced by the quasiparabolic structure (Remark 4.12).
+Besides perfect matchings, rook configurations, and Fano labellings, the quasiparabolic sets $\Omega_{U}$ constructed in sections 4.1 and 4.2 also enjoy several intriguing connections with certain matroids and Riemannian symmetric spaces.
+In all our examples, every element in $\Omega_{U}$ turns out to be equipped with the structure of a binary matroid that can be used to produce the level generating function of $\Omega_{U}$.
+For the sets $U=U(\Gamma,p)$, the quasiparabolic sets $\Omega_{U}$ correspond to Riemannian spaces whose associated symmetric pair consists of two simply-laced Lie algebras of the same rank, and $U$ coincides with the set of noncompact roots of the symmetric pair.
+For example, for $U=U(E_{8},8)$, every element of $\Omega_{U}$ carries the matroid structure of the affine space $AG(2,2)$ (Example 4.13), and $\Omega_{U}$ is a quasiparabolic set corresponding to the symmetric space of type EIX (Example 4.14).
+The precise correspondences between the sets $\Omega_{U}$, binary matroids, and symmetric spaces are explained in Section 4.3, and we believe that deeper relationships between these objects deserve further study (Remark 6.3).
+In Section 5, we focus on two important examples arising from the constructions of Section 4 in type $E$.
+The sets $U$ in these cases are given by a certain set $U(E_{8},8)$ of $56$ roots in type $E_{8}$ and a certain set $U(E_{7},7)$ of $27$ roots in type $E_{7}$, each of which appears naturally as the set of weights of a suitable minuscule representation.
+The set $U=U(E_{7},7)$ gives rise to a quasiparabolic set $\Omega_{U}$ whose elements correspond to tritangent planes on a del Pezzo surface of degree 3 and whose generalized Pfaffian gives the well-known invariant cubic polynomial in type $E_{6}$ (Theorem 5.19).
 The quasiparabolic order on $\Omega_{U}$ endows the tritangent planes with a natural partial order and yields a non-recursive interpretation of the coefficients of the invariant polynomial, both of which are new to our knowledge.
@@
 In Section 3, we introduce the key definitions and summarize the theory of quasiparabolic sets developed by Rains–Vazirani, and we give details of the two motivating examples.
-Section 4 exhibits many other examples of our construction, including the cases where $U$ is the set of all positive roots of type $E_{7}$, $E_{8}$, or $D_{n}$ for $n$ even.
+Section 4 exhibits many other examples of our construction, as well as their connections to binary matroids and symmetric spaces.
 Section 5 is an in-depth study of the two examples in type $E$ mentioned in the last paragraph.
@@
 We recall the necessary background concerning root systems in this subsection.
-While many definitions in this paper work more generally for all (finite) root systems, we will focus on irreducible simply laced root systems, i.e., on the root systems of types $ADE$.
+While many definitions in this paper work more generally for all (finite) root systems, we will focus on irreducible simply-laced root systems, i.e., on the root systems of types $ADE$.
 We label the vertices in the Dynkin diagrams of root systems as shown in Figure 2 until Section 5.2, where we will switch to a different convention in type $E$ for reasons we will explain.
@@
 (e) $E_{8}$
-Figure 2. Dynkin diagrams of irreducible simply laced Weyl groups
-Let $\Phi=\Phi(\Gamma)$ be a simply laced root system with Dynkin diagram $\Gamma$.
+Figure 2. Dynkin diagrams of irreducible simply-laced Weyl groups
+Let $\Phi=\Phi(\Gamma)$ be a simply-laced root system with Dynkin diagram $\Gamma$.
 The vertices of $\Gamma$ index the simple roots $\Pi=\{\alpha_{i}:i\in\Gamma\}$, and the rank of $\Phi$ is $\text{\rm rank}(\Phi):=\lvert\Pi\rvert$.
@@
 For each subset $I\subseteq S$, the standard parabolic subgroup $W_{I}\leq W$ is the subgroup $\langle I\rangle$ generated by $I$.
-Every root is of the form $\alpha=\sum_{i\in\Gamma}c_{i}\alpha_{i}$ for some integers $c_{i}$ that are either all all nonnegative or all nonpositive.
+Every root is of the form $\alpha=\sum_{i\in\Gamma}c_{i}\alpha_{i}$ for some integers $c_{i}$ that are either all nonnegative or all nonpositive.
 We say $\alpha$ is positive or negative in these cases, respectively.
@@
 Equivalently, a subsystem of $\Phi$ is a nonempty subset that is closed under the operations $\{s_{\alpha}:\alpha\in\Psi\}$.
-The positive roots of $\Phi$ naturally induce a positive system, $\Psi_{+}$, and a simple system, $\Pi_{\Psi}$, in $\Psi$ (in the sense of [10, Section 1.3]; see also [8, Section 2.2]).
-We have $\Psi_{+}=\Psi\cap\Phi_{+}$, and $\Pi_{\Psi}$ can be characterized as the set the elements of $\Psi_{+}$ that cannot be expressed as nonnegative linear combinations of other elements of $\Psi_{+}$.
+The positive roots of $\Phi$ naturally induce a positive system, $\Psi_{+}$, and a simple system, $\Pi_{\Psi}$, in $\Psi$ (in the sense of [15, Section 1.3]; see also [12, Section 2.2]).
+We have $\Psi_{+}=\Psi\cap\Phi_{+}$, and $\Pi_{\Psi}$ can be characterized as the set of elements of $\Psi_{+}$ that cannot be expressed as nonnegative linear combinations of other elements of $\Psi_{+}$.
 The elements of $\Pi_{\Psi}$ may or may not be elements of $\Pi$, although it is true that $\Pi\cap\Psi\subseteq\Pi_{\Psi}$.
@@
 The bilinear form $B$ on ${\mathbb{Z}}\Pi$ is nondegenerate.
-We have $B(\alpha,\beta)\in\{2,1,0,-1,-2\}$ for all roots $\alpha,\beta\in\Phi$ , with $B(\alpha,\beta)=2$ if and only if $\alpha=\beta$ ([10, §6.6]).
-We say two roots $\alpha$ and $\beta$ are orthogonal if $B(\alpha,\beta)=0.$ The roots orthogonal to a fixed root in any root system $\Phi$ naturally form a subsystem, and it is known that when $\Phi$ has type $E_{8}$, $E_{7}$, $E_{6}$, and $D_{n}$ for $n>4$, such a subsystem has type $E_{7}$, $D_{6}$, $A_{5}$, and $A_{1}+D_{n-2}$, respectively, for all roots in $\Phi$ ([7, Theorem 2.5]).
+We have $B(\alpha,\beta)\in\{2,1,0,-1,-2\}$ for all roots $\alpha,\beta\in\Phi$ , with $B(\alpha,\beta)=2$ if and only if $\alpha=\beta$ ([15, §6.6]).
+We say two roots $\alpha$ and $\beta$ are orthogonal if $B(\alpha,\beta)=0.$ The set $\Phi_{\alpha}$ of roots orthogonal to a fixed root $\alpha\in\Phi$ naturally form a subsystem.
 The form $B$ is $W$-invariant in the sense that $B(\alpha,\beta)=B(w(\alpha),w(\beta))$ for all $w\in W$ and all $\alpha,\beta\in\Phi$.
 In particular, the action of $W$ on $\Phi$ preserves orthogonality.
+Remark 2.1.
+It is known that when $\Phi$ has type $E_{8}$, $E_{7}$, $E_{6}$, or $D_{n}$ for $n>4$, for each $\alpha\in\Phi$ the subsystem $\Phi_{\alpha}$ has type $E_{7}$, $D_{6}$, $A_{5}$, or $A_{1}+D_{n-2}$, respectively ([11, Theorem 2.5]).
 The root systems of types $A$, $D$, and $E$ have the following well-known coordinate representations.
@@
 Also note that since orthogonal sets in $R$ are linearly independent in the reflection representation $V$, $k$-roots cannot exist for $k>\dim(V)=\text{\rm rank}(\Phi)$.
-The notion of $k$-roots generalize the notions of $2$-roots and $n$-roots studied in [7, 8], where $n=\text{\rm rank}(\Phi)$; see [7, Remark 3.7] and [8, Section 1].
+The notion of $k$-roots generalizes the notions of $2$-roots and $n$-roots studied in [11, 12], where $n=\text{\rm rank}(\Phi)$; see [11, Remark 3.7] and [12, Section 1].
 Lemma 2.4.
-Let $\Phi$ be an irreducible simply laced root system of rank $n$, and let $W$ be the Weyl group of $\Phi$.
+Let $\Phi$ be an irreducible simply-laced root system of rank $n$, and let $W$ be the Weyl group of $\Phi$.
   • (i)
@@
 Proof of Lemma 2.4.
-Part (i) is well known, and can be proved similarly to [8, Lemma 3.2] by an inductive argument. The first assertion of (ii) is [8, Lemma 3.2], and the second assertion then follows immediately. ∎
+Part (i) is well known, and can be proved similarly to [12, Lemma 3.2] by an inductive argument. The first assertion of (ii) is [12, Lemma 3.2], and the second assertion then follows immediately. ∎
 The following example on the $4$-roots in type $D_{4}$ turns out to be crucial for the general theory of $n$-roots.
@@
 We define $\mathcal{Q}(R)$ to be the set of all $D_{4}$-subsystems $\Psi$ of $\Phi$ such that $\Psi\cap R\in D_{4}(R)$.
-The following proposition expands [8, Proposition 3.3] and details many useful properties of coplanar quadruples.
+The following proposition expands [12, Proposition 3.3] and details many useful properties of coplanar quadruples.
 Proposition 2.10.
@@
 Proof of Proposition 2.10.
-Part (i) follows from [8, Proposition 3.8] and the first paragraph of its proof.
+Part (i) follows from [12, Proposition 3.8] and the first paragraph of its proof.
 To prove (ii), first note that (a) immediately implies (b).
@@
 Remark 2.14.
-Definition 2.12 resembles but differs from [8, Definition 3.9], where the classification of coplanar quadruples into the three types of possible features is characterized in a less concise way.
+Definition 2.13 resembles but differs from [12, Definition 3.9], where the classification of coplanar quadruples into the three types of possible features is characterized in a less concise way.
 However, the definitions are equivalent: in the notation of Example 2.8, each coplanar quadruple is necessarily one of $Q_{1},Q_{2}$ and $Q_{3}$, and both definitions classify $Q_{1}$ as an alignment, $Q_{2}$ as a crossing, and $Q_{3}$ as an alignment.
@@
 Proof.
-Part (i) and (ii) a restatement of [8, Proposition 3.20].
-Part (iii) follows from [8, Lemma 3.18], which states that the features of $R$ form the blocks of a Steiner quadruple system $S(3,4,8)$, meaning that every $3$-element subset of $R$ is contained in a unique feature.
-More precisely, there is a unique Steiner quadruple system on $8$ points ([8, Remark 3.16]) and its blocks are closed under taking complements, so (iii) follows.
+Parts (i) and (ii) are a restatement of [12, Proposition 3.20].
+Part (iii) follows from [12, Lemma 3.18], which states that the features of $R$ form the blocks of a Steiner quadruple system $S(3,4,8)$, meaning that every $3$-element subset of $R$ is contained in a unique feature.
+More precisely, there is a unique Steiner quadruple system on $8$ points ([12, Remark 3.16]) and its blocks are closed under taking complements, so (iii) follows.
 ∎
-3. Residues and quasiparabolic sets
-In Section 3.1, we establish the key new definitions of residues, Rothe diagrams, the level statistic, as well as generalized Hafnians and Pfaffians and their $q$-analogues.
-We also recall the notion of a quasiparabolic set, due to Rains and Vazirani.
-Then, in Section 3.2, we give details of two of the motivating examples of residues: one based on perfect matchings of a set of size $2k$ (Theorem 3.10), and one based on permutations of $k$ objects (Theorem 3.13).
+3. Generalized Rothe diagrams and quasiparabolic sets
+In Section 3.1, we introduce the definitions of generalized Rothe diagrams and several related objects, including generalized quantum Hafnians (Definition 3.2).
+We also recall Rains and Vazirani’s notion of quasiparabolic sets, which provides a common framework for the sets of $k$-roots that we will study in the remaining parts of the paper.
+In Section 3.2, we give details of two of the motivating examples of generalized Rothe diagrams: one based on perfect matchings of a set of size $2k$ (Theorem 3.11), and one based on permutations of $k$ objects (Theorem 3.14).
 3.1. Key definitions
-The following definition introduces the key notion of residues, which applies to all root systems. This notion is new to the best of our knowledge.
+The following two definitions are central to this paper. They make sense for all root systems, and they are new to the best of our knowledge.
+Definition 3.1.
+Let $\Phi$ be an arbitrary root system and let $\alpha,\beta\in\Phi_{+}$. We say that $\alpha$ dominates $\beta$ if $s_{\beta}(\alpha)$ is a positive root.
 Definition 3.2.
@@
   2. (ii)
-    For any orthogonal set $R\in\Omega_{U}$, we define the set of residues of $R$ with respect to $U$ to be
-    $\text{\rm Res}_{U}(R)=\{\gamma\in U:s_{\beta}(\gamma)>0\text{\ for\ all\ }\beta\in R\},$
-    and we will also refer to $\text{\rm Res}_{U}(R)$ as the (generalized) Rothe diagram of $R$ with respect to $U$.
-We define $\rho_{U}(R):=\lvert\text{\rm Res}_{U}(R)\rvert$ to be the level of $R$ with respect to $U$.
+    For any orthogonal set $R\in\Omega_{U}$, we say that a positive root $\alpha\in\Phi_{+}$ dominates $R$, or that $\alpha$ is $R$-dominating, if $\alpha$ dominates every element of $R$.
+We define the generalized Rothe diagram of $R$ with respect to $U$ to be the set
+    $\mathscr{D}_{U}(R)=\{\alpha\in U:\alpha\text{\;dominates\;}R\}=\{\alpha\in U:s_{\beta}(\alpha)>0\text{\ for\ all\ }\beta\in R\}.$
+    We define the level of $R$ with respect to $U$ to be $\rho_{U}(R):=\lvert\mathscr{D}_{U}(R)\rvert$.
   3. (iii)
-    We identify an element $R=\{\beta_{1},\beta_{2},\ldots,\beta_{k}\}\in\Omega_{U}$ with the monomial $\prod_{\beta\in R}x_{\beta}$, where $\{x_{\beta}:\beta\in U\}$ is a set of commuting indeterminates.
-The (generalized) quantum Hafnian of $U$, $\text{\rm QHf}(U)=\text{\rm QHf}(U,q)$, is the element of ${\mathbb{Z}}[U][q]$ given by
+    We associate to each element $R=\{\beta_{1},\beta_{2},\ldots,\beta_{k}\}\in\Omega_{U}$ the monomial $\prod_{\beta\in R}x_{\beta}$, where $\{x_{\beta}:\beta\in U\}$ is a set of commuting indeterminates.
+The generalized quantum Hafnian of $U$, $\text{\rm QHf}(U)=\text{\rm QHf}(U,q)$, is the element of ${\mathbb{Z}}[U][q]$ given by
     $\sum_{R\in\Omega_{U}}q^{\rho(R)}\prod_{\beta_{i}\in R}x_{\beta_{i}}.$
-    The specializations $\text{\rm QPf}(U):=\text{\rm QHf}(U,-q)$, $\text{\rm Hf}(U):=\text{\rm QHf}(U,1)$, and $\text{\rm Pf}(U):=\text{\rm QHf}(U,-1)$ are called the (generalized) quantum Pfaffian, the (generalized) Hafnian, and the (generalized) Pfaffian of $U$, respectively.
+    The specializations $\text{\rm QPf}(U):=\text{\rm QHf}(U,-q)$, $\text{\rm Hf}(U):=\text{\rm QHf}(U,1)$, and $\text{\rm Pf}(U):=\text{\rm QHf}(U,-1)$ are called the generalized quantum Pfaffian, the generalized Hafnian, and the generalized Pfaffian of $U$, respectively.
   4. (iv)
@@
   1. (i)
-    Since $s_{\gamma}(\gamma)=-\gamma<0$ for any positive root $\gamma$ in a root system, it follows that $\text{\rm Res}_{U}(R)$ cannot contain any element of $R$ in the setting of Definition 3.1.
+    Since $s_{\gamma}(\gamma)=-\gamma<0$ for any root $\alpha\in\Phi$, no positive root dominates itself, so the set $\mathscr{D}_{U}(R)$ cannot contain any element of $R$ in the setting of Definition 3.2.
   2. (ii)
@@
 Direct computation shows that
-$\text{\rm Res}_{U}(Q_{1})=\emptyset,\quad\text{\rm Res}_{U}(Q_{2})=\{\theta\}=\{\varepsilon_{1}+\varepsilon_{2}\},\quad\text{and}\quad\text{\rm Res}_{U}(Q_{3})=\{\theta-\alpha_{2},\ \theta\}=\{\varepsilon_{1}+\varepsilon_{2},\ \varepsilon_{1}+\varepsilon_{3}\},$
+$\mathscr{D}_{U}(Q_{1})=\emptyset,\quad\mathscr{D}_{U}(Q_{2})=\{\theta\}=\{\varepsilon_{1}+\varepsilon_{2}\},\quad\text{and}\quad\mathscr{D}_{U}(Q_{3})=\{\theta-\alpha_{2},\ \theta\}=\{\varepsilon_{1}+\varepsilon_{2},\ \varepsilon_{1}+\varepsilon_{3}\},$
 from which it follows that $\rho_{U}(Q_{1})=0$, $\rho_{U}(Q_{2})=1$, and $\rho_{U}(Q_{3})=2$.
@@
 Note that the sets $Q_{i}$ for $1\leq i\leq 3$ are exactly the three coplanar quadruples partitioning $\Phi_{+}$} from Example 2.8, with $Q_{1}$ being an alignment, $Q_{2}$ a crossing, and $Q_{3}$ a nesting.
-The highest root $\theta$ and the second highest root $\theta-\alpha_{2}$ are the only roots that may appear in $\text{\rm Res}_{U}(R)$ for any $R\in\Omega_{U}$.
-Note also that if $\alpha$ is a residue and $\alpha\in\{\theta,\theta-\alpha_{2}\}$, then $\alpha$ has full support with respect to any of the orthogonal bases $Q_{i}=\{\beta_{1},\beta_{2},\beta_{3},\beta_{4}\}$.
-The next definition considers an important special case of Definition 3.1 (ii).
+The highest root $\theta$ and the second highest root $\theta-\alpha_{2}$ are the only roots that may appear in $\mathscr{D}_{U}(R)$ for any $R\in\Omega_{U}$.
+Note also that if a root $\alpha\in\{\theta,\theta-\alpha_{2}\}$ dominates $Q_{i}$ for some $i$, then $\alpha$ has full support with respect to the orthogonal basis $Q_{i}$.
+The next definition considers an important special case of Definition 3.2 (ii), where the reference set $U$ in Definition 3.2 no longer needs to be specified independently.
 Note that by Section 2.1, if $R$ is a $k$-root of a root system $\Phi$, then the set $\Psi:=\text{\rm Span}(R)\cap\Phi$ is a subsystem of $\Phi$, and the positive roots $\Phi_{+}$ of $\Phi$ induce a positive system $\Psi_{+}=\Psi\cap\Phi_{+}=\text{\rm Span}(R)\cap\Phi_{+}$ of $\Psi$.
@@
 Let $\Phi$ be an arbitrary root system, let $R\subseteq\Phi_{+}$ be a $k$-root of $\Phi$, and let $\Psi_{R}=\text{\rm Span}(R)\cap\Phi$ be the subsystem of $\Phi$ induced by $R$.
-We define the set of residues of $R$ to be the set
-$\text{\rm Res}(R)=\text{\rm Res}_{\Psi_{+}}(R),$
-where $\Psi_{+}=\text{\rm Span}(R)\cap\Phi_{+}$ is the positive system of $\Psi$ induced by $\Phi_{+}$.
-We refer to $\text{\rm Res}(R)$ as the (generalized) Rothe diagram of $R$, and we define $\rho(R):=\lvert\text{\rm Res}(R)\rvert$ to be the level of $R$.
+We defined the generalized Rothe diagram of $R$ to be the set
+$\mathscr{D}(R)=\mathscr{D}_{\Psi_{+}}(R),$
+where $\Psi_{+}=\text{\rm Span}(R)\cap\Phi_{+}$ is the positive system of $\Psi$ induced by $\Phi_{+}$. We define $\rho(R):=\lvert\mathscr{D}(R)\rvert$ to be the level of $R$.
 Remark 3.6.
   1. (i)
-    It follows from Definitions 3.1 (ii) and 3.4 that for any $k$-root $R$ of a root system $\Phi$, we have $\text{\rm Res}(R)=\text{\rm Res}_{\Phi_{+}}(R)\cap\text{\rm Span}(R)$.
-In particular, if $R$ is an $n$-root where $n=\text{\rm rank}(\Phi)$, then $\text{\rm Res}(R)=\text{\rm Res}_{\Phi_{+}}(R)$ and $\rho(R)=\rho_{\Phi_{+}}(R)$.
+    It follows from Definitions 3.2 (ii) and 3.5 that for any $k$-root $R$ of a root system $\Phi$, we have $\mathscr{D}(R)=\mathscr{D}_{\Phi_{+}}(R)\cap\text{\rm Span}(R)$.
+In particular, if $R$ is an $n$-root where $n=\text{\rm rank}(\Phi)$, then $\mathscr{D}(R)=\mathscr{D}_{\Phi_{+}}(R)$ and $\rho(R)=\rho_{\Phi_{+}}(R)$.
   2. (ii)
-    In the setting of Example 3.3, we have $\text{\rm Res}(Q_{i})=\text{\rm Res}_{\Phi_{+}}(Q_{i})=\text{\rm Res}_{U}(Q_{i})$ and $\rho(Q_{i})=\rho_{\Phi_{+}}(Q_{i})=\rho_{U}(Q_{i})$ for each $i$.
-Since the calculation of the residue sets in this example depends only on the choice of a simple system for a type $D_{4}$ root system, it follows that whenever $Q$ is a feature in a root system of type $E_{7},E_{8}$ or $D_{n}$ for $n$ even, we have $\rho(Q)=0$ if $Q$ is an alignment, $\rho(Q)=1$ if $Q$ is a crossing, and $\rho(Q)=2$ if $Q$ is a nesting.
-Quasiparabolic sets were introduced by Rains and Vazirani for a general Coxeter system as follows.
+    In the setting of Example 3.4, we have $\mathscr{D}(Q_{i})=\mathscr{D}_{\Phi_{+}}(Q_{i})=\mathscr{D}_{U}(Q_{i})$ and $\rho(Q_{i})=\rho_{\Phi_{+}}(Q_{i})=\rho_{U}(Q_{i})$ for each $i$.
+Since the calculation of $\mathscr{D}(Q_{i})$ in this example depends only on the choice of a simple system for a type $D_{4}$ root system, it follows that whenever $Q$ is a feature in a root system of type $E_{7},E_{8}$ or $D_{n}$ for $n$ even, we have $\rho(Q)=0$ if $Q$ is an alignment, $\rho(Q)=1$ if $Q$ is a crossing, and $\rho(Q)=2$ if $Q$ is a nesting.
+The rest of the paper will study sets of the form $\Omega_{U}$ that are quasiparabolic sets under their associated level functions, in the following sense.
 Definition 3.7.
-[14, Section 2, Section 5] Let $(W,S)$ be a Coxeter system with set of reflections $T$.
+[19, Section 2, Section 5] Let $(W,S)$ be a Coxeter system with set of reflections $T$.
 A scaled $W$-set is a pair $(X,\lambda)$ where $X$ is a $W$-set and $\lambda:X\rightarrow{\mathbb{Z}}$ is a function satisfying $|\lambda(sx)-\lambda(x)|\leq 1$ for all $s\in S$.
@@
 Rains and Vazirani call $\lambda(x)$ the height of $x$ and $\leq_{Q}$ the Bruhat order, but we will refer to them as the level of $x$ and the quasiparabolic order because of the potential for confusion in the context of this paper.
-For any quasiparabolic $W$-set $(X,\lambda)$, it follows from [14, Corollary 2.10] that each $W$-orbit in $X$ contains at most one $W$-minimal element, so if $X$ is a finite and transitive $W$-set then $X$ has a unique $W$-minimal element, $x_{0}$.
-Moreover, when this is the case, the level of any element $x\in X$ equals the minimum length of the elements $w\in W$ such that $w(x_{0})=x$ by [14, Corollary 2.13].
-We may associate a quantum Hafnian and a Poincaré polynomial to a general quasiparabolic set by defining $\text{\rm QHf}(X)=\sum_{x\in X}q^{\rho(R)}x$ and $PS_{X}(q)=\sum_{x\in X}q^{\rho(x)}.$ However, we will not require this level of generality, because our focus for the rest of the paper will be on quasiparabolic sets that arise from sets of orthogonal roots via the following construction.
+For any quasiparabolic $W$-set $(X,\lambda)$, it follows from [19, Corollary 2.10] that each $W$-orbit in $X$ contains at most one $W$-minimal element, so if $X$ is a finite and transitive $W$-set then $X$ has a unique $W$-minimal element, $x_{0}$.
+Moreover, when this is the case, the level of any element $x\in X$ equals the minimum length of the elements $w\in W$ such that $w(x_{0})=x$ by [19, Corollary 2.13].
+We may associate a quantum Hafnian and a Poincaré polynomial to a general quasiparabolic set by defining $\text{\rm QHf}(X)=\sum_{x\in X}q^{\lambda(x)}x$ and $PS_{X}(q)=\sum_{x\in X}q^{\lambda(x)}.$ However, we will not require this level of generality, because our focus for the rest of the paper will be on quasiparabolic sets that arise from sets of orthogonal roots via the following construction.
 Definition 3.8.
@@
 3.2. Motivating examples
-We now present the two motivating examples of residues: one based on perfect matchings, and the other based on permutations.
-The following definition will be helpful in the context of the perfect matching example.
+We now present the two motivating examples of Definition 3.2: one based on perfect matchings, and the other based on permutations.
+The following definition will be helpful in the context of the first example.
 Definition 3.9.
-Let $M=\{\{i_{1},j_{1}\},\ldots,\{i_{k},j_{k}\}\}$ be a perfect matching of the set $[2k]$, where we have $k\geq 2$ and $i_{1}<\ldots<i_{k}$ and $i_{t}<j_{t}$ for all $1\leq t\leq k$.
-Following [12, §2], we define $\sigma_{M}\in S_{2k}$ to be the permutation whose two-line notation is given by
+Let $M=\{\{i_{1},j_{1}\},\ldots,\{i_{k},j_{k}\}\}$ be a perfect matching of the set $[2k]$, where we have $k\geq 2$, $i_{1}<\ldots<i_{k}$, and $i_{t}<j_{t}$ for all $1\leq t\leq k$.
+Following [17, §2], we define $\sigma_{M}\in S_{2k}$ to be the permutation whose two-line notation is given by
 $\sigma_{M}=\begin{bmatrix}1&2&\ldots&2k-1&2k\\ i_{1}&j_{1}&\ldots&i_{k}&j_{k}\end{bmatrix}$.
-We define $\tau_{M}\in S_{2k}$ to be the fixed point free involution whose disjoint cycle notation is given by
+We define $\tau_{M}\in S_{2k}$ to be the fixed-point-free involution whose disjoint cycle notation is given by
 $\tau_{M}=(i_{1},j_{1})(i_{2},j_{2})\ldots(i_{k},j_{k}).$
@@
     For each $M\in\mathcal{M}$, we have
-    $\text{\rm Res}_{U}(R(M))=\{u_{ij}:1\leq i<j\leq 2k,\tau_{M}(i)>j,\text{and }\tau_{M}(j)>i\}.$
+    $\mathscr{D}_{U}(R(M))=\{u_{ij}:1\leq i<j\leq 2k,\tau_{M}(i)>j,\text{and }\tau_{M}(j)>i\}.$
   • (iv)
     For each $M\in\mathcal{M}$, the set
-    $\{(i,j):j<i\text{\ and\ }u_{ji}\in\text{\rm Res}_{U}(R(M))\}=\{(i,j):(j,i)\in D_{M}\}$
-    is the Rothe diagram $\hat{D}_{\tt FPF}(\tau_{M})$ of the fixed point free involution $\tau_{M}$ in the sense of Hamaker–Marberg–Pawlowski [9, (3.3)].
+    $\{(i,j):j<i\text{\ and\ }u_{ji}\in\mathscr{D}_{U}(R(M))\}=\{(i,j):(j,i)\in D_{M}\}$
+    is the Rothe diagram $\hat{D}_{\tt FPF}(\tau_{M})$ of the fixed-point-free involution $\tau_{M}$ in the sense of Hamaker–Marberg–Pawlowski [13, (3.3)].
   • (v)
@@
   • (vi)
-    The generalized quantum Hafnian $\text{\rm QHf}(U)$ is the $q$-Hafnian in the sense of Jing–Zhang [13, Proposition 3.6], and the generalized quantum Pfaffian $\text{\rm QPf}(U)$ is the $q$-Pfaffian in the sense of Strickland [16] and Jing–Zhang [13, Proposition 3.4].
+    The generalized quantum Hafnian $\text{\rm QHf}(U)$ is the $q$-Hafnian in the sense of Jing–Zhang [18, Proposition 3.6], and the generalized quantum Pfaffian $\text{\rm QPf}(U)$ is the $q$-Pfaffian in the sense of Strickland [21] and Jing–Zhang [18, Proposition 3.4].
   • (vii)
-    The Hafnian $\text{\rm Hf}(U)$ is the Hafnian of a $2k\times 2k$ symmetric matrix $(a_{ij})$ satisfying $a_{ij}=u_{ij}$ for all $1\leq i<j\leq 2k$, and the Pfaffian $\text{\rm Pf}(U)$ is the Pfaffian of a $2k\times 2k$ skew-symmetric matrix $(a_{ij})$ satisfying $a_{ij}=u_{ij}$ for all $1\leq i<j\leq 2k$.
+    The generalized Hafnian $\text{\rm Hf}(U)$ is the Hafnian of a $2k\times 2k$ symmetric matrix $(a_{ij})$ satisfying $a_{ij}=u_{ij}$ for all $1\leq i<j\leq 2k$, and the generalized Pfaffian $\text{\rm Pf}(U)$ is the Pfaffian of a $2k\times 2k$ skew-symmetric matrix $(a_{ij})$ satisfying $a_{ij}=u_{ij}$ for all $1\leq i<j\leq 2k$.
   • (viii)
@@
 If $\beta,u_{ij}\in U$ satisfy $s_{\beta}(u_{ij})<0$, then we must either have $\beta=u_{aj}$ for some $a\leq i$, or $\beta=u_{ib}$ for some $b\leq j$.
-It follows that $u_{ij}$ is a residue of $R(M)$ if and only if we have both $\tau_{M}(i)>j$ and $\tau_{M}(j)>i$, proving (iii).
+It follows that $u_{ij}$ dominates $R(M)$ if and only if we have both $\tau_{M}(i)>j$ and $\tau_{M}(j)>i$, proving (iii).
 The set
 $\{(j,i):j<i,\tau(i)>j\text{\ and\ }\tau(j)>i\}$
-is the Rothe diagram $\hat{D}_{\tt FPF}(\tau)$ of the fixed point free involution $\tau$ as defined by Hamaker–Marberg–Pawlowski [9, (3.3)]. Part (iv) follows by applying (iii) to this definition.
-It follows from [9, Proposition 3.6] that the level, $\rho_{U}(R(M))=|\hat{D}_{\tt FPF}(\tau_{M})|$, of $R$ is equal to $(\ell(\tau_{M})-k)/2$.
+is the Rothe diagram $\hat{D}_{\tt FPF}(\tau)$ of the fixed-point-free involution $\tau$ as defined by Hamaker–Marberg–Pawlowski [13, (3.3)]. Part (iv) follows by applying (iii) to this definition.
+It follows from [13, Proposition 3.6] that the level, $\rho_{U}(R(M))=|\hat{D}_{\tt FPF}(\tau_{M})|$, of $R$ is equal to $(\ell(\tau_{M})-k)/2$.
 Lemma 3.10 (iii) implies that $\rho_{U}(R(M))=\ell(\sigma_{M}$), and this completes the proof of (v).
-Rains–Vazirani [14, §4] proved that the function $(\ell(\tau_{M})-k)/2$ is a quasiparabolic level function for the set of fixed point free involutions of $S_{2k}$, regarded as an $S_{2k}$-set in the natural way.
+Rains–Vazirani [19, §4] proved that the function $(\ell(\tau_{M})-k)/2$ is a quasiparabolic level function for the set of fixed-point-free involutions of $S_{2k}$, regarded as an $S_{2k}$-set in the natural way.
 Combining this result with part (v) proves (i).
@@
 where the sum is over all perfect matchings of $[2k]$.
-This agrees with Jing–Zhang’s construction of the quantum Hafnian from [13, Proposition 3.6] All the assertions of (vi) and (vii) now follow from the relevant definitions.
-Part (viii) is a standard result that has been rediscovered several times; see [8, Proposition 7.1]. ∎
+This agrees with Jing–Zhang’s construction of the quantum Hafnian from [18, Proposition 3.6] All the assertions of (vi) and (vii) now follow from the relevant definitions.
+Part (viii) is a standard result that has been rediscovered several times; see [12, Proposition 7.1]. ∎
 Example 3.12.
@@
 Remark 3.13.
-In certain cases, the product of the residues of a perfect matching, when regarded as a polynomial in the $\varepsilon_{i}$, has an interpretation in terms of Schubert polynomials; see [9, Theorem 1.3].
+In certain cases, the product of the elements in $\mathscr{D}_{U}(R(M))$ associated to a perfect matching $M$, when regarded as a polynomial in the $\varepsilon_{i}$, has an interpretation in terms of Schubert polynomials; see [13, Theorem 1.3].
 Theorem 3.14.
@@
     For each $\pi\in S_{k}$, we have
-    $\text{\rm Res}_{U}(R(\pi))=\{v_{ij}:i,j\in[k],\pi(i)>j,\text{and }\pi^{-1}(j)>i\}.$
+    $\mathscr{D}_{U}(R(\pi))=\{v_{ij}:i,j\in[k],\pi(i)>j,\text{and }\pi^{-1}(j)>i\}.$
   • (iv)
     The set
-    $\{(i,j):v_{ij}\in\text{\rm Res}_{U}(R(\pi))\}$
+    $\{(i,j):v_{ij}\in\mathscr{D}_{U}(R(\pi))\}$
     is the Rothe diagram $D(\pi)$ of the permutation $\pi$.
@@
   • (vi)
-    The generalized quantum Hafnian $\text{\rm QHf}(U)$ is the $q$-permanent of the matrix $A=(v_{ij})$ in the sense of [12, (2.7)], and the generalized quantum Pfaffian $\text{\rm QHf}(U)$ is the quantum row permanent $\text{\rm per}_{q}(A)$ of $A$ in the sense of [13, (3.18)].
+    The generalized quantum Hafnian $\text{\rm QHf}(U)$ is the $q$-permanent of the matrix $A=(v_{ij})$ in the sense of [17, (2.7)], and the generalized quantum Pfaffian $\text{\rm QHf}(U)$ is the quantum row permanent $\text{\rm per}_{q}(A)$ of $A$ in the sense of [18, (3.18)].
   • (vii)
-    The Hafnian $\text{\rm Hf}(U)$ and the Pfaffian $\text{\rm Pf}(U)$ are the permanent $\text{\rm per}(A)$, and the determinant $\det(A)$, respectively.
+    The generalized Hafnian $\text{\rm Hf}(U)$ and the generalized Pfaffian $\text{\rm Pf}(U)$ are the permanent $\text{\rm per}(A)$, and the determinant $\det(A)$, respectively.
   • (viii)
@@
 Parts (ii) and (iii) follow by restriction from Theorem 3.11 (ii) and (iii), with each $k$-root $R(\pi)$ defined in (ii) corresponding to the perfect matching $M\{\{i,\pi(i)+k\}:i\in[k]\}$.
-Part (iv) then follows from the definition of the Rothe diagram $D(\pi)$ [9, (2.7)].
+Part (iv) then follows from the definition of the Rothe diagram $D(\pi)$ [13, (2.7)].
 Part (v), which is well known, follows from the fact that there is a bijection between the inversions of $\pi$ and the Rothe diagram of $\pi$ given by $(i,j)\mapsto(i,\pi(j))$.
@@
 $(w_{1},w_{2})\cdot w=w_{1}w(w_{2})^{-1}.$
-By [14, Theorem 3.1], this makes $S_{k}$ into a quasiparabolic set with height function $\ell$, proving (i).
+By [19, Theorem 3.1], this makes $S_{k}$ into a quasiparabolic set with height function $\ell$, proving (i).
 Parts (vi) and (vii) follow from the relevant definitions, and (viii) follows from the well-known identity $\sum_{\pi\in S_{k}}q^{\ell(\pi)}=[k]_{q}!$. ∎
@@
     It is sometimes helpful to identify $U$ as the positions on a $k\times k$ board and $\Omega_{U}$ with configurations of $k$ nonattacking rooks on the board, with each root $v_{ij}=\varepsilon_{i}\pm\varepsilon_{j+k}$ corresponding to the $(i,j)$-th position.
-A position on the board then corresponds to a residue if and only if it is vacant and not attacked either from the left or from above.
-Plotting a $k$-root and its residues on the board yields the original Rothe diagram of the permutation correspoding to the rook configuration; see Figure 1.
+A position on the board then dominates a $k$-root if and only if it is vacant and not attacked either from the left or from above.
+Plotting a $k$-root $R$ and the set $\mathscr{D}_{U}(R)$ on the board yields the original Rothe diagram of the permutation corresponding to the rook configuration; see Figure 1.
   2. (ii)
-    Instead of regarding $\Omega_{U}$ as an $S_{k}\times S_{k}$-set, we may regard it as an $(S_{k},S_{k})$-biset, as in the proof of part (v) of the theorem, and from this perspective, $\Omega_{U}$ has a left $S_{k}$-action permuting the rows, and a right $S_{k}$-action permuting the columns.
-The actions commute with each other, and $\Omega_{U}$ is a quasiparaboic $S_{k}$-set under each action.
+    Instead of regarding $\Omega_{U}$ as an $S_{k}\times S_{k}$-set, we may regard it as an $(S_{k},S_{k})$-biset, as in the proof of part (v) of the theorem.
+From this perspective, $\Omega_{U}$ has a left $S_{k}$-action permuting the rows, and a right $S_{k}$-action permuting the columns.
+The actions commute with each other, and $\Omega_{U}$ is a quasiparabolic $S_{k}$-set under each action.
 4. Further examples of QP-triples
-In this section, we give details of many other examples of QP-triples.
-Section 4.1 studies the root systems of rank $n$ where $n$-roots exist, and we show that for these root systems we can form a QP-triple by taking $U$ to be the entire set of positive roots (Theorem 4.3).
-In Section 4.2, we show that it is often possible to construct examples of QP-triples by selecting those positive roots in which particular simple roots appear with odd coefficients.
-4.1. Orthogonal bases of roots
-Recall from Lemma 2.3 that a Weyl group $W$ of rank $n$ and type $ADE$ has a set of $n$ orthogonal roots if and only if $W$ has type $E_{7}$, $E_{8}$, or $D_{n}$ for $n$ even, and we will show that $(W,S,{\Phi_{+}})$ is a QP-triple in these cases.
+We give details of two families of examples of QP-triples in this section.
+The first family is associated to the root systems of rank $n$ where $n$-roots exist, and we show in Section 4.1 that for these root systems we can form a QP-triple by taking $U$ to be the entire set of positive roots.
+In Section 4.2, we show that for each root system of simply-laced type $\Gamma$, it is often possible to construct a QP-triple $T(\Gamma,p)$ by taking $U$ to be the set of all positive roots in which a particular simple root $\alpha_{p}$ appears with odd coefficients (Definition 4.6).
+This construction furnishes a large number of QP-triples that form our second family of examples, and we will explain their rich connections to the combinatorics of perfect matchings, rook configurations, and Fano plane labellings.
+Two examples from this family, $T(E_{8},8)$ and $T(E_{7},7)$, also have natural connections to minuscule representations and del Pezzo surfaces, and a technical study of these two examples will be given in Section 5.
+In Section 4.3, we explain how each QP-triple from sections 4.1 and 4.2 gives rise to a matroid, as well as how this matroid be used to predict the degrees of the QP-triple.
+We also explain how the examples $T(\Gamma,p)$ in the second family relate to symmetric spaces and symmetric pairs.
+4.1. Orthogonal bases
+Recall from Lemma 2.4 that a Weyl group $W$ of rank $n$ and type $ADE$ has a set of $n$ orthogonal roots if and only if $W$ has type $E_{7}$, $E_{8}$, or $D_{n}$ for $n$ even.
+We will show that $(W,S,{\Phi_{+}})$ is a QP-triple in these cases.
 These examples are particularly useful because many other examples of QP-triples can be constructed from them by restriction.
-The next result develops some key properties of residues of $n$-roots in these cases.
-Recall from Remark 3.5 (i) that for each $n$-root $R$ of $\Phi$, we have $\text{\rm Res}(R)=\text{\rm Res}_{\Phi_{+}}(R)$ and $\rho(R)=\rho_{\Phi_{+}}(R)$.
+The next result develops some key properties of the generalized Rothe diagrams of $n$-roots in these cases.
+Recall from Remark 3.6 (i) that for each $n$-root $R$ of $\Phi$, we have $\mathscr{D}(R)=\mathscr{D}_{\Phi_{+}}(R)$ and $\rho(R)=\rho_{\Phi_{+}}(R)$.
 Proposition 4.1.
@@
   • (i)
-    The set of residues of $R$ can be expressed as the disjoint union
-    $\text{\rm Res}(R)=\dot{\bigcup}_{Q\in D_{4}(R)}\text{\rm Res}(Q).$
+    The generalized Rothe diagram $\mathscr{D}(R)$ can be expressed as the disjoint union
+    $\mathscr{D}(R)=\dot{\bigcup}_{Q\in D_{4}(R)}\mathscr{D}(Q).$
   • (ii)
     If $\Phi$ is of type $E_{8}$, then with respect to each fixed feature $Q\in D_{4}(R)$, we have
-    $\text{\rm Res}(R)=\text{\rm Res}(Q)\quad\dot{\cup}\quad\dot{\bigcup}_{Q\subset H\in D_{6}(R)}(\text{\rm Res}(H)\backslash\text{\rm Res}(Q))\quad\dot{\cup}\quad\text{\rm Res}(R\backslash Q).$
+    $\mathscr{D}(R)=\mathscr{D}(Q)\quad\dot{\cup}\quad\left(\dot{\bigcup}_{Q\subset H\in D_{6}(R)}\mathscr{D}(H)\backslash\mathscr{D}(Q)\right)\quad\dot{\cup}\quad\mathscr{D}(R\backslash Q).$
 Proof.
-Remark 3.2 (i) and Proposition 2.9 (ii) imply that every $\gamma\in\text{\rm Res}(R)$ lies in the support of a unique $Q\in D_{4}(R)$.
-It follows that $\gamma\in\text{\rm Res}(Q)$ in this case, which proves $\text{\rm Res}(R)\subseteq\dot{\bigcup}_{Q\in D_{4}(R)}\text{\rm Res}(Q)$.
-Conversely, any element $\gamma\in\text{\rm Res}(Q)$ is fixed by all reflections $s_{\beta}$ such that $\Q\beta\in R\backslash$ Q, which implies that $\gamma\in\text{\rm Res}(R)$ and establishes the reverse containment.
+Remark 3.3 (i) and Proposition 2.10 (ii) imply that every $\gamma\in\mathscr{D}(R)$ lies in the support of a unique $Q\in D_{4}(R)$.
+It follows that $\gamma\in\mathscr{D}(Q)$ in this case, which proves $\mathscr{D}(R)\subseteq\dot{\bigcup}_{Q\in D_{4}(R)}\mathscr{D}(Q)$.
+Conversely, any element $\gamma\in\mathscr{D}(Q)$ is fixed by all reflections $s_{\beta}$ such that $\Q\beta\in R\backslash$ Q, which implies that $\gamma\in\mathscr{D}(R)$ and establishes the reverse containment.
 The disjointness assertion follows from Proposition 2.10 (iii), which completes the proof of (i).
@@
 Applying part (i) to the type $D_{6}$ subsystem $\Xi=\text{\rm Span}(H)\cap\Phi$ yields
-$\text{\rm Res}(H)=\text{\rm Res}(Q)\ \dot{\cup}\ \text{\rm Res}(Q^{\prime})\ \dot{\cup}\ \text{\rm Res}(Q^{\prime\prime})$
+$\mathscr{D}(H)=\mathscr{D}(Q)\ \dot{\cup}\ \mathscr{D}(Q^{\prime})\ \dot{\cup}\ \mathscr{D}(Q^{\prime\prime})$
 and thus
-$\text{\rm Res}(H)\backslash\text{\rm Res}(Q)=\text{\rm Res}(Q^{\prime})\ \dot{\cup}\ \text{\rm Res}(Q^{\prime\prime}).$
+$\mathscr{D}(H)\backslash\mathscr{D}(Q)=\mathscr{D}(Q^{\prime})\ \dot{\cup}\ \mathscr{D}(Q^{\prime\prime}).$
 The conclusions of (ii) now follow from (i). ∎
@@
 By Lemma 2.4 and the paragraph preceding it, the set $\Omega_{\Phi_{+}}$ consists of the positive $n$-roots of $\Phi$, and the standard action of $W_{S}=W$ on $\Omega_{\Phi_{+}}$ is transitive.
-It is known [8, Theorem 4.5] that $\Omega_{\Phi_{+}}$ is a quasiparabolic set for $W$ under the standard action and level function $C+2N$, where $C$ and $N$ count the numbers of crossings and nestings in each $n$-root.
+It is known [12, Theorem 4.5] that $\Omega_{\Phi_{+}}$ is a quasiparabolic set for $W$ under the standard action and level function $C+2N$, where $C$ and $N$ count the numbers of crossings and nestings in each $n$-root.
 Since $C(R)+2N(R)=\rho(R)=\rho_{\Phi_{+}}(R)$ for all $R\in\Omega_{\Phi_{+}}$ by Corollary 4.2, it follows that $(W,S,\Phi_{+})$ is a transitive QP-triple, which completes the proof.
 ∎
-Corollary 4.4.
+Remark 4.4.
+  • (i)
+    The Poincaré polynomial of $PS_{\Phi_{+}}(q)$ factors into quantum integers, with the multiset of degrees being $\{3,5,\ldots,2k-1\}$ in type $D_{2k}$, $\{3,5,9\}$ in type $E_{7}$, and $\{3,5,9,15\}$ in type $E_{8}$ ([12, Proposition 7.1 (i)]).
+We will explain how the degrees can be obtained combinatorially for type $D_{2k}$ in Example 4.7, and we will give a conceptual proof of the degrees in types $E_{7}$ and $E_{8}$, which were previously obtained only computationally, in the forthcoming paper [10].
+In addition, we will also explain an algorithm for computing the degrees by using a matroid associated with $\Omega_{\Phi_{+}}$ in Section 4.3; see also Example 4.13.
+  • (ii)
+    Table 1 summarizes the key numerical quantities associated with the examples of Theorem 4.3, where $\kappa=n$ is the common cardinality of the sets in $\Omega_{U}$.
+The last column of the table concerns the matroid we can naturally associate to $\Omega_{\Phi_{+}}$, whose construction we will elaborate on in Section 4.3.
+Note that the cardinality of the set $\Omega_{U}$ is simply the product of the degrees by Definition 3.2 (iv), so $\lvert\Omega_{U}\rvert$ equals $(2k-1)!!,3\cdot 5\cdot 9=135$, and $3\cdot 5\cdot 9\cdot 15=2025$ in types $D_{2k},E_{7}$, and $E_{8}$, respectively.
+    $\Gamma$ $\kappa$ $|U|$ degrees matroid
+    $D_{2k}$ $2k$ 2k(2k1)2k(2k-1) $\{3,5,7,\ldots,2k-1\}$ $U^{(2)}_{k-1,k}$
+    $E_{7}$ 77 $63$ $\{3,5,9\}$ PG(2,2)
+    $E_{8}$ 88 $120$ $\{3,5,9,15\}$ AG(3,2)
+    Table 1. QP-triples of the form $(W,S,{\Phi_{+}}_{+})$
+  • (iii)
+    In type $D_{2k}$, the quasiparabolic set $\Omega_{\Phi_{+}}$ for the Weyl group $W=W(D_{2k})$ is essentially the same as the quasiparabolic set $\Omega_{U}$ for the parabolic subgroup $W_{I}\cong S_{2k}$ from Theorem 3.11 (where $I=S(D_{2k}\setminus\{s_{2k}\}$), in the sense that there is a natural bijection from $\Omega_{U}$ to $\Omega_{\Phi_{+}}$ that is $W_{I}$-equivariant and preserves level, namely, the map
+    $\phi:\Omega_{U}\rightarrow\Omega_{\Phi_{+}},\;\{\varepsilon_{i_{1}}+\varepsilon_{j_{1}},\ldots,\varepsilon_{i_{k}}+\varepsilon_{j_{k}}\}\mapsto\{\varepsilon_{i_{1}}+\varepsilon_{j_{1}},\varepsilon_{i_{1}}-\varepsilon_{j_{1}},\ldots,\varepsilon_{i_{k}}+\varepsilon_{j_{k}},\varepsilon_{i_{k}}-\varepsilon_{j_{k}}\}.$
+  • (iv)
+    The unique $W$-minimal element in type $E_{8}$ is the set
+    $R=\{\alpha_{2},\alpha_{3},\alpha_{5},\alpha_{7},\theta_{4},\theta_{6},\theta_{7},\theta_{8}\},$
+    where $\theta_{4}$, $\theta_{6}$, $\theta_{7}$, and $\theta_{8}$ are the highest roots in the standard parabolic subsystems of types $D_{4}$, $D_{6}$, $E_{7}$, and $E_{8}$ of $\Phi$, respectively.
+This follows from [12, §6.3], and can also be checked directly using the definition of $\mathscr{D}(R)$.
+Since the lowest-degree term of $PS_{\Phi_{+}}(q)$ is the constant 1 by (ii), we have $\rho(R)=0$.
+We close this subsection with the following lemma concerning the action of reflections on $\Omega_{\Phi_{+}}$. It will be used frequently in Section 5.
+Lemma 4.5.
 Maintain the assumptions of Theorem 4.3.
@@
 By that remark and Remark 3.6 (ii), if $\rho(s_{\alpha}Q)>\rho(Q)$ then we have (a) $Q$ is an alignment and $s_{\alpha}Q$ is a crossing, or (b) $Q$ is a crossing and $s_{\alpha}Q$ is a nesting, or (c) $Q$ is an alignment and $s_{\alpha}Q$ is a nesting.
-In these cases, the value of $\rho(s_{\alpha}(Q))-\rho(Q)$ is 1, 1, and 2, respectively, and $d$ is a nonnegative even number by parts (ii), (iii), and (iv) of [8, Proposition 4.7], respectively.
+In these cases, the value of $\rho(s_{\alpha}(Q))-\rho(Q)$ is 1, 1, and 2, respectively, and $d$ is a nonnegative even number by parts (ii), (iii), and (iv) of [12, Proposition 4.7], respectively.
 ∎
-Remark 4.5.
-The paper [8] details many interesting properties of the quasiparabolic sets produced by Theorem 4.3. We highlight some of them that are directly related to this paper below.
-  • (i)
-    In type $D_{2k}$, the quasiparabolic set $\Omega_{\Phi_{+}}$ for the Weyl group $W=W(D_{2k})$ is essentially the same as the quasiparabolic set $\Omega_{U}$ for the parabolic subgroup $W_{I}\cong S_{2k}$ from Theorem 3.10 (where $I=S(D_{2k}\setminus\{s_{2k}\}$), in the sense that there is a natural bijection from $\Omega_{U}$ to $\Omega_{\Phi_{+}}$ that is $W_{I}$-equivariant and preserves level, namely, the map
-    $\phi:\Omega_{U}\rightarrow\Omega_{\Phi_{+}},\;\{\varepsilon_{i_{1}}+\varepsilon_{j_{1}},\ldots,\varepsilon_{i_{k}}+\varepsilon_{j_{k}}\}\mapsto\{\varepsilon_{i_{1}}+\varepsilon_{j_{1}},\varepsilon_{i_{1}}-\varepsilon_{j_{1}},\ldots,\varepsilon_{i_{k}}+\varepsilon_{j_{k}},\varepsilon_{i_{k}}-\varepsilon_{j_{k}}\}.$
-  • (ii)
-    The Poincaré polynomial of $PS_{\Phi_{+}}(q)$ factors into quantum integers, with the multiset of degrees being $\{3,5,\ldots,2k-1\}$ in type $D_{2k}$, $\{3,5,9\}$ in type $E_{7}$, and $\{3,5,9,15\}$ in type $E_{8}$ ([8, Proposition 7.1 (i)]).
-In particular, the cardinality of $\Omega_{\Phi_{+}}$ is $3\cdot 5\cdot 9=135$ in type $E_{7}$ and $3\cdot 5\cdot 9\cdot 15=2025$ in type $E_{8}$.
-  • (iii)
-    The unique $W$-minimal element in type $E_{8}$ is the set
-    $R=\{\alpha_{2},\alpha_{3},\alpha_{5},\alpha_{7},\theta_{4},\theta_{6},\theta_{7},\theta_{8}\},$
-    where $\theta_{4}$, $\theta_{6}$, $\theta_{7}$, and $\theta_{8}$ are the highest roots in the standard parabolic subsystems of types $D_{4}$, $D_{6}$, $E_{7}$, and $E_{8}$ of $\Phi$, respectively.
-This follows from [8, §6.3], and can also be checked directly using the definition of residues.
-Since the lowest-degree term of $PS_{\Phi_{+}}(q)$ is the constant 1 by (ii), the $8$-root $R$ has no residues with respect to $\Phi_{+}$.
 4.2. An effective construction
-The constructions of the following definition turn out to provide extremely effective methods for constructing QP-triples from the root systems of type $ADE$, and many of the resulting QP-triples provide constructions of known objects, in some cases providing extra insight into their structure.
-We will record the examples without giving any detailed proofs in this section, but we note that the examples in types $A$ and $D$ can all be interpreted in terms of the combinatorics of rook configurations, and we will study the examples $T(E_{8},8)$ and $T(E_{7},7)$ in full detail in Section 5.
+The following definition produces a highly effective method for constructing QP-triples from simply-laced root systems.
+As our examples will show, many of the QP-triples constructed using the definition recover known objects, in some cases providing extra insight into their structure.
+The sets $U(\Gamma,p)$ in the definition correspond to symmetric spaces in a certain precise way that we will explain in Section 4.3.
 Definition 4.6.
@@
 […]
-Example 4.7.
-  • (i)
-    The triple $(W,I,U)$ appearing in Theorem 3.10 is $T(D_{2k},2k)$.
-In type $D_{n}$, a simple root can only appear with coefficient $\pm 2$, $\pm 1$, or 0, so $U(D_{2k},2k)$ is the set of roots in which $\alpha_{2k}$ appears with coefficient $1$.
-  • (ii)
-    The triples $T(D_{2k},2k-1),T(D_{2k-1},2k-1),$ and $T(D_{2k-1},2k-2)\}$ are also QP-triples.
-    The above three triples all give rise to essentially the same quasiparolic set as $T(D_{2k},2k)$, in the sense that their associated quasiparabolic sets $\Omega_{U(\Gamma,p)}$ all admit canonical level-preserving bijections to the quasiparabolic set $\Omega(D_{2k},2k)$ from (i).
+Note that when we expand a root in $\Phi$ into a linear combination of simple roots, the coefficient of each simple root lies in the set $\{\pm 1,0\}$ if $\Phi$ is of type $A$ and in the set $\{\pm 2,\pm 1,0\}$ if $\Phi$ has type $D_{n}$.
+Thus, in these two types, the condition in Definition 4.6 that $\alpha_{p}$ appears with odd positive coefficient is equivalent to the condition that $\alpha_{p}$ appears with coefficient 1.
+In the rest of this subsection, we will describe all QP-triples of the form $T(\Gamma,p)$ where $\Gamma$ is of type $ADE$.
+We will explain how to interpret the examples in types $A$ and $D$ using the combinatorics of rook configurations, and we will study the examples $T(E_{8},8)$ and $T(E_{7},7)$ in full detail in Section 5.
+The key numerical quantities associated with the QP-triples are summarized in Table 2 below, where $U=U(\Gamma,p)$, $\kappa$ stands for the maximum cardinality of orthogonal subsets of $U$, $k$ stands for an integer larger than $1$, $n$ denotes the rank of $\Phi$, and $\lfloor\cdot\rfloor$ denotes the floor function.
+The Poincaré polynomial factors as a multiplicity-free product of quantum integers in all cases, so each example has a well-defined, multiplicity-free set of degrees; we will express this set in the notation
+$a]_{b}:=\{a,a-1,a-2,...,a-b+1\}$
+in some rows of the table. Recall that the cardinality of the quasiparabolic set $\Omega_{U}$ can be deduced from the degrees, as the product of the degrees, by Definition 3.2 (iv).
+$\Gamma$ $p$ $\kappa$ $\lvert U\rvert$ degrees matroid
+$A_{2k-1}$ $k$ $k$ $k^{2}$ $p]_{p-1}$ $U_{p-1,p}$
+$A_{n}$ $n]\setminus\{(n+1)/2\}$ $m:=\min(p,n+1-p)$ m(n+1m)m(n+1-m) $n+1-m]_{m}$ $U_{m,m}$
+$D_{2k}$ $k$ $2k$ $2k^{2}$ $p]_{p-1}$ $U^{(2)}_{p-1,p}$
+$D_{n}$ $1,2,\ldots,\lfloor n/2\rfloor$ $2p$ 2p(np)2p(n-p) $n-p]_{p}$ $U^{(2)}_{p,p}$
+$D_{2k}$ 2k1,2k2k-1,2k $k$ k(2k1)k(2k-1) $\{3,5,\ldots,2k-1\}$ $U_{k-1,k}$
+$D_{2k+1}$ $2k,2k+1$ $k$ $k(2k+1)$ $\{3,5,\ldots,2k+1\}$ $U_{k,k}$
+$E_{6}$ 22 44 2020 $\{2,3,5\}$ $AG(2,2)$
+$E_{6}$ $1,6$ 22 $16$ $\{5,8\}$ $AG(1,2)$
+$E_{7}$ $2,5$ 77 $35$ $\{2,3,5\}$ $PG(2,2)$
+$E_{7}$ 11 44 $32$ $\{3,5,8\}$ $AG(2,2)$
+$E_{7}$ 77 33 $27$ $\{5,9\}$ $PG(1,2)$
+$E_{8}$ $1,2,5,6$ 88 $64$ $\{2,3,5,8\}$ $AG(3,2)$
+$E_{8}$ 88 44 $56$ $\{5,9,14\}$ $AG(2,2)$
+Table 2. QP-triples $T(\Gamma,p)$ of type ADEADE
+Example 4.7 (Type $A$).
+The triple $T(A_{n},p)$ is a QP-triple for all integers $n\geq 1$ and $1\leq p\leq n$.
+For such a triple, the set $U=U(A_{n},p)$ consists of the roots of the form $\varepsilon_{i}-\varepsilon_{j}$ where $1\leq i\leq p<j\leq n+1$, so we may identify $U$ with the positions on a $p\times(n+1-p)$ board with rows labeled by 1 to $p$ from top to bottom and columns labelled by $p+1$ to n+1pn+1-p from left to right.
+Each element of $R\in\Omega_{U(A_{n},p)}$ consists of $m:=\min(p,n+1-p)$ orthogonal roots, and if we view $R$ as an involution in the natural way, as the product $w_{R}\in S_{n+1}$ of the 2-cycles $(i,j)$ for which $\varepsilon_{i}-\varepsilon_{j}\in R$, then a root $\varepsilon_{a}-\varepsilon_{b}\in U(A_{n},p)$ dominates $R$ if and only if we have both $w_{R}(a)<b$ and $w_{R}(b)>a$.
+Equivalently, a position on the board lies in $\mathscr{D}_{U}(R)$ if and only if it is not attacked by a rook to its right or above it in the rook configuration corresponding to $R$, as shown in Figure 4, where the elements of $\mathscr{D}_{U}(R)$ are marked by stars as in Figure 1.
+Thus, the generalized Rothe diagrams associated to $T(A_{n},p)$ highly resemble the Rothe diagrams mentioned in Remark 3.15 (i).
+[…]
+Figure 4. The generalized Rothe diagram of $\{\varepsilon_{1}-\varepsilon_{6},\varepsilon_{2}-\varepsilon_{7},\varepsilon_{3}-\varepsilon_{5}\}\in\Omega_{U(A_{7},3)}$
+Using the rook configuration model, one can show using an elementary combinatorial argument that the Poincaré polynomial of $U(A_{n},p)$ is $\prod_{d\in[n+1-m]_{m}}[d]_{q}$, as asserted in Table 2.
+The row–column symmetry of the model implies that there is a natural level-preserving bijection between the quasiparabolic sets $\Omega(A_{n},p)$ and $\Omega(A_{n},n+1-p)$; in particular, the Poincaré polynomials of $U(A_{n},p)$ and $U(A_{n},n+1-p)$ agree.
+Remark 4.8.
+The rook configuration model described in Example 4.7 is inspired by a similar model used by Garsia and Remmel in [5] to study a $q$-analogue of the Stirling numbers of the second kind.
+Garsia and Remmel place rooks on a general Ferrers board, and our model corresponds to the special case where the board is a $a\times b$ rectangle containing $\min(a,b)$ non-attacking rooks, with the set $\mathscr{D}_{U}(R)$ in our construction corresponding to the positions marked with a circle on the board; see [5, Figure (I.7)].
+Example 4.9 (Type $D$).
+For all integers $n\geq 4$ and $1\leq p\leq n$, the triple $T(D_{n},p)$ is a QP-triple whenever we have $1\leq p\leq\lfloor n/2\rfloor$, but not when $\lfloor n/2\rfloor<p<n-2$.
+However, if we relax the setting of Definition 4.6, then the triple
+$T^{\prime}(\Gamma,p)=(W(D_{n}),S(D_{n})\setminus\{s_{p}\},U^{\prime}_{p}:=U(D_{n},p)\cap U(D_{n},n)),$
+forms a QP-triple for all $1\leq p\leq n-2$. Note that the QP-triple $(W,I,U)$ from Theorem 3.14 is a special case of the triple $T^{\prime}(\Gamma,p)$ from (v): it is $T^{\prime}(D_{2k},k)$.
+We note that for all the QP-triples mentioned in the last paragraph, the maximum-cardinality orthogonal sets $R$ and their generalized Rothe diagrams can be interpreted in terms of rook configurations as in Example 4.7, but on a $p\times(n-p)$ board and using the rules of the Rothe diagrams of Theorem 3.14, in which the attacks come from above and to the left.
+As with Example 4.7, the rook configuration model can be used to give an elementary combinatorial proof of the factorization of the Poincaré polynomial, but this comes with the caveat that a position on the board may now correspond to either one of two roots $\varepsilon_{i}\pm\varepsilon_{j}$.
+The fact that $T(D_{n},p)$ fails to be a QP-triple when $\lfloor n/2\rfloor<p<n-2$ has to do with this caveat and the fact that in these cases $R$ does not contribute a rook to every row.
+By contrast, intersecting $U(D_{n},p)$ with $U(D_{n},n)$ to produce $U^{\prime}_{p}$ eliminates this complication for $T^{\prime}(D_{n},p)$: all elements of $U(D_{n},n)$ are of the form $\varepsilon_{i}+\varepsilon_{j}$, so the positions on the board are now in a natural one-to-one correspondence with the elements of $U^{\prime}_{p}$, which is why $T^{\prime}(D_{n},p)$ does form a QP-triple for all $1\leq p\leq n$ in a similar way to Example 4.7.
+Example 4.10 (Type $D$).
+For any integer $k\geq 2$, the triples $T(D_{2k},2k)$, $T(D_{2k},2k-1)$, $T(D_{2k-1},2k-1)$, and $T(D_{2k-1},2k-2)$ are all QP-triples. Note that they are different from the QP-triples in Example 4.9.
+Of the above four triples, the first triple $T(D_{2k},2k)$ is precisely the triple $(W,I,U)$ appearing in Theorem 3.11.
+The other three triples give rise to essentially the same quasiparabolic set as $T(D_{2k},2k)$, in the sense that their associated quasiparabolic sets $\Omega_{U(\Gamma,p)}$ all admit canonical level-preserving bijections to the quasiparabolic set $\Omega(D_{2k},2k)$.
 This bijection is induced by the symmetry of the Dynkin diagram for $\Omega(D_{2k},2k-1)$.
 For the case of $\Omega=\Omega(D_{2k-1},2k-1)$, each element of $\Omega$ corresponds to a (non-perfect) matching of $\{1,2,\dots,2k-1\}$ into (k1)(k-1) pairs and one unmatched point, and the canonical bijection from $\Omega$ to $\Omega(D_{2k},2k)$ sends such an element to the $2k$-root in $\Omega(D_{2k},2k)$ corresponding to the unique perfect matching containing those (k1)(k-1) pairs as blocks.
-  • (iii)
-    The triple $T(A_{n},p)$ is a QP-triple for all integers $n\geq 1$ and $1\leq p\leq n$.
-    The residue sets arising from the triple $T(A_{n},p)$ resemble the Rothe diagrams mentioned in Remark 3.14 (i).
-More specifically, the set $U(A_{n},p)$ consists of the roots of the form $\varepsilon_{i}-\varepsilon_{j}$ where $1\leq i\leq p<j\leq n+1$, so we may identify $U$ with the positions on a $p\times(n+1-p)$ board with rows labeled by 1 to $p$ from top to bottom and columns labelled by $p+1$ to n+1pn+1-p from left to right.
-Each element of $R\in\Omega_{U(A_{n},p)}$ consists of $\min(p,n+1-p)$ orthogonal roots, and if we view $R$ as an involution in the natural way, as the product $w_{R}\in S_{n+1}$ of the 2-cycles $(i,j)$ for which $\varepsilon_{i}-\varepsilon_{j}\in R$, then a root $\varepsilon_{a}-\varepsilon_{b}\in U(A_{n},p)$ is a residue of $R$ with respect to $U(A_{n},p)$ if and only if we have both $w_{R}(a)<b$ and $w_{R}(b)>a$.
-Equivalently, a position $(a,b)$ on the board is a residue $R$ if and only if it is not attacked by a rook to its right or above it in the rook configuration corresponding to $R$, as shown in Figure 4, where the residues are marked by stars as in Figure 1.
-[…]
-    Figure 4. The generalized Rothe diagram of $\{\varepsilon_{1}-\varepsilon_{6},\varepsilon_{2}-\varepsilon_{7},\varepsilon_{3}-\varepsilon_{5}\}\in\Omega_{U(A_{7},3)}$
-    Using the rook configuration model, it is possible to deduce that the Poincaré polynomial of $\Omega(A_{n},p)$ factors into quantum integers, as the product $PS_{U(A_{n},p)}=\prod_{D}[d]$ where $D$ is the set (with no repetition) containing each of the $\min(p,n+1-p)$ consecutive integers leading up to (and including) $\max(p,n+1-p)$.
-The row–column symmetry of the model also implies that there is a natural level-preserving bijection between the quasiparabolic sets $\Omega(A_{n},p)$ and $\Omega(A_{n},n+1-p)$; in particular, the Poincaré polynomials of $U(A_{n},p)$ and $U(A_{n},n+1-p)$ agree.
-  • (iv)
-    For all integers $n\geq 4$ and $1\leq p\leq n$, the triple $T(D_{n},p)$ is a QP-triple whenever we have $1\leq p\leq\lfloor n/2\rfloor$ (where $\lfloor\cdot\rfloor$ denotes the floor function), but not when $\lfloor n/2\rfloor<p<n-2$.
-However, if we relax the setting of Definition 4.6 and consider the set $U^{\prime}_{p}=U(D_{n},p)\cap U(D_{n},n)$, then the triple $T^{\prime}(\Gamma,P)=(W(D_{n}),S(D_{n})\setminus\{s_{p}\},U^{\prime}_{p})$ forms a QP-triple for all $1\leq p\leq n-2$.
-Note that the QP-triple $(W,I,U)$ from Theorem 3.13 is a special case of the triple $T^{\prime}(\Gamma,p)$ from (v), namely, $T^{\prime}(D_{2k},k)$.
-    We note that for all the triples $T(D_{n},p)$ mentioned in the last paragraph, the elements $R\in\Omega_{U(D_{n},p)}$ and their residues can be interpreted in terms of rook configurations on a $p\times(n-p)$ board as in (iii), with the caveat that a position on the board may now correspond to either one of two roots $\varepsilon_{i}\pm\varepsilon_{j}$.
-The fact that $T(D_{n},p)$ fails to be a QP-triple when $\lfloor n/2\rfloor<p<n-2$ has to do with this caveat and the fact that in these cases $R$ does not contribute a rook to every row.
+Example 4.11 (Type $E$).
+The last seven rows of Table 2 give the complete list of cases for which $T(E_{n},p)$ is a QP-triple for some $n\in\{6,7,8\}$ and $1\leq p\leq n$.
+We have verified that this is the complete list by computation, and we will give a detailed treatment of the triples $T(E_{7},7)$ and $T(E_{8},8)$, two examples that have natural connections to minuscule representations and del Pezzo surfaces.
+We also note that by our computation, in the cases where $T(E_{n},p)$ is not a QP-triple, it always happens that $T(E_{n},p)$ is not even a scaled $W_{I}$-set.
 Remark 4.12.
-  • (i)
-    To summarize Example 4.7, every simple root $\alpha_{p}$ in a root system of type $A_{n}$ or $D_{n}$ gives rise to one or two QP-triples: $T(A_{n},p)$ in type $A_{n}$, $T^{\prime}(D_{n},p)$ in type $D_{n}$, and $T(D_{n},p)$ if $p\leq\lfloor n/2\rfloor$.
-  • (ii)
-    The rook configuration model described in Example 4.7 (iii) is inspired by a similar model used by Garsia and Remmel in [3] to study a $q$-analogue of the Stirling numbers of the second kind.
-Garsia and Remmel place rooks on a general Ferrers board, and our model corresponds to the special case where the board is a $a\times b$ rectangle containing $\min(a,b)$ non-attacking rooks, with the residues in our construction corresponding to the positions marked with a circle on the board; see [3, Figure (I.7)].
-  • (iii)
-    The rook configuration models mentioned in Example 4.7 allow combinatorial proofs of the Poincaré polynomials of the quasiparabolic triples.
-In particular, it follows from the models that for all the quasiparabolic sets mentioned in the example, the Poincaré polynomials factor into quantum integers (that have combinatorial interpretations).
-For example, the quasiparabolic sets associated to $T(A_{n},p)$ and $T(A_{n},n+1-p)$ have the same Poincaré polynomial for any $1\leq p\leq\lfloor n+1-p\rfloor$, namely, $\prod_{d=n-2p+2}^{n+1-p}[d]_{q}$.
-Example 4.9.
   1. (i)
-    Table 1 gives the complete list of cases for which $T(E_{n},p)$ is a QP-triple, where $n\in\{6,7,8\},1\leq p\leq n$, and $\kappa$ stands for the cardinality of the elements in $\Omega_{U}$ in each case.
-We have verified that this is the complete list by computation, and we will give a detailed treatment of $T(E_{8},8)$ and $T(E_{7},7)$, two examples that have natural connections to minuscule representations and del Pezzo surfaces, but we do not have a combinatorial model or a conceptual proof of the quasiparabolic structure for all the examples.
-We also note that by our computation, in the cases where $T(E_{n},p)$ is not a QP-triple, it always happens that $T(E_{n},p)$ is not even a scaled $W_{I}$-set.
-    $T(\Gamma,p)$ $|U|$ $|\Omega_{U}|$ $\kappa$ degrees
-    $T(E_{6},1)$ 16 40 2 5, 8
-    $T(E_{6},2)$ 20 30 4 2, 3, 5
-    $T(E_{6},6)$ 16 40 2 5, 8
-    $T(E_{7},1)$ 32 120 4 3, 5, 8
-    $T(E_{7},2)$ 35 30 7 2, 3, 5
-    $T(E_{7},5)$ 35 30 7 2, 3, 5
-    $T(E_{7},7)$ 27 45 3 5, 9
-    $T(E_{8},1)$ 64 240 8 2, 3, 5, 8
-    $T(E_{8},2)$ 64 240 8 2, 3, 5, 8
-    $T(E_{8},5)$ 64 240 8 2, 3, 5, 8
-    $T(E_{8},6)$ 64 240 8 2, 3, 5, 8
-    $T(E_{8},8)$ 56 630 4 5, 9, 14
-    Table 1. Table of QP-triples of the form $T(E_{n},p)$
-  2. (ii)
-    The structures of the QP-triple $T(E_{7},2)$ and the associated quasiparolic set $\Omega(E_{7},2)$ are known by [8, Section 6.2]: the parabolic subgroup $W_{I}$ has type $A_{6}$, the set $\Omega(E_{7},2)$ coincides with the set of positive $7$-roots of $\Phi(E_{7})$ containing no alignments, and $W_{I}\cong S_{7}$ acts transitively on $\Omega(E_{7},2)$ in the standard action.
-The set $\Omega(E_{7},2)$ is also in a natural bijection with the $30$ inequivalent labellings of the Fano plane.
-Under the correspondence between the $7$-roots and labelled Fano planes, the generalized Rothe diagram of a labelled Fano plane has a simple combinatorial description: a residue is a triple of non-collinear points such that each point has a higher label than the point collinear with the other two.
+    The structures of the QP-triple $T(E_{7},2)$ and the associated quasiparabolic set $\Omega(E_{7},2)$ are known by [12, Section 6.2]: the parabolic subgroup $W_{I}$ has type $A_{6}$, the set $\Omega(E_{7},2)$ coincides with the set of positive $7$-roots of $\Phi(E_{7})$ containing no alignments, and $W_{I}\cong S_{7}$ acts transitively on $\Omega(E_{7},2)$ in the standard action.
+To each root $\alpha\in U(E_{7},2)$ we may naturally associate a triple $t_{\alpha}$ of integers from the set $7]:=\{1,2,...,7\}$, and $\Omega(E_{7},2)$ can be naturally identified with the $30$ inequivalent labellings of the Fano plane using the set $[7]$.
+It can be shown that a root $\alpha\in U(E_{7},2)$ dominates a labelling $L$ if and only if the elements of $t_{\alpha}$ do not form a line in $L$ and each element in $t_{\alpha}$ is larger than the unique element collinear with the other two in $L$.
 For example, the unique $W_{I}$-minimal labelling
     $\{136,145,127,235,246,347,567\}$
-    is the only labelling with no residues, and the unique $W_{I}$-maximal labelling
+    is the only labelling dominated by no root, and the unique $W_{I}$-maximal labelling
     $\{123,145,246,257,347,356,167\}$
-    has seven residues: $267$, $357$, $367$, $456$, $457$, $467$, and $567$.
+    is dominated by seven roots: $267$, $357$, $367$, $456$, $457$, $467$, and $567$.
 The generalized Rothe diagram of each labelling can be interpreted as the set of non-positions in the context of games arising from the Steiner system $S(2,3,7)$; see [16].
-  3. (iii)
-    For the triple $T(E_{8},2)$, the results of [8, Section 6.3] show that the corresponding subgroup $W_{I}$ has type $A_{7}$, the set $\Omega(E_{8},2)$ coincides with the set of positive $8$-roots of $\Phi(E_{8})$, and $W_{I}\cong S_{8}$ acts transitively on $\Omega(E_{8},2)$ in the standard action.
+  2. (ii)
+    For the QP-triple $T(E_{8},2)$, the results of [12, Section 6.3] show that the corresponding subgroup $W_{I}$ has type $A_{7}$, the set $\Omega(E_{8},2)$ coincides with the set of positive $8$-roots of $\Phi(E_{8})$, and $W_{I}\cong S_{8}$ acts transitively on $\Omega(E_{8},2)$ in the standard action.
 The set $\Omega(E_{8},2)$ is in a natural bijection with the set of $240$ inequivalent labellings of the Fano plane by the set $\{1,2,\ldots,8\}$, as well as with the 240 packings of the projective space $PG(3,2)$.
-The generalized Rothe diagram of each $8$-root in $\Omega(E_{8},2)$ has a characterization similar to the one in (ii), although the details are more complicated.
+The generalized Rothe diagram of each $8$-root in $\Omega(E_{8},2)$ has a characterization similar to the one in (i), although the details are more complicated.
 This particularly intricate example is the subject of the paper [8].
+4.3. Binary matroids and symmetric spaces
+The goal of this subsection is to explain several intriguing connections that the QP-triples from Table 1 and Table 2 have with matroids and with symmetric spaces.
+We refer the reader to [7, Chapter 1 and Chapter 2] for basic definitions concerning matroids.
+For background on symmetric spaces, including Cartan’s classification of them, we refer the reader to [2, Chapter 28] and [14, Chapter X].
+Let $T=(W,I,U)$ be a QP-triple described in Table 1 or Table 2.
+It turns out that every $R\in\Omega_{U}$ gives rise to a matroid, $M_{R}$, whose ground set is $R$ and whose set of flats is given by the set
+(4.1) $\{R\cap\left(\bigcap_{R^{\prime}\in Y}R^{\prime}\right):Y\subseteq\Omega_{U}\}.$
+Furthermore, the isomorphism type of $M_{R}$ depends only on the triple $T$ and not on $R$.
+The isomorphism type of the matroid $M_{R}(R\in\Omega_{U})$ is given in the last columns of Tables 1 and 2, whose notation we now explain below.
+For nonnegative integers $j\leq k$, we use $U_{j,k}$ to denote the uniform matroid with ground set $[k]$ and rank $j$, where a subset of the ground set is independent if and only if it has size at most $j$.
+The notation $U^{(2)}_{j,k}$ stands for the doubling of $U_{j,k}$, where the ground set and independent sets are given by the doublings of those of $U_{j,k}$.
+Here, by a doubling of a set $A\subseteq[k]$ we mean the set $A\cup\{a^{\prime}:a\in A\}$, where $1,1^{\prime},2,2^{\prime},...,k,k^{\prime}$ are understood to be $2k$ distinct elements.
+For each positive integer $n$, the matroid $AG(n,2)$ is the $n$-dimensional affine space over ${\mathbb{F}}_{2}$, where the nonempty flats are given by the affine subspaces.
+Finally, the matroid $PG(n,2)$ is the $n$-dimensional projective space over ${\mathbb{F}}_{2}$, with the nonempty flats given by the projective subspaces.
+All of these matroids are examples of binary matroids, in the sense that they are representable over the field ${\mathbb{F}}_{2}$.
+More details about $U_{j,k},AG(n,2),PG(n,2),$ and binary matroids can be found in sections 1.2, 5.1, 5.4, and 6.3 of [7], respectively.
+Remarkably, we can use the matroid $M_{R}$ to predict the degrees of $U$ in all cases, in the following way.
+If we list the possible sizes for the flats of $M_{R}$ as $k_{0},k_{1},...,k_{r}$, in increasing order, then it turns out that each flat of size $k_{i}$ is a subset of precisely $d_{i}$ elements of $\Omega_{U}$ for some number $d_{i}$ that depends only on $i$ and not on the specific choice of the flat (or $R$), and that the sequence $(d_{0}/d_{1},d_{1}/d_{2},\ldots,d_{r-1}/d_{r})$ produces the sequence of degrees of the QP-set.
+Example 4.13.
+For the triple $(W,S,\Phi_{+})$ of type $E_{8}$ and for each $8$-root $R\in\Omega_{\Phi_{+}}$, the possible sizes for the flats of $M_{R}$ are $0,1,2,4$ and $8$, and each flat of these sizes is contained in precisely $d_{0}=2025,d_{1}=135,d_{2}=15,d_{3}=3,$ and $d_{4}=1$ element(s) of $\Omega_{U}$, respectively; it is possible to deduce the numbers $d_{1},d_{2},d_{3}$, and $d_{4}$ from Remark 2.1.
+The quotients $d_{i}/d_{i+1}$ are $2025/135=15$, $135/15=9$, $15/3=5$, and $3/1=3$, which agree with the degrees of $U$ by Remark 4.4 (i).
+As is the case in examples 4.74.11, the assertions from the three paragraphs preceding Example 4.13 can be checked using combinatorial arguments in types $A$ or $D$ and via computer verification (or by careful case-by-case analysis) in type $E$.
+However, we do not have a conceptual explanation for why the matroids can predict the degrees of $U$ via the quotients $d_{i}/d_{i+1}$.
+Indeed, we found it surprising that the predictions work for all our examples, because the degrees arise from the level generating function of $\Omega_{U}$ but the definition of the matroids $M_{R}$ does not directly involve generalized Rothe diagrams at all.
+It would be interesting to know whether the intersections in the set given by (4.1) have any inherent connections with generalized Rothe diagrams.
+We now restrict our attention to the QP-triples $T(\Gamma,p)$ described in Table 2 and discuss their connection with symmetric spaces.
+Recall that symmetric spaces are classified by Cartan, who reduced the classification to that of real simple Lie algebras ([3], [14, Chapter X]).
+To each irreducible symmetric space $G/K$ one may associate a so-called symmetric pair $({\mathfrak{g}},{\mathfrak{k}})$, which is a pair of real semisimple Lie algebras satisfying a Cartan decomposition ${\mathfrak{g}}={\mathfrak{k}}\oplus{\mathfrak{p}}$, where $\mathfrak{k}$} is the fixed-point subalgebra of a Lie algebra involution $\theta$ and $\mathfrak{p}$} is the (1)(-1)-eigenspace of $\theta$.
+The pair $({\mathfrak{g}},{\mathfrak{k}})$ is not necessarily unique, but the isomorphism types of $\mathfrak{g}$} and $\mathfrak{k}$} are uniquely determined by the symmetric space $G/K$ ([14, Theorem 6.1]).
+The roots of $\mathfrak{g}$} whose root spaces lie in $\mathfrak{p}$} are the noncompact roots of $({\mathfrak{g}},{\mathfrak{k}})$ ([14, Section VIII.7]).
+Using Cartan’s classification of symmetric spaces and symmetric pairs ([14, Table X.V]), it can be checked case by case that for every symmetric space $G/K$ whose associated symmetric pair $({\mathfrak{g}},{\mathfrak{k}})$ consists of two simply-laced Lie algebras of the same rank, we may choose $\mathfrak{k}$} in such a way that there is a QP-triple $(W,I,U)=T(\Gamma,p)$ from Table 2 with the property that $\mathfrak{g}$} is of type $\Gamma$, $U=U(\Gamma,p)$ consists precisely of the noncompact roots of $({\mathfrak{g}},{\mathfrak{k}})$, and we have $I=S\cap W^{\prime}$ where $W^{\prime}$ is the Weyl group of $\mathfrak{k}$}.
+Moreover, every QP-triple $(W,I,U)=T(\Gamma,p)$ in Table 2 satisfies the above property for a suitable symmetric pair $({\mathfrak{g}},{\mathfrak{k}})$ of equal-rank simply-laced Lie algebras associated to a symmetric space.
+Example 4.14.
+The QP-triple $(W,I,U):=T(E_{8},8)=(W(E_{8}),S(E_{8})\setminus\{s_{8}\},U(E_{8},8))$ corresponds to the symmetric space of Cartan type EIX.
+In this case, $\mathfrak{g}$} is the Lie algebra of type $E_{8}$, the positive roots of $\mathfrak{p}$} are the roots in $U$, i.e., the positive roots in which the simple root $\alpha_{8}$ appears with odd coefficient.
+The positive roots of $\mathfrak{k}$} are the positive roots in which the simple root $\alpha_{8}$ appears with even coefficient, which include the highest root $\theta$ of $\Phi$.
+The Weyl group $W^{\prime}$ of $\mathfrak{k}$} is of type $E_{7}+A_{1}$; the set $I=\{s_{1},s_{2},\ldots,s_{7}\}=S\cap W^{\prime}$ generates the $E_{7}$ component of $W^{\prime}$, while the reflection $s_{\theta}$, which commutes with all elements of $I$, generates the $A_{1}$ component.
+The common size $\kappa$ of the elements of $\Omega_{U}$ is $4$, which agrees with the rank of the symmetric space in question.
+Table 3 details the correspondences between the Cartan types of the symmetric spaces, the equal-rank simply-laced symmetric pairs, and all QP-triples from Table 2.
+The second to last column of the table gives the rank of the symmetric space, which turns out to be equal to the size $\kappa$ of the elements in $\Omega_{U}$ in all cases.
+The last column describes the decomposition of $\mathfrak{k}$}, where we consider the one-dimensional abelian Lie algebra $\mathbb{R}$} to be of simply-laced Cartan type $\mathbb{R}$}”.
+Note that while we have to distinguish the first two rows in Table 2 because the matroid for the QP-triple $T(A_{2k-1},k)$ differs from those associated with the other QP-triples of type $A$, all the QP-triples of type $A$ correspond to a single type of symmetric space, regardless of the parity of the rank of $\Phi$ or the value of $p$.
+Similarly, the QP-triples of type $D$ correspond only to two types of symmetric spaces, even though they correspond to four distinct types of matroids in Table 2.
+$G/K$ $\Gamma$ $p$ $\text{\rm rank}(G/K)$ $\mathfrak{k}$}
+AIII $A_{n}$ $1,2,\ldots,n$ $\min(p,n+1-p)$ $A_{p-1}+A_{n-p}+{\mathbb{R}}$
+DI $D_{n}$ $1,2,\ldots,\lfloor n/2\rfloor$ $2p$ $D_{p}+D_{n-p}$
+DIII $D_{n}$ n1,nn-1,n $\lfloor n/2\rfloor$ $A_{n-1}+{\mathbb{R}}$
+EII $E_{6}$ 22 44 $A_{5}+A_{1}$
+EIII $E_{6}$ $1,6$ 22 $D_{5}+{\mathbb{R}}$
+EV $E_{7}$ $2,5$ 77 $A_{7}$
+EVI $E_{7}$ 11 44 $D_{6}+A_{1}$
+EVII $E_{7}$ 77 33 $E_{6}+{\mathbb{R}}$
+EVIII $E_{8}$ $1,2,5,6$ 88 $D_{8}$
+EIX $E_{8}$ 88 44 $E_{7}+A_{1}$
+Table 3. Correspondence between symmetric spaces and QP-triples $T(\Gamma,p)$
 5. The QP-triples $T(E_{8},8)$ and $T(E_{7},7)$
@@
 $T(E_{8},8)=(W(E_{8}),S(E_{8})\setminus\{s_{8}\},U(E_{8},8))\quad\text{and}\quad T(E_{7},7)=(W(E_{7}),S(E_{7})\setminus\{s_{7}\},U(E_{7},7))$
-from Example 4.9 in detail.
+from Section 4.2 in detail.
 The sets $U(E_{8},8)$ and $U(E_{7},7)$ have interesting natural connections to minuscule representations and del Pezzo surfaces.
-We give a non-computational proof that $T(E_{8},8)$ is a QP-triple in Section 5.1, and we use this result to deduce that $T(E_{7},7)$ is a QP-triple in Section 5.2.
-In Section 5.3, we show that the Pfaffian of $T(E_{7},7)$ coincides with the invariant polynomial associated to a 27-dimensional minuscule representation of type $E_{6}$, and we also describe the generalized Rothe diagrams of the elements of $\Omega(E_{7},7)$, which we expect may have applications to Schubert varieties or del Pezzo surfaces.
+Recall from Section 4.3 that we can verify by computation that $T(E_{8},8)$ corresponds to the matroid $AG(2,2)$ and the symmetric space of type EIX, and that $T(E_{7},7)$ corresponds to the matroid $PG(1,2)$ and the symmetric space of type EVII.
+Our goal in this section is to give non-computational proofs of four theorems: we show in Theorem 5.9 that $T(E_{8},8)$ is a QP-triple; we deduce in Theorem 5.13 that $T(E_{7},7)$ is a QP-triple; we explicitly describe in Theorem 5.17 the generalized Rothe diagrams of the elements of $\Omega(E_{7},7)$, which we expect may have applications to Schubert varieties or del Pezzo surfaces; and, we prove in Theorem 5.19 that the generalized Pfaffian of $T(E_{7},7)$ coincides with the invariant polynomial associated to a 27-dimensional minuscule representation of type $E_{6}$.
 5.1. $T(E_{8},8)$
@@
 Our strategy for proving the quasiparabolic structure of $\Omega_{U_{1}}$ is to leverage the quasiparabolic structure of the set $\Omega_{\Phi_{+}}$ of the $8$-roots, so inclusions of the form $Q\subseteq R$ where $Q\in\Omega_{U_{1}}$ and $R\in\Omega_{\Phi_{+}}$ will play an important role in our arguments, as will chains of the form $Q\subseteq H\subseteq R$ where $R$ is an $8$-root, $H\in D_{6}(R)$, and $Q$ is a feature in $R$.
-More specifically, it will be useful to study the relationship between the residues of $Q,H,$ and $R$ with various $8$-heights in such a chain.
+More specifically, given such a chain it will be useful to study the relationship between the roots with various $8$-heights that dominate $Q,H,$ or $R$.
 We will do so in lemmas 5.25.3 and 5.55.8.
@@
 Let $\Phi$ be the root system of type $E_{8}$, and let $R$ be a positive $8$-root of $\Phi$. Suppose that $H\in D_{4}(R)$, or $H\in D_{6}(R)$, or $H=R$. For each $i\in\{0,1,2\}$, we define
-$\mathrm{Res}^{(i)}(H):=\text{\rm Res}(H)\cap U_{i}\text{\; and \;}\rho^{({i})}(H)=\lvert\mathrm{Res}^{(i)}(H)\rvert$.
+$\mathscr{D}^{(i)}(H):=\mathscr{D}(H)\cap U_{i}\text{\; and \;}\rho^{({i})}(H)=\lvert\mathscr{D}^{(i)}(H)\rvert$.
 Recall from Section 2.1 that for any $H\in D_{6}(R)$ where $R$ is an $8$-root, the $D_{6}$ subsystem $\Xi:=\text{\rm Span}(H)\cap\Phi$ has positive roots $\{\varepsilon_{i}\pm\varepsilon_{j}:1\leq i<j\leq 6\}$, simple roots $\{\varepsilon_{1}-\varepsilon_{2},\varepsilon_{2}-\varepsilon_{3},\ldots,\varepsilon_{5}-\varepsilon_{6},\varepsilon_{5}+\varepsilon_{6}\}$, and highest root $\varepsilon_{1}+\varepsilon_{2}$ in the coordinate realization of $\Xi$.
@@
   1. (i)
-    We have $\mathrm{Res}^{(2)}(R)=\{\theta\}$ and $\theta\in\text{\rm Res}_{\Phi_{+}}(H)$.
+    We have $\mathscr{D}^{(2)}(R)=\{\theta\}$ and $\theta\in\mathscr{D}_{\Phi_{+}}(H)$.
   2. (ii)
-    If $\lvert H\cap U_{0}\rvert<4$, then $\mathrm{Res}^{(0)}(H)=\emptyset$.
+    If $\lvert H\cap U_{0}\rvert<4$, then $\mathscr{D}^{(0)}(H)=\emptyset$.
   3. (iii)
-    If $\lvert H\cap U_{1}\rvert<4$, then $\theta\notin\Xi$ and $\mathrm{Res}^{(2)}(H)=\emptyset$.
+    If $\lvert H\cap U_{1}\rvert<4$, then $\theta\notin\Xi$ and $\mathscr{D}^{(2)}(H)=\emptyset$.
 Proof.
@@
 We prove (ii) by proving its contrapositive.
-Suppose that $\mathrm{Res}^{(0)}(H)$ contains a root $\gamma$.
+Suppose that $\mathscr{D}^{(0)}(H)$ contains a root $\gamma$.
 Applying Proposition 4.1 (i) to $\Xi=\text{\rm Span}(H)\cap\Phi$ shows that $\gamma$ is the highest or second highest root in a $D_{4}$ subsystem $\Psi\subseteq\Xi$ corresponding to some feature $Q\subseteq H$.
@@
 This contradicts the assumption that $|H\cap U_{1}|<4$, so we have $\theta\notin\Xi$.
-It follows that $\theta\notin\text{\rm Res}(H)$, so we have $\mathrm{Res}^{(2)}(H)=\emptyset$, proving (iii).
+It follows that $\theta\notin\mathscr{D}(H)$, so we have $\mathscr{D}^{(2)}(H)=\emptyset$, proving (iii).
 ∎
@@
   • (ii)
-    In the setting of (i), we have $\theta\in\text{\rm Res}(Q_{1})$. We also have
-    $\mathrm{Res}^{(2)}(R)=\{\theta\},\quad\mathrm{Res}^{(1)}(R)=\text{\rm Res}_{U_{1}}(Q_{1}),\ \text{and}\quad\mathrm{Res}^{(0)}(R)=\text{\rm Res}(Q_{0}).$
+    In the setting of (i), we have $\theta\in\mathscr{D}(Q_{1})$. We also have
+    $\mathscr{D}^{(2)}(R)=\{\theta\},\quad\mathscr{D}^{(1)}(R)=\mathscr{D}_{U_{1}}(Q_{1}),\ \text{and}\quad\mathscr{D}^{(0)}(R)=\mathscr{D}(Q_{0}).$
 Proof.
@@
 The last paragraph shows that $\theta$ lies in the $D_{4}$-subsystem $\Phi\cap\text{\rm Span}(Q_{1})$.
-Since $\theta\in\text{\rm Res}_{\Phi_{+}}(Q_{1})$ by Lemma 5.2 (i), it follows that $\theta\in\text{\rm Res}(Q_{1})$.
+Since $\theta\in\mathscr{D}_{\Phi_{+}}(Q_{1})$ by Lemma 5.2 (i), it follows that $\theta\in\mathscr{D}(Q_{1})$.
 This proves the first assertion of (ii).
-The equation $\mathrm{Res}^{(2)}(R)=\{\theta\}$ holds by Lemma 5.2 (ii).
-The fact that $Q_{1}\subseteq R$ implies that $\mathrm{Res}^{(1)}(R)=\text{\rm Res}_{\Phi_{+}}(R)\cap U_{1}\subseteq\text{\rm Res}_{U_{1}}(Q_{1})$.
-Conversely, we have $\text{\rm Res}_{U_{1}}(Q_{1})\subseteq\text{\rm Res}_{\Phi_{+}}(R)\cap U_{1}$ because $s_{\beta}(\gamma)>0$ for all $\beta\in Q_{0}\subseteq U_{0}$ and $\gamma\in U_{1}$, so we have $\mathrm{Res}^{(1)}(R)=\text{\rm Res}_{U_{1}}(Q_{1})$.
-Finally, we have $\text{\rm Res}(R)=\bigcup_{Q\in D_{4}(R)}\text{\rm Res}(Q)$ by Proposition 4.1 (i) and therefore $\mathrm{Res}^{(0)}(R)=\bigcup_{Q\in D_{4}(R)}\mathrm{Res}^{(0)}(Q)$.
-Every feature $Q$ other than $Q_{0}$ appearing in the union satisfies $\lvert Q\cap U_{0}\rvert<4$ and thus $\mathrm{Res}^{(0)}(Q)=\emptyset$ by Lemma 5.2 (ii), so it follows that $\mathrm{Res}^{(0)}(R)=\mathrm{Res}^{(0)}(Q_{0})$.
+The equation $\mathscr{D}^{(2)}(R)=\{\theta\}$ holds by Lemma 5.2 (ii).
+The fact that $Q_{1}\subseteq R$ implies that $\mathscr{D}^{(1)}(R)=\mathscr{D}_{\Phi_{+}}(R)\cap U_{1}\subseteq\mathscr{D}_{U_{1}}(Q_{1})$.
+Conversely, we have $\mathscr{D}_{U_{1}}(Q_{1})\subseteq\mathscr{D}_{\Phi_{+}}(R)\cap U_{1}$ because $s_{\beta}(\gamma)>0$ for all $\beta\in Q_{0}\subseteq U_{0}$ and $\gamma\in U_{1}$, so we have $\mathscr{D}^{(1)}(R)=\mathscr{D}_{U_{1}}(Q_{1})$.
+Finally, we have $\mathscr{D}(R)=\bigcup_{Q\in D_{4}(R)}\mathscr{D}(Q)$ by Proposition 4.1 (i) and therefore $\mathscr{D}^{(0)}(R)=\bigcup_{Q\in D_{4}(R)}\mathscr{D}^{(0)}(Q)$.
+Every feature $Q$ other than $Q_{0}$ appearing in the union satisfies $\lvert Q\cap U_{0}\rvert<4$ and thus $\mathscr{D}^{(0)}(Q)=\emptyset$ by Lemma 5.2 (ii), so it follows that $\mathscr{D}^{(0)}(R)=\mathscr{D}^{(0)}(Q_{0})$.
 This completes the proof of (ii).
@@
 It follows that $\varepsilon_{1}\pm\varepsilon_{j}\in U_{1},\varepsilon_{2}\pm\varepsilon_{k}\in U_{0}$, and $\beta_{1}\in U_{1}$, as claimed.
-Since every simple root of $\Xi$ appears in the decomposition of $\psi$ into simple roots and $\psi$ and $\beta_{1}$ have the same $8$-height, it follows that all simple roots of $\Xi$ other than $\beta$ lie in $U_{0}$, proving (iii).
+Since every simple root of $\Xi$ appears in the decomposition of $\psi$ into simple roots and $\psi$ and $\beta_{1}$ have the same $8$-height, it follows that all simple roots of $\Xi$ other than $\beta_{1}$ lie in $U_{0}$, proving (iii).
 Part (iv) follows from (iii) by expressing $\gamma$ as a linear combination of simple roots. ∎
@@
   • (i)
-    All residues of $Q$ have $8$-height 1, so that $\text{\rm Res}(Q)=\mathrm{Res}^{(1)}(Q)$.
+    All roots that dominate $Q$ have $8$-height 1, so that $\mathscr{D}(Q)=\mathscr{D}^{(1)}(Q)$.
   • (ii)
     We have $H\in D_{6}(R)$, and we also have
-    $\mathrm{Res}^{(2)}(H)=\{\theta\},\quad\mathrm{Res}^{(1)}(H)=\text{\rm Res}_{U_{1}}(Q),\ \text{and}\quad\mathrm{Res}^{(0)}(H)=\text{\rm Res}(Q_{0}).$
+    $\mathscr{D}^{(2)}(H)=\{\theta\},\quad\mathscr{D}^{(1)}(H)=\mathscr{D}_{U_{1}}(Q),\ \text{and}\quad\mathscr{D}^{(0)}(H)=\mathscr{D}(Q_{0}).$
 Proof.
-The assumptions that $\lvert Q\cap Q_{0}\rvert=2<4$ and $\lvert Q\cap Q_{1}\rvert=2<4$ imply that $\mathrm{Res}^{(0)}(Q)=\emptyset$ and $\mathrm{Res}^{(2)}(Q)=\emptyset$ by Lemma 5.2 (ii) and (iii), respectively, and (i) follows.
+The assumptions that $\lvert Q\cap Q_{0}\rvert=2<4$ and $\lvert Q\cap Q_{1}\rvert=2<4$ imply that $\mathscr{D}^{(0)}(Q)=\emptyset$ and $\mathscr{D}^{(2)}(Q)=\emptyset$ by Lemma 5.2 (ii) and (iii), respectively, and (i) follows.
 The fact that $H\in D_{6}(R)$ follows from Lemma 2.15 (i).
 Let $\Xi=\text{\rm Span}(H)\cap\Phi$ be the corresponding $D_{6}$ subsystem in $\Phi$.
-We have $\theta\notin\Xi$ by Lemma 5.5 (i), so $\theta\notin\text{\rm Res}(H)$ and $\mathrm{Res}^{(2)}(H)=\emptyset$, proving the first equality in (ii).
+We have $\theta\notin\Xi$ by Lemma 5.5 (i), so $\theta\notin\mathscr{D}(H)$ and $\mathscr{D}^{(2)}(H)=\emptyset$, proving the first equality in (ii).
 The two remaining equalities follow from Proposition 4.1 (i) and Lemma 5.2 (ii) in the same way as the last two equalities in Lemma 5.3 (ii).
@@
 Let $\beta\in\Xi\cap U_{1}$.
-If we have $\beta=\varepsilon_{1}-\varepsilon_{i}$ for some $2\leq i\leq 6$ or $\beta=\varepsilon_{1}+\varepsilon_{i}$ for some $i>a$, then we have $s_{\gamma}(\beta)<0$ for $\gamma=\varepsilon_{1}+\varepsilon_{a}\in H$, so $\beta\notin\text{\rm Res}(H)$.
-On the other hand, if $\beta=\varepsilon_{1}+\varepsilon_{i}$ for some $1<i<a$ then we have $s_{\gamma}(\beta)>0$ for all $\gamma\in H$ and hence $\beta\in\text{\rm Res}(H)$.
-It follows that $\mathrm{Res}^{(1)}(H)=\{\varepsilon_{1}+\varepsilon_{i}:1<i<a\}$ and $\rho^{({1})}(H)=a-2$.
+If we have $\beta=\varepsilon_{1}-\varepsilon_{i}$ for some $2\leq i\leq 6$ or $\beta=\varepsilon_{1}+\varepsilon_{i}$ for some $i>a$, then we have $s_{\gamma}(\beta)<0$ for $\gamma=\varepsilon_{1}+\varepsilon_{a}\in H$, so $\beta\notin\mathscr{D}(H)$.
+On the other hand, if $\beta=\varepsilon_{1}+\varepsilon_{i}$ for some $1<i<a$ then we have $s_{\gamma}(\beta)>0$ for all $\gamma\in H$ and hence $\beta\in\mathscr{D}(H)$.
+It follows that $\mathscr{D}^{(1)}(H)=\{\varepsilon_{1}+\varepsilon_{i}:1<i<a\}$ and $\rho^{({1})}(H)=a-2$.
 Similarly, we have $\rho^{({1})}(rH)=a^{\prime}-2$, so that $\rho^{({1})}(rH)-\rho^{({1})}(H)=a^{\prime}-a$.
-By Remarks 2.13 and 3.5 (ii), the assumption that $\rho(rQ)>\rho(Q)$ implies that we have one of the following situations:
+By Remarks 2.14 and 3.6 (ii), the assumption that $\rho(rQ)>\rho(Q)$ implies that we have one of the following situations:
   1. (1)
@@
 Now suppose that $\alpha$ is a simple root of $\Phi$.
-This implies that $\alpha$ is also a simple root of $\Xi$, so the reflection $r$ induces a basic transposition $s_{i}:=(i,i+1)$ on the perfect matching $M:=\{\{1,a\},\{b,c\}\}$ corresponding to $Q$, with $\{i,i+1\}\subseteq\{1,a,b,c\}$ because $Q$ is the support of $\alpha$.
+This implies that $\alpha$ is also a simple root of $\Xi$, so the reflection $r$ induces an adjacent transposition $s_{i}:=(i,i+1)$ on the perfect matching $M:=\{\{1,a\},\{b,c\}\}$ corresponding to $Q$, with $\{i,i+1\}\subseteq\{1,a,b,c\}$ because $Q$ is the support of $\alpha$.
 If $\{i,i+1\}\in M$, then $rH=H$ and (ii) follows.
@@
 To prove $\rho_{U_{1}}(Q_{1})<\rho_{U_{1}}(rQ_{1})$, it suffices to prove that $\rho^{({1})}(R)<\rho^{({1})}(rR)$ by Lemma 5.3 (ii). We have
-$\text{\rm Res}(R)=\text{\rm Res}(Q)\ \dot{\cup}\ \text{\rm Res}(R\setminus Q)\ \dot{\cup}\ \dot{\bigcup}_{Q\subseteq H^{\prime}\in D_{6}(R)}(\text{\rm Res}(H^{\prime})\setminus\text{\rm Res}(Q))$
+$\mathscr{D}(R)=\mathscr{D}(Q)\ \dot{\cup}\ \mathscr{D}(R\setminus Q)\ \dot{\cup}\ \dot{\bigcup}_{Q\subseteq H^{\prime}\in D_{6}(R)}(\mathscr{D}(H^{\prime})\setminus\mathscr{D}(Q))$
 by Proposition 4.1 (ii). Taking intersections with $U_{1}$ then implies that
-$\mathrm{Res}^{(1)}(R)=\mathrm{Res}^{(1)}(Q)\ \dot{\cup}\ \mathrm{Res}^{(1)}(R\setminus Q)\ \dot{\cup}\ \dot{\bigcup}_{Q\subseteq H^{\prime}\in D_{6}(R)}(\mathrm{Res}^{(1)}(H^{\prime})\setminus\text{\rm Res}(Q),$
-where we have $\mathrm{Res}^{(1)}(Q)=\text{\rm Res}(Q)$ by Lemma 5.6 (i) and $\text{\rm Res}(Q)=\mathrm{Res}^{(1)}(Q)\subseteq\mathrm{Res}^{(1)}(H^{\prime})$ for all $Q\subseteq H^{\prime}\in D_{6}(R)$ by Proposition 4.1 (i).
+$\mathscr{D}^{(1)}(R)=\mathscr{D}^{(1)}(Q)\ \dot{\cup}\ \mathscr{D}^{(1)}(R\setminus Q)\ \dot{\cup}\ \dot{\bigcup}_{Q\subseteq H^{\prime}\in D_{6}(R)}\mathscr{D}^{(1)}(H^{\prime})\setminus\mathscr{D}(Q),$
+where we have $\mathscr{D}^{(1)}(Q)=\mathscr{D}(Q)$ by Lemma 5.6 (i) and $\mathscr{D}(Q)=\mathscr{D}^{(1)}(Q)\subseteq\mathscr{D}^{(1)}(H^{\prime})$ for all $Q\subseteq H^{\prime}\in D_{6}(R)$ by Proposition 4.1 (i).
 Taking cardinalities of the sets in the last equation then yields the equation
@@
 by Lemma 5.7 (i).
-For any $Q\subseteq H^{\prime}\subseteq D_{6}(R)\setminus\{H\}$, we have $\lvert H^{\prime}\cap U_{0}\rvert=\lvert rH^{\prime}\cap U_{0}\rvert<4$ and therefore that $\mathrm{Res}^{(0)}(H^{\prime})=\mathrm{Res}^{(0)}(rH^{\prime})=\emptyset$ by Lemma 5.2 (ii).
+For any $Q\subseteq H^{\prime}\subseteq D_{6}(R)\setminus\{H\}$, we have $\lvert H^{\prime}\cap U_{0}\rvert=\lvert rH^{\prime}\cap U_{0}\rvert<4$ and therefore that $\mathscr{D}^{(0)}(H^{\prime})=\mathscr{D}^{(0)}(rH^{\prime})=\emptyset$ by Lemma 5.2 (ii).
 Furthermore, we have either $H^{\prime}=Q_{1}\cup Q$ or $\lvert H^{\prime}\cap Q_{1}\rvert<4$.
-In the former case, Lemma 5.3 (ii) and Proposition 4.1 (i) imply that $\theta\in\text{\rm Res}(Q_{1})\subseteq\text{\rm Res}(H^{\prime})$ and similarly $\theta\in\text{\rm Res}(rQ_{1})\subseteq\text{\rm Res}(rH^{\prime})$, so we have $\mathrm{Res}^{(2)}(H^{\prime})=\mathrm{Res}^{(2)}(rH^{\prime})=\{\theta\}$.
-In the latter case, we have $\mathrm{Res}^{(2)}(H^{\prime})=\mathrm{Res}^{(2)}(rH^{\prime})=\emptyset$ by Lemma 5.2 (iii).
+In the former case, Lemma 5.3 (ii) and Proposition 4.1 (i) imply that $\theta\in\mathscr{D}(Q_{1})\subseteq\mathscr{D}(H^{\prime})$ and similarly $\theta\in\mathscr{D}(rQ_{1})\subseteq\mathscr{D}(rH^{\prime})$, so we have $\mathscr{D}^{(2)}(H^{\prime})=\mathscr{D}^{(2)}(rH^{\prime})=\{\theta\}$.
+In the latter case, we have $\mathscr{D}^{(2)}(H^{\prime})=\mathscr{D}^{(2)}(rH^{\prime})=\emptyset$ by Lemma 5.2 (iii).
 It follows that $\rho^{({0})}(H^{\prime})=\rho^{({0})}(rH^{\prime})$ and $\rho^{({2})}(H^{\prime})=\rho^{({2})}(H^{\prime})$ in all cases, so we have
 $(\rho^{({1})}(rH^{\prime})-\rho^{({1})}(H^{\prime}))-(\rho(rQ)-\rho(Q))=(\rho(rH^{\prime})-\rho(H^{\prime}))-(\rho(rQ)-\rho(Q))>0$
-for any $Q\subseteq H^{\prime}\in D_{6}(R)\setminus\{H\}$, where the inequality holds by an application of Corollary 4.4 to the $6$-root $H^{\prime}$ of the $D_{6}$ system $\text{\rm Span}(H^{\prime})\cap\Phi$.
+for any $Q\subseteq H^{\prime}\in D_{6}(R)\setminus\{H\}$, where the inequality holds by an application of Corollary 4.5 to the $6$-root $H^{\prime}$ of the $D_{6}$-system $\text{\rm Span}(H^{\prime})\cap\Phi$.
 Equation (5.5) now implies that $\rho^{({1})}(rR)>\rho^{({1})}(R)$, as desired.
@@
   1. (i)
-    Using Remark 4.5 (ii) and Lemma 5.3 (i), one can show that exactly 135 of the positive $8$-roots in $\Omega_{\Phi_{+}}$ contain $\theta$ and that the remaining 1890 positive $8$-roots are in a 3-to-1 correspondence with the elements of the set $\Omega_{U}$ from Theorem 5.9.
+    By Example 4.13, the set $\Omega_{\Phi_{+}}$ contains exactly $2025$ positive $8$-roots, and exactly 135 of them contain $\theta$.
+One can show using Lemma 5.3 that the remaining 1890 positive $8$-roots are in a 3-to-1 correspondence with the elements of the set $\Omega_{U}$ from Theorem 5.9.
 The 33-to-11 ratio comes from the fact the each element $Q_{1}\in\Omega_{U}$ can be extended in three ways to a positive $8$-root $R=Q_{1}\ \dot{\cup}\ Q_{0}$, where $Q_{0}$ is one of the three features in the $D_{4}$-subsystem given by $\Psi=\{\beta\in\Phi:B(\beta,\gamma)=0\text{ for all }\gamma\in Q_{1}\}$.
-Combining this fact with Lemma 5.3 (ii), Remark 3.5 (ii), and Remark 4.5 (ii) shows that the Poincaré polynomial of $\Omega_{U}$ is $PS_{U}(q)=[5]_{q}[9]_{q}[14]_{q}$, which agrees with Table 1.
+Combining this fact with Lemma 5.3 (ii), Remark 3.6 (ii), and Table 1 shows that the Poincaré polynomial of $\Omega_{U}$ is $PS_{U}(q)=[5]_{q}[9]_{q}[14]_{q}$, which agrees with Table 2.
 In particular, the set $\Omega_{U}$ contains $5\cdot 9\cdot 14=630$ triples.
   2. (ii)
-    Recall from Remark 4.5 (iii) that the $8$-root
+    Recall from Remark 4.4 (iv) that the $8$-root
     $R=\{\alpha_{2},\alpha_{3},\alpha_{5},\alpha_{7},\theta_{4},\theta_{6},\theta_{7},\theta_{8}\}$
-    is the unique $W(E_{8})$-minimal element of $\Omega_{\Phi_{+}}$ and has no residues with respect to $\Phi_{+}$.
-It follows from Theorem 4.3 that $s_{8}(R)$ has a single residue, namely, $\theta_{8}$.
+    is the unique $W(E_{8})$-minimal element of $\Omega_{\Phi_{+}}$ and has level $\rho(R)=0$.
+It follows from Theorem 4.3 that a single root dominates $s_{8}(R)$, namely, $\theta_{8}$.
 Lemma 5.3 (ii) then implies that the unique $W(E_{7})$-minimal element of $\Omega_{U}$ is the quadruple $\{s_{8}(\alpha_{7}),s_{8}(\theta_{6}),s_{8}(\theta_{7}),s_{8}(\theta)\}$.
 5.2. $T(E_{7},7)$
-In this subsection, we maintain the notation of Theorem 5.9, identify the root system $\Phi(E_{7})$ with the subsystem $\Phi^{\prime}=\Phi_{8,0}$ of $\Phi$ in the usual way, and identify the set $U^{\prime}=U(E_{7},7)$ with the subset of $\Phi^{\prime}_{+}$ consisting of roots of positive odd $7$-height.
+In this subsection, we maintain the notation of Theorem 5.9, we identify the root system $\Phi(E_{7})$ with the subsystem $\Phi^{\prime}=\Phi_{8,0}$ of $\Phi$ in the usual way, and we identify the set $U^{\prime}=U(E_{7},7)$ with the subset of $\Phi^{\prime}_{+}$ consisting of roots of positive odd $7$-height.
 Let $\theta^{\prime}=\theta_{7}$ be the highest root of $\Phi^{\prime}$.
@@
   • (i)
-    We have $\alpha_{8}\not\in\text{\rm Res}_{U}(Q)$.
+    We have $\alpha_{8}\not\in\mathscr{D}_{U}(Q)$.
   • (ii)
-    Every element of $\text{\rm Res}_{U}(Q)$ is orthogonal to $\eta$.
+    Every element of $\mathscr{D}_{U}(Q)$ is orthogonal to $\eta$.
 Proof.
@@
 This can only happen if every $\beta\in Q\backslash\{\eta\}$ satisfies $B(\alpha_{8},\beta)=+1$, which implies that $s_{\beta}(\alpha_{8})<0$ for all $\beta\in Q\setminus\{\eta\}$, proving (i).
-Let $\gamma\in\text{\rm Res}_{U}(Q)$.
+Let $\gamma\in\mathscr{D}_{U}(Q)$.
 Remark 3.3 (i) implies that $\gamma\neq\eta$, so we have $B(\eta,\gamma)\neq 2$.
@@
 This would imply that $\gamma=\alpha_{8}$, contradicting (i).
-If we had $B(\eta,\gamma)=+1$, it would follow that $s_{\eta}(\gamma)<0$, and $\gamma$ would not be a residue.
+If we had $B(\eta,\gamma)=+1$, it would follow that $s_{\eta}(\gamma)<0$, and $\gamma$ would not dominate $Q$.
 We conclude that $B(\eta,\gamma)=0$, and (ii) follows.
@@
   • (v)
-    For every $T\in\Omega_{U^{\prime}}$, we have a bijection $\text{\rm Res}_{U^{\prime}}(T)\rightarrow\text{\rm Res}_{U}(\iota(T))$ sending each $\gamma\in\text{\rm Res}_{U^{\prime}}(T)$ to $s_{8}(\gamma)$.
+    For every $T\in\Omega_{U^{\prime}}$, we have a bijection $\mathscr{D}_{U^{\prime}}(T)\rightarrow\mathscr{D}_{U}(\iota(T))$ sending each $\gamma\in\mathscr{D}_{U^{\prime}}(T)$ to $s_{8}(\gamma)$.
 In particular, we have $\rho_{U^{\prime}}(T)=\rho_{U}(\iota(T))$.
@@
 This proves (iv).
-Let $T=\{\beta_{1},\beta_{2},\beta_{3}\}\in\Omega_{U^{\prime}}$ and $\gamma\in\text{\rm Res}_{U^{\prime}}(T)$.
+Let $T=\{\beta_{1},\beta_{2},\beta_{3}\}\in\Omega_{U^{\prime}}$ and $\gamma\in\mathscr{D}_{U^{\prime}}(T)$.
 Since $B(\eta,s_{8}(\gamma))=B(\theta,\gamma)=0$, we have $s_{s_{8}(\gamma)}(\eta)=\eta>0$.
-For each $i\in\{1,2,3\}$, we have $\gamma\neq\beta_{i}$ by Remark 3.2 (i), and we have $s_{\beta_{i}}(\gamma)>0$ since $\gamma\in\text{\rm Res}_{U^{\prime}}(T)$, so (iv) implies that $s_{s_{8}(\beta_{i})}(s_{8}(\gamma))>0$.
-It follows that $s_{8}(\gamma)\in\text{\rm Res}_{U}(\iota(T))$.
-Conversely, for any $\delta\in\text{\rm Res}_{U}(\iota(T))$, we have $\delta\in\eta_{U}^{\perp}$ by Lemma 5.11 (ii), so we have $\delta=s_{8}(\gamma)$ for some $\gamma\in U^{\prime}$ by (i).
-The assumption that $\delta\in\text{\rm Res}_{U}(\iota(T))$ implies that $\delta\notin\iota(T)$ and, therefore, that $\gamma\notin T$.
-Since $\delta=s_{8}(\gamma)\in\text{\rm Res}_{U}(\iota(T))$, it follows from (iv) that $\gamma\in\text{\rm Res}_{U^{\prime}}(T)$, which proves the first assertion of (v).
+For each $i\in\{1,2,3\}$, we have $\gamma\neq\beta_{i}$ by Remark 3.3 (i), and we have $s_{\beta_{i}}(\gamma)>0$ since $\gamma\in\mathscr{D}_{U^{\prime}}(T)$, so (iv) implies that $s_{s_{8}(\beta_{i})}(s_{8}(\gamma))>0$.
+It follows that $s_{8}(\gamma)\in\mathscr{D}_{U}(\iota(T))$.
+Conversely, for any $\delta\in\mathscr{D}_{U}(\iota(T))$, we have $\delta\in\eta_{U}^{\perp}$ by Lemma 5.11 (ii), so we have $\delta=s_{8}(\gamma)$ for some $\gamma\in U^{\prime}$ by (i).
+The assumption that $\delta\in\mathscr{D}_{U}(\iota(T))$ implies that $\delta\notin\iota(T)$ and, therefore, that $\gamma\notin T$.
+Since $\delta=s_{8}(\gamma)\in\mathscr{D}_{U}(\iota(T))$, it follows from (iv) that $\gamma\in\mathscr{D}_{U^{\prime}}(T)$, which proves the first assertion of (v).
 The second assertion follows from the first by taking cardinalities.
@@
   1. (i)
-    Recall from Remark 4.5 (ii) that there are $135$ positive $7$-roots in type $E_{7}$, and that the Poincaré polynomial of the associated to these 7-roots is $PS_{\Phi^{\prime}_{+}}(q)=[3]_{q}[5]_{q}[9]_{q}$.
+    Recall that by Remark 4.4 (ii), there are $135$ positive $7$-roots in type $E_{7}$, and the Poincaré polynomial of the associated to these 7-roots is $PS_{\Phi^{\prime}_{+}}(q)=[3]_{q}[5]_{q}[9]_{q}$.
 The $7$-roots are in a 3–to–1 correspondence with the elements of $\Omega_{U}^{\prime}$ similar to the correspondence mentioned in Remark 5.10 (i).
@@
 The sets $U=U(E_{8},8)$ and $U^{\prime}=U(E_{7},7)$ from Theorems 5.9 and 5.13 have representation-theoretic interpretations.
-We recall that the exceptional Lie algebras of types $E_{6}$, $E_{7}$, and $E_{8}$ have three minuscule representations between them: two $27$-dimensional representations in type $E_{6}$, $L(E_{6},\omega_{1})$ and $L(E_{6},\omega_{6})$, and a $56$-dimensional representation in type $E_{7}$, $L(E_{7},\omega_{7})$, where each $\omega_{i}$ indicates the fundamental weight that appears as the highest weight of the corresponding representation (see, for example, [5, Theorem 5.1.5]).
-If one restricts the adjoint representation of the simple Lie algebra ${\mathfrak{g}}(E_{8})$ to ${\mathfrak{g}}(E_{7})$, then the set $U$ corresponds to the weights of $L(E_{7},\omega_{7})$ [5, Proposition 10.5.5].
+We recall that the exceptional Lie algebras of types $E_{6}$, $E_{7}$, and $E_{8}$ have three minuscule representations between them: two $27$-dimensional representations in type $E_{6}$, $L(E_{6},\omega_{1})$ and $L(E_{6},\omega_{6})$, and a $56$-dimensional representation in type $E_{7}$, $L(E_{7},\omega_{7})$, where each $\omega_{i}$ indicates the fundamental weight that appears as the highest weight of the corresponding representation (see, for example, [9, Theorem 5.1.5]).
+If one restricts the adjoint representation of the simple Lie algebra ${\mathfrak{g}}(E_{8})$ to ${\mathfrak{g}}(E_{7})$, then the set $U$ corresponds to the weights of $L(E_{7},\omega_{7})$ [9, Proposition 10.5.5].
 Similarly, restricting the adjoint representation of ${\mathfrak{g}}(E_{7})$ to ${\mathfrak{g}}(E_{6})$ identifies $U^{\prime}$ with the set of weights of $L(E_{6},\omega_{1})$.
-The sets $U$ and $U^{\prime}$ also have natural connections to algebraic geometry: the roots in $U$ are in bijection with the 56 exceptional curves on a del Pezzo surface $\mathcal{S}$ of degree 2, and the roots in $U^{\prime}$ are in bijection with the 27 lines on a del Pezzo surface $\mathcal{S}^{\prime}$ of degree 3 [5, Sections 9.3 and 10.1].
+The sets $U$ and $U^{\prime}$ also have natural connections to algebraic geometry: the roots in $U$ are in bijection with the 56 exceptional curves on a del Pezzo surface $\mathcal{S}$ of degree 2, and the roots in $U^{\prime}$ are in bijection with the 27 lines on a del Pezzo surface $\mathcal{S}^{\prime}$ of degree 3 [9, Sections 9.3 and 10.1].
 Under these bijections, two roots are orthogonal if and only if the intersection number between their corresponding curves is equal to $1$, so the triples of $\Omega_{U^{\prime}}$ are in a bijective correspondence with the tritangent planes of $\mathcal{S}^{\prime}$.
-The quadruples of $\Omega_{U}$ corresponds bijectively to the what are sometimes called “tetrads” on $\mathcal{S}$; see, for example, [4, Proposition 5.4].
+The quadruples of $\Omega_{U}$ correspond bijectively to the what are sometimes called “tetrads” on $\mathcal{S}$; see, for example, [6, Proposition 5.4].
 Theorems 5.9 and 5.13 imply that the tritangent planes and tetrads carry natural partial orders induced by the quasiparabolic orders of $\Omega_{U^{\prime}}$ and $\Omega_{U}$, respectively.
-The existence of these induced orders are new to the best of our knowledge.
+The existence of these induced orders is new to the best of our knowledge.
 In the rest of this subsection, we discuss two applications of the quasiparabolic set $\Omega_{U^{\prime}}$ arising from the QP-triple $T(E_{7},7)$.
 First, we give explicit descriptions of the generalized Rothe diagrams associated to the elements $\Omega_{U^{\prime}}$ in Theorem 5.17.
-These descriptions are also new to our knowledge, and we expect they may have applications to Schubert varieties or del Pezzo surfaces.
-Secondly, we show in Theorem 5.19 that the Pfaffian of $\Omega_{U^{\prime}}$ agrees with the well-known invariant cubic polynomial associated to the minuscule representation $L(E_{6},\omega_{1})$.
+We expect that these diagrams, which are also new to our knowledge, may have applications to Schubert varieties or del Pezzo surfaces.
+Secondly, we show in Theorem 5.19 that the generalized Pfaffian of $\Omega_{U^{\prime}}$ agrees with the well-known invariant cubic polynomial associated to the minuscule representation $L(E_{6},\omega_{1})$.
 To state the theorems, we will use the following labelling conventions for the roots of $U$ adapted from the context of del Pezzo surfaces.
@@
 In this realization, the $126$ roots of type $E_{7}$ consist of the $56$ coordinate permutations of $4(\varepsilon_{1}-\varepsilon_{2})$ and the $70$ coordinate permutations of $\alpha^{\prime}_{7}$; a root is positive if and only if its rightmost nonzero coefficient is negative.
-More details about this realization can be found in [5, Example 8.1.7].
+More details about this realization can be found in [9, Example 8.1.7].
 11223344556677
@@
 where $\tau(i)$ refers to the same vertex labeled by $i$ in the Dynkin diagram in the convention of Figure 2 (d).
-It is known that the map $f$ is a bijection between $U$ and the set of vectors $\{\pm v_{i,j}:1\leq i<j\leq 8\}$, where $v_{i,j}=4(\varepsilon_{i}+\varepsilon_{j})-\sum_{k=1}^{8}\varepsilon_{k}$ for all $i,j$ ([5, Example 8.1.7]).
+It is known that the map $f$ is a bijection between $U$ and the set of vectors $\{\pm v_{i,j}:1\leq i<j\leq 8\}$, where $v_{i,j}=4(\varepsilon_{i}+\varepsilon_{j})-\sum_{k=1}^{8}\varepsilon_{k}$ for all $i,j$ ([9, Example 8.1.7]).
 We will write $X_{ij}$ for $v_{i,j}$ and $Y_{ij}$ for $v_{i,j}$.
@@
     If $1<i<j<6$, then the triple $a_{i}c_{ij}b_{j}$ has level 5+ij5+i-j, and we have
-    $\text{\rm Res}_{U^{\prime}}(a_{i}c_{ij}b_{j})=\{a_{h}:1\leq h<i\}\ \cup\ \{c_{ik}:j<k\leq 6\}.$
+    $\mathscr{D}_{U^{\prime}}(a_{i}c_{ij}b_{j})=\{a_{h}:1\leq h<i\}\ \cup\ \{c_{ik}:j<k\leq 6\}.$
   • (ii)
     If $1<j<i<6$, then the triple $a_{i}c_{ij}b_{j}$ has level 4+ij4+i-j, and we have
-    $\text{\rm Res}_{U^{\prime}}(a_{i}c_{ij}b_{j})=\{a_{h}:1\leq h<i\}\ \cup\ \{c_{ki}:j<k\leq 6\text{\ and\ }k\neq i\}.$
+    $\mathscr{D}_{U^{\prime}}(a_{i}c_{ij}b_{j})=\{a_{h}:1\leq h<i\}\ \cup\ \{c_{ki}:j<k\leq 6\text{\ and\ }k\neq i\}.$
   • (iii)
-    If $M:=\{\{i_{1},j_{1}\},\{i_{2},j_{2}\},\{i_{3},j_{3}\}\}$ is a perfect matching of $[6]$, then the triple $c_{i_{1}j_{1}}c_{i_{2}j_{2}}c_{i_{3}j_{3}}$ has level $(\ell(\tau)+9)/2$, where $\tau$ is the fixed point free involution associated to $M$ and $\ell(\tau)$ is the length of $\tau$.
+    If $M:=\{\{i_{1},j_{1}\},\{i_{2},j_{2}\},\{i_{3},j_{3}\}\}$ is a perfect matching of $[6]$, then the triple $c_{i_{1}j_{1}}c_{i_{2}j_{2}}c_{i_{3}j_{3}}$ has level $(\ell(\tau)+9)/2$, where $\tau$ is the fixed-point-free involution associated to $M$ and $\ell(\tau)$ is the length of $\tau$.
 We have
-    $\text{\rm Res}_{U^{\prime}}(c_{i_{1}j_{1}}c_{i_{2}j_{2}}c_{i_{3}j_{3}})=\{a_{i}:1\leq i\leq 6\}\ \cup\ \{c_{pq}:\tau(p)<q\text{\ and\ }\tau(q)<p\}.$
+    $\mathscr{D}_{U^{\prime}}(c_{i_{1}j_{1}}c_{i_{2}j_{2}}c_{i_{3}j_{3}})=\{a_{i}:1\leq i\leq 6\}\ \cup\ \{c_{pq}:\tau(p)<q\text{\ and\ }\tau(q)<p\}.$
 Proof.
 For each triple $T\in\Omega_{U^{\prime}}$, we have
-$\text{\rm Res}_{U^{\prime}}(T)=\{\beta\in U^{\prime}:s_{\gamma}(\beta)>0\text{ for all }\gamma\in T\}=\bigcap_{\gamma\in T}(U^{\prime}\setminus\{\beta\in U^{\prime}:s_{\gamma}(\beta)<0\}),$
+$\mathscr{D}_{U^{\prime}}(T)=\{\beta\in U^{\prime}:s_{\gamma}(\beta)>0\text{ for all }\gamma\in T\}=\bigcap_{\gamma\in T}(U^{\prime}\setminus\{\beta\in U^{\prime}:s_{\gamma}(\beta)<0\}),$
 where each set appearing in the intersection can be calculated using Lemma 5.16.
-The descriptions of the residues follow from the lemma by straightforward computations.
+The descriptions of the generalized Rothe diagrams follow from the lemma by straightforward computations.
 Taking cardinalities of the generalized Rothe diagrams yields the levels of the triples in (i) and (ii), and the claim about the level in (iii) follows from Theorem 3.11 (v).
@@
 First, we identify the 27 roots of $U^{\prime}$ with the 27 square cells in the array shown in Figure 6 (a) as follows: identify the cells in the top row with $a_{1},a_{2},\ldots,$ and $a_{6}$ from left to right, the cells in the rightmost column with $b_{1},b_{2},\ldots,$ and $b_{6}$ from top to bottom, and the unique cell in the $(a_{j},b_{i})$-position (i.e., the cell in the same row as $b_{i}$ and same column as $a_{j}$) with $c_{ij}=c_{ji}$.
-If we represent each $T_{i}$ by marking the elements of $T$ with solid black dots, then the residues of $T_{i}$ are the cells marked with a star.
-For an element of the form $T=a_{i}c_{ij}b_{j}$ where $j<i$, such as $T_{1}$, Theorem 5.17 (ii) implies that $\text{\rm Res}_{U^{\prime}}(T)$ consists of the cells to the left of $a_{i}$, or below $c_{ji}$, or to the left of $b_{i}$ (which is below $b_{j}$ and not in $T$).
-For an element of the form $T=c_{i_{1}j_{1}}c_{i_{2}j_{2}}c_{i_{3}j_{3}}\in\Omega_{U^{\prime}}$, such as $T_{2}$, one can show using Theorem 5.17 (iii) that $\text{\rm Res}_{U^{\prime}}(T)$ consists of all the six roots in the top row and the roots $c_{ij}$ in the 27-cell grid that are neither attacked from the right nor from below by any of the roots in $T$ or their mirror images across the diagonal consisting of the $(a_{k},b_{k})$-positions for $k\in\{1,2,\ldots,6\}$.
+If we represent each $T_{i}$ by marking the elements of $T$ with solid black dots, then the roots dominating $T_{i}$ are the cells marked with a star.
+For an element of the form $T=a_{i}c_{ij}b_{j}$ where $j<i$, such as $T_{1}$, Theorem 5.17 (ii) implies that $\mathscr{D}_{U^{\prime}}(T)$ consists of the cells to the left of $a_{i}$, or below $c_{ji}$, or to the left of $b_{i}$ (which is below $b_{j}$ and not in $T$).
+For an element of the form $T=c_{i_{1}j_{1}}c_{i_{2}j_{2}}c_{i_{3}j_{3}}\in\Omega_{U^{\prime}}$, such as $T_{2}$, one can show using Theorem 5.17 (iii) that $\mathscr{D}_{U^{\prime}}(T)$ consists of all the six roots in the top row and the roots $c_{ij}$ in the 27-cell grid that are neither attacked from the right nor from below by any of the roots in $T$ or their mirror images across the diagonal consisting of the $(a_{k},b_{k})$-positions for $k\in\{1,2,\ldots,6\}$.
 We have shaded the cell $b_{i}=b_{4}$ in Figure 6 (a) and the aforementioned mirror images in Figure 6 (b).
 $665544332211\bigstar\bigstar\bigstar\bigstar\bigstar\bigstar$
-(a) $\text{\rm Res}_{U^{\prime}}(a_{4}c_{24}b_{2})$
+(a) $\mathscr{D}_{U^{\prime}}(a_{4}c_{24}b_{2})$
 $665544332211\bigstar\bigstar\bigstar\bigstar\bigstar\bigstar\bigstar\bigstar$
-(b) $\text{\rm Res}_{U^{\prime}}(c_{13}c_{25}c_{46})$
+(b) $\mathscr{D}_{U^{\prime}}(c_{13}c_{25}c_{46})$
 Figure 6. The generalized Rothe diagrams of two elements in $\Omega_{U^{\prime}}$
-Theorem 5.17 allows us to compute the quantum Hafnian $\text{\rm QHf}(U^{\prime})$ of $\Omega_{U^{\prime}}$ explicitly. We have
+Theorem 5.17 allows us to compute $\text{\rm QHf}(U^{\prime})$ explicitly. We have
 $\text{\rm QHf}(U^{\prime})=$ $\hphantom{+}q^{0}a_{1}c_{16}b_{6}+q^{1}a_{2}c_{26}b_{6}+q^{1}a_{1}c_{15}b_{5}+q^{2}a_{2}c_{25}b_{5}+q^{2}a_{3}c_{36}b_{6}+q^{2}a_{1}c_{14}b_{4}+q^{3}a_{2}c_{24}b_{4}$
@@
 $+q^{10}c_{14}c_{26}c_{35}+q^{11}c_{16}c_{24}c_{35}+q^{11}c_{15}c_{26}c_{34}+q^{12}c_{16}c_{25}c_{34}.$
-We now show that the Pfaffian $\text{\rm Pf}({U^{\prime}})=\text{\rm QHf}(U^{\prime},-1)$ associated to the QP-triple $T(E_{7},7)$ agrees with the invariant cubic polynomial associated with the Lie algebra (or Chevalley group) of type $E_{6}$ as constructed by Vavilov–Luzgarev–Pevzner [17].
+We now show that the generalized Pfaffian $\text{\rm Pf}({U^{\prime}})=\text{\rm QHf}(U^{\prime},-1)$ associated to the QP-triple $T(E_{7},7)$ agrees with the invariant cubic polynomial associated with the Lie algebra (or Chevalley group) of type $E_{6}$ as constructed by Vavilov–Luzgarev–Pevzner [22].
 This provides a nonrecursive combinatorial interpretation of the coefficients of the polynomial in terms of the parity of the cardinalities of the associated generalized Rothe diagrams.
 Theorem 5.19.
-The Pfaffian $\text{\rm Pf}(U^{\prime})$ associated to the QP-triple $T(E_{7},7)$ is the invariant cubic polynomial $F$ constructed in Vavilov–Luzgarev–Pevzner [17, §12].
+The generalized Pfaffian $\text{\rm Pf}(U^{\prime})$ associated to the QP-triple $T(E_{7},7)$ is the invariant cubic polynomial $F$ constructed in Vavilov–Luzgarev–Pevzner [22, §12].
 Proof.
-In [17, §12], Vavilov–Luzgarev–Pevzner define the minimal $W(E_{6})$-element $\{b_{6},c_{16},a_{1}\}\in\Omega_{U^{\prime}}$ from Example 5.15 to be the “distinguished triple”, which they denote as $x_{0}=(\lambda_{0},\mu_{0},\nu_{0})$.
+In [22, §12], Vavilov–Luzgarev–Pevzner define the minimal $W(E_{6})$-element $\{b_{6},c_{16},a_{1}\}\in\Omega_{U^{\prime}}$ from Example 5.15 to be the “distinguished triple”, which they denote as $x_{0}=(\lambda_{0},\mu_{0},\nu_{0})$.
 They define the invariant cubic polynomial $F$ to be the linear combination of the $45$ elements in the $W(E_{6})$-orbit of $x_{0}$ in which the coefficient of any monomial $x_{T}$ is $(-1)^{\ell(w)}$, where $w$ is an element of minimum length such that $w(x_{0})=x_{T}$.
@@
   • (i)
-    Vavilov–Luzgarev–Pevzner note in [17, §12] that if we apply a simple reflection $s_{i}$ on one of the monomials $x_{T}$ in the invariant cubic polynomial, then either $s_{i}$ fixes $x_{T}$, or the coefficients of $x_{T}$ and $s_{i}(x_{T})$ are the negatives of each other.
+    Vavilov–Luzgarev–Pevzner note in [22, §12] that if we apply a simple reflection $s_{i}$ to one of the monomials $x_{T}$ in the invariant cubic polynomial, then either $s_{i}$ fixes $x_{T}$, or the coefficients of $x_{T}$ and $s_{i}(x_{T})$ are the negatives of each other.
 This property also characterizes $\text{\rm Pf}(U^{\prime})$ and provides an alternative proof of Theorem 5.19.
   • (ii)
-    The cubic polynomial $F$ is written down explicitly in [17, Table 11]. This makes it possible, after translating notation, to verify Theorem 5.19 by hand.
+    The cubic polynomial $F$ is written down explicitly in [22, Table 11]. This makes it possible, after translating notation, to verify Theorem 5.19 by hand.
   • (iii)
@@
 In particular, $\Gamma$ is strongly regular with parameters $(45,32,22,24)$.
+Remark 5.21.
+It is possible to establish an analogue of Theorem 5.19 for $\Omega_{U}$ as follows.
+The 56 vectors $X_{ij}$ and $Y_{ij}$ arising from the set $U$ can be grouped into 28 pairs $i,j]:=\{X_{ij},Y_{ij}\}$.
+These pairs correspond naturally to the 28 bitangents to a plane quartic curve, and the 630 quadruples of $\Omega_{U}$ correspond to the 315 “tetrads of bitangents” of the quartic curve [6, Remark 5.5].
+One can modify the generalized Pfaffian of $\Omega_{U}$ in a precise way to produce an invariant quartic polynomial known as Dickson’s form (see [6, Definition 5.3] and [4]), thus providing the promised analogue of Theorem 5.19 for $\Omega_{U}$.
+We have chosen not to describe the modification here because it requires elaborate rescalings of coefficients for which we have no conceptual explanation.
+However, we note that the study of geometric objects such as the 630 tetrads corresponding to $\Omega_{U}$ and the 315 tetrads of bitangents has a long history dating back to the works of Cayley and Salmon (see, for example, [6, Remark 5.5] and [20, Section 262]).
+It would be interesting to know if the set $\Omega_{U}$ has further implications for these classical objects.
 6. Concluding remarks
 Remark 6.1.
-The original Rothe diagrams and the Rothe diagrams of fixed point free involutions are useful in the study of Schubert varieties.
+The original Rothe diagrams and the Rothe diagrams of fixed-point-free involutions are useful in the study of Schubert varieties.
 It would be interesting to know if this also applies to the generalized Rothe diagrams of Theorem 5.17, or to the generalized Rothe diagrams associated to the sets of all positive roots of types $E_{7}$ and $E_{8}$.
 Remark 6.2.
-In the setting of Section 5.3, the 56 vectors $X_{ij}$ and $Y_{ij}$ arising from the set $U$ can be grouped into 28 pairs $i,j]:=\{X_{ij},Y_{ij}\}$.
-These pairs correspond naturally to the 28 bitangents to a plane quartic curve, and the 630 quadruples of $\Omega_{U}$ correspond to the 315 “tetrads of bitangents” of the quartic curve [4, Remark 5.5].
-It is possible to modify the Pfaffian of $\Omega_{U}$ in a precise way to produce an invariant quartic polynomial known as Dickson’s form (see [4, Definition 5.3] and [2]), and this fact may be viewed as an analogue of Theorem 5.19 for $\Omega_{U}$.
-We have chosen not to describe the modification here because it requires elaborate rescalings of coefficients for which we have no conceptual explanation.
-However, we note that the study of geometric objects such as the 630 tetrads corresponding to $\Omega_{U}$ and the 315 tetrads of bitangents has a long history dating back to the works of Cayley and Salmon (see, for example, [4, Remark 5.5] and [15, Section 262]).
-It would be interesting to know if the set $\Omega_{U}$ has further implications for these classical objects.
-Remark 6.3.
 For any Coxeter system $(W,S)$ and any subset $I\subseteq S$, the left cosets of the parabolic subgroup $W_{I}\leq W$ form a quasiparabolic $W$-set whose level function $\ell_{I}$ measures the lengths of shortest coset representatives, and using $\ell_{I}$ one can deform the natural action of $W$ on the set $W/W_{I}$ to construct a module for the Iwahori–Hecke algebra $\mathcal{H}$ of $W$.
-The defining properties of a quasiparabolic set $X$ generalize the properties of $W/W_{I}$ and allows a similar deformation of the $W$-action on $X$ [14, Theorem 7.1].
+The defining properties of a quasiparabolic set $X$ generalize the properties of $W/W_{I}$ and allows a similar deformation of the $W$-action on $X$ [19, Theorem 7.1].
 It would be interesting to use $k$-roots to study the $\mathcal{H}$-modules obtained via the deformation of the quasiparabolic sets $X=\Omega_{U}$ constructed in this paper.
 Remark 6.3.
-The examples of quasiparabolic sets given in this paper all share some additional structural features that are not required by the definitions or needed for the proofs, which suggests the possible existence of an undiscovered underlying structure.
-For example, we have no general conceptual explanation for why the Poincaré polynomial $\sum_{R\in\Omega_{U}}q^{\rho_{U}(R)}$ always factorizes as the product of quantum integers.
-Furthermore, the largest such integer always turns out to be equal to $|U|/\kappa$, where $\kappa$ is the size of a maximal orthogonal subset of $U$.
-It is often (but not always) the case that $U$ has a partition whose blocks are elements of $\Omega_{U}$, which gives a combinatorial proof that $\kappa$ divides $|U|$.
-In particular, such a partition always exists for the quasiparabolic sets of the form $T(E_{n},p)$, except in the cases $T(E_{6},2)$, $T(E_{7},2)$, and $T(E_{8},5)$.
-Remark 6.5.
-One can create a directed acyclic graph (DAG) from a QP-triple $(W,I,U)$ by taking $U$ to be the set of vertices, and installing an edge from $u_{1}$ to $u_{2}$ if $u_{2}-u_{1}$ is a positive root of $W_{I}$.
-If $T\in\Omega_{U}$ satisfies $\rho(T)=0$, then $T$ is a a kernel of the DAG, meaning that (a) no two elements of $T$ are joined by an edge, and (b) for all $v\in U\backslash T$, there is an edge from $v$ to some element $u\in U$.
-We can use the DAG to create a two-player game, as in Example 4.9 (ii), so that the kernel is the kernel in the sense of game theory [18, §51].
-The vertices of the graph are the game positions, the directed edges are the legal moves, and playing a move that leads into the kernel constitutes a winning strategy.
+For all the quasiparabolic sets $\Omega_{U}$ constructed in this paper, the Poincaré polynomial factors as a multiplicity-free product of quantum integers.
+The significance of the corresponding integers, i.e., the degrees of $U$, may deserve further study.
+In particular, it would be interesting to know why the degrees can be computed via the matroid associated to $\Omega_{U}$ as explained in Section 4.3.
+For the QP-triples $\Omega_{U}$ from Section 4.2, it is also natural to wonder whether the degrees of $U$ have interpretations in terms of the corresponding symmetric spaces and symmetric pairs.